### v1

# Chance Constrained PDE-Constrained Optimal Design Strategies Under High-Dimensional Uncertainty

Abstract
This study presents an advanced computational framework for the optimal design of thermal insulation components in buildings, utilizing silica aerogel porous materials. The framework aims to achieve superior thermal insulation while maintaining structural integrity of the component under stress concentrations. A multiphase continuum model is employed to simulate the thermomechanical behavior of the insulation system, governed by a set of coupled partial differential equations (PDEs). The design process explicitly accounts for spatially correlated uncertainties in the design parameters, particularly the spatially varying aerogel porosity, resulting in a high-dimensional, PDE-constrained optimization under uncertainty. The optimization problem is formulated using a risk-averse cost functional to balance insulation performance with uncertainty mitigation, incorporating statistical moments of the design objective. Additionally, chance constraints are imposed to limit the probability of stress exceeding critical thresholds. To address the challenges arising from high-dimensional parameters, the optimization leverages a second-order Taylor expansion of both the design objective and the chance constraint functions, combined with a low-rank approximation of the Hessian matrix for efficient evaluation of the generalized eigenvalue problem. This approach supports scalable computations, either directly or as a variance-reduction control variate for Monte Carlo estimations. Combined with a gradient-based optimization approach, we achieve a scalable solution algorithm with dimension-independent computational costs in terms of number of PDE solved. Two- and three-dimensional numerical experiments on the design of thermal breaks in buildings showcase the features of the proposed design under uncertainty framework with respect to accuracy, scalability, and computational cost.
keywords:
Thermal insulation , Optimal design , PDE-constrained optimization , chance constraint , high-dimensional parameters , uncertainty quantification
journal: International Journal for Numerical Methods in Engineering –
[…]
[inst1]organization=Department of Mechanical and Aerospace Engineering, University at Buffalo,city=Buffalo, state=NY, country=USA
1 Introduction
Silica aerogels have emerged as a premier class of porous materials for high-performance thermal insulation, widely utilized across diverse engineering applications due to their exceptionally low thermal conductivity [1, 2, 3]. Among these applications, aerogel-based insulation components for building envelopes hold significant promise, offering the potential to enhance energy efficiency, reduce carbon emissions, and promote environmental sustainability. However, the adoption of silica aerogels in such systems is hindered by their inherently low mechanical strength and susceptibility to failure under localized stress concentrations. Recent advancements in extrusion-based additive manufacturing may be able to address these limitations by enabling precise spatial control over material deposition, thereby enhancing the mechanical stability of components. These developments have paved the way for integrating silica aerogels into lightweight thermal break insulation elements [2, 4], which effectively interrupt heat transfer pathways and mitigate the adverse effects of thermal bridging at assembly interfaces [5, 6]. Nevertheless, significant uncertainties remain regarding the thermal and mechanical performance of additively manufactured aerogel components, posing challenges to their adoption as reliable and sustainable solutions for thermal break applications in building envelopes.
In recent years, significant efforts have been dedicated to developing computational methods for topology optimization governed by partial differential equations (PDEs) (e.g., [7, 8, 9, 10, 11, 12, 13, 14, 15]). However, the optimal design of additively manufactured aerogel thermal breaks poses two major computational challenges. The first challenge arises from uncertainties in design parameters, stemming from variations in material properties, external loads, and environmental conditions. In particular, the layer-by-layer manufacturing process of aerogel components introduces spatially correlated uncertainties along the layer deposition direction, necessitating the representation of these uncertainties as spatially varying random fields. When discretized, the resulting high-dimensional parameter space makes the optimization problem computationally prohibitive due to the “curse of dimensionality” and the need for repeated evaluations of nonlinear PDEs governing material behavior. The second challenge involves incorporating stress constraints into the optimization process, which introduces nonlinearity and numerical instability. Stress gradients often exhibit discontinuities, further complicating the numerical implementation. To mitigate these issues, global stress constraints that aggregate stress measures (e.g., von Mises stress) across the design domain are commonly employed, significantly reducing the number of constraints. However, selecting an efficient stress aggregation function with appropriate parameters is crucial to achieving computational efficiency while ensuring the detection of localized stress concentrations. Furthermore, stress measures are influenced by uncertainties in material properties and operating conditions, necessitating the incorporation of randomness into stress constraints during the optimization process.
To address uncertainty in optimization, one approach is to utilize a risk-averse formulation of the cost functional, that takes into account the statistical moments of the objective, e.g., [16, 17, 18, 19, 20, 21, 22, 11, 23, 24]. However, these methods face significant computational challenges when evaluating the statistical properties of objectives or conditional probabilities, especially in the context of PDE-constrained optimization within high-dimensional parameter spaces. Various approximations have been proposed to address these challenges. Ghattas et al. [25, 20, 19, 18] utilized Taylor expansions for scalable, dimension-independent solutions with significant computational savings. This linearization, previously applied in robust design [26, 27] and topology optimization [28], was extended by Kriegesmann et al. [21, 29, 30] with a first-order second-moment method for efficient variance gradient evaluation. Tan et al. [31] extended the Taylor approximation method to a multi-objective PDE-constrained optimization framework for multi-material systems under uncertainty. Their approach incorporates phase-field regularization to effectively manage spatial sparsity and the material interface. Chen and Royset [32] provided further analysis of approximation-based algorithms for PDE-constrained optimization under uncertainty. In the context of stress constraints, several studies have investigated the integration of these constraints into optimization frameworks for diverse materials and manufacturing processes [33, 34, 35, 36, 37, 14, 38, 39, 40, 41]. Significant efforts have also focused on incorporating uncertainty into stress-constrained optimization [42], addressing factors such as variable loading and heat sources. For instance, Silva et al. [35] developed a method to account for uncertainties in boundary conditions during manufacturing. Similarly, Oh et al. [43] proposed a topology optimization approach that incorporates uncertainties in loading positions.
Building on previous work, this study presents a scalable computational framework for risk-averse PDE-constrained optimization under spatially correlated and high-dimensional uncertainty, while incorporating chance constraints. The framework is applied to the design of thermal breaks in building envelope, aiming to maximize thermal insulation performance and ensure mechanical stability by mitigating stress concentrations through chance constraints. The forward model is based on a thermomechanical continuum mixture theory for porous materials. The PDE model parameters, such as solid and fluid conductivities and solid elastic modulus, have been determined through model calibration in prior studies. Material porosity, modeled as a spatially correlated random field, serves as the design parameter. Discretization of this field transforms the problem into a high-dimensional optimization under uncertainty To achieve target performance levels while mitigating the uncertainty, the framework employs a risk-averse formulation that integrates the mean and variance of the design objective. The computational burden of Monte Carlo moment estimation is mitigated using second-order Taylor approximations for the design objective and chance constraints. An inexact Newton-Conjugate Gradient algorithm with convergence properties independent of problem dimension is utilized to ensure computational scalability—measured by the number of PDE solutions—regardless of the dimensionality of the design parameter space. The effectiveness and computational cost of the framework are explored through two- and three-dimensional numerical examples, including a case study of a beam-thermal break component in the building envelope. Compared to our earlier work in [31], the key contributions of this study are as follows: (1) Incorporation of probabilistic constraints in the optimization framework to mitigate stress concentrations. (2) Utilization of second-order Taylor approximations as a control variate for Monte Carlo estimation of the design objective, enabling more accurate optimal designs in high-variance uncertainty scenarios. (3) Implementation of an inexact Newton-Conjugate Gradient algorithm, facilitating parallelization and accelerating PDE-constrained optimization calculations. (4) Extension of the numerical framework to three-dimensional design of insulation components.
The rest of the paper is organized as follows: Section 2 presents the optimal design under uncertainty problem, including the thermomechanical forward model, the modeling of spatially correlated uncertainty in the design parameter, and the formulation of mean-variance optimization. Section 3 is dedicated to the approximation methods, encompassing Taylor approximation and the Monte Carlo estimate with control variate of the moments of the design objective. Section 4 describes the gradient-based optimization method and explains the computation of the gradient of the cost functional concerning the design parameter. Section 5 presents the numerical experiments on the optimal design of thermal breaks, followed by the conclusions and future works in Section 6.
2 Problem formulation: Design of aerogel thermal break under uncertainty
This study investigates the optimal design of building insulation components using silica aerogel materials. Figure 1 illustrates the schematic representation of a building thermal break, consisting of a concrete beam integrated with an aerogel insulation system, along with the thermal and mechanical boundary conditions employed in the numerical experiments. Concrete beams significantly influence the thermal performance of building envelopes due to their high thermal conductivity, which creates preferential pathways for heat transfer. The inclusion of silica aerogel thermal breaks disrupts this conduction, effectively mitigating thermal bridging from the interior to the exterior of the building envelope. However, thermal expansion of the concrete beam can induce stress concentrations within the silica aerogel insulation component, potentially leading to mechanical failure and compromising the thermal break’s effectiveness. In this section, we introduce the problem of optimal design under uncertainty, leveraging a forward model of aerogel and concrete materials governed by a system of partial differential equations (PDEs). We then describe how the uncertainty in design parameters are modeled, as well as formulation of design problem.
Figure 1: The domain of a beam-insulator system used for numerical experiments, indicating both the mechanical (left) and thermal (right) boundary conditions.
2.1 Forward thermomechanical model and design objectives
For silica aerogel materials, we utilize a thermomechanical model based on continuum mixture theory [44]. The model accounts for both an incompressible solid aerogel phase and a compressible fluid (gaseous) phase within the material. According to the mixture theory, each phase independently adheres to its own motion, balance laws, internal energy, and entropy, while collectively following the overall balance laws governing the entire mixture. As a result, a system of four partial differential equations is derived, which governs the heat transfer in the solid and fluid phases,and the deformation in the solid phase and the evolution of pore pressure. For our problem of designing thermal breaks, we focus on steady-state and static conditions, which simplify the governing equations to the following PDEs,
$\begin{cases}-\nabla\cdot(\phi_{s}\,\kappa_{s}\,\nabla\theta_{s})+h(\theta_{s}-\theta_{f})=0\quad&\text{in}\quad\Omega,\\ -\nabla\cdot(\phi_{f}\,\kappa_{f}\,\nabla\theta_{f})-h(\theta_{s}-\theta_{f})=0\quad&\text{in}\quad\Omega,\\ -\phi_{s}\,\kappa_{s}\,\nabla\theta_{s}=h_{\mathrm{air}}(\theta_{s}-\theta_{\mathrm{amb}})\quad&\text{on}\quad\Gamma_{1}\cup\Gamma_{4},\\ -\phi_{f}\,\kappa_{f}\,\nabla\theta_{f}=h_{\mathrm{air}}(\theta_{f}-\theta_{\mathrm{amb}})\quad&\text{on}\quad\Gamma_{1}\cup\Gamma_{4},\\ \end{cases}$ (1)
$\begin{cases}D\,p+(\nabla\cdot\mathbf{u}_{s})=0\quad&\text{in}\quad\Omega,\\ -\nabla\cdot\mathbf{T}^{\prime}_{s}-(2\,\phi_{f}-1)\,\nabla p=0\quad&\text{in}\quad\Omega,\\ \mathbf{u}_{s}=\bar{\mathbf{u}}_{s}\quad&\text{on}\quad\Gamma_{1}\cup\Gamma_{3},\\ \nabla\mathbf{u}_{s}.\mathbf{n}=0\quad&\text{on}\quad\Gamma_{2}\end{cases}$ (2)
where model parameters are $\theta=(\kappa_{s},\kappa_{f},D,\mathbf{E}_{s})$ and states are $u=(\theta_{s},\theta_{f},\mathbf{u_{s}},p)$. The state variables $\theta_{s}$, $\theta_{f}$, $\mathbf{u}_{s}$, and $p$ represent the solid and fluid temperatures, solid displacement, and fluid pressure, respectively, defined over the domain $\Omega$. Here, $\phi_{f}(\mathbf{x})$ represents the volume fraction of fluid (porosity), and $\phi_{s}(\mathbf{x})$ represents the volume fraction of solid, which can alternatively be written as $1-\phi_{f}(\mathbf{x})$. The solid displacement is subject to fixed values $\bar{\mathbf{u}}_{s}$ at the boundaries $\Gamma_{1}$ and $\Gamma_{3}$. A roller boundary condition is implemented at $\Gamma_{2}$. At $\Gamma_{1}$ and $\Gamma_{4}$, we consider convective heat transfer with a coefficient $h_{\mathrm{air}}$ and ambient temperature $\theta_{\mathrm{amb}}$. A Neumann type boundary condition with no heat flux is implemented at $\Gamma_{2}$. In the above equations, $\Omega\subset\mathbb{R}^{d}$ is an open and bounded domain with Lipschitz boundary $\partial\Omega$, and $\Gamma_{1},\Gamma_{3},\Gamma\subset\partial\Omega$ denote the Dirichlet and $\Gamma_{2}\subset\partial\Omega$ indicate the Neumann portions of $\partial\Omega$ such that $\Gamma_{1}\cup\Gamma_{2}\cup\Gamma_{4}=\partial\Omega$, $\Gamma_{1}\cap\Gamma_{2}=\emptyset$, $\Gamma_{3}\cap\Gamma_{4}=\Gamma_{3}$, and $\Gamma_{1}\cap\Gamma_{4}=\emptyset$. In (2), the effective solid stress is
$\mathbf{T}^{\prime}_{s}=2\,\mu\,\mathbf{E}_{s}+\lambda\,\mathrm{tr}(\mathbf{E}_{s})\,\mathbf{I}\quad\text{with}\quad\mathbf{E}_{s}=\frac{1}{2}\big{(}\nabla\mathbf{u}_{s}+(\nabla\mathbf{u}_{s})^{T}\big{)},$ (3)
where $\mathbf{E}_{s}$ is the solid strain and $\lambda$ and $\mu$ are the Lamé constants. Determining the model parameters $\kappa_{s}$, $\kappa_{f}$, $h$, $K$, $C$, and $\mu$ from experimental data of silica aerogel is demonstrated in [45] and assumed to be known. for the concrete beam we consider the following thermoelastic model,
$\begin{cases}-\nabla\cdot(\,\kappa_{b}\,\nabla\theta_{b})=0\quad&\text{in}\quad\Omega_{b},\\ \kappa_{b}\,\nabla\theta_{b}=h_{\mathrm{air}}(\theta_{b}-\theta_{\mathrm{amb}})\quad&\text{on}\quad\Gamma_{3},\\ \nabla.\mathbf{T}_{b}=0&\text{in}\quad\Omega_{b}\\ \end{cases}$ (4)
where $\kappa_{b}$ is thermal conductivity of beam, $\theta_{b}$ represents temperature of beam and $\mathbf{T}_{b}$ is the beam stress. In (4), the stress in beam is
$\mathbf{T}_{b}=2\mu_{b}\mathbf{E}_{b}+\lambda_{b}\mathrm{tr}(\mathbf{E}_{b})\mathbf{I}-\alpha_{T}(3\lambda_{b}+2\mu_{b})(\theta_{b}-\theta_{0})\quad\text{with}\quad\mathbf{E}_{b}=\frac{1}{2}\big{(}\nabla\mathbf{u}_{b}+(\nabla\mathbf{u}_{b})^{T}\big{)}$ (5)
where $\mathbf{E}_{b}$ is strain in the beam, and $\lambda_{b}$ and $\mu_{b}$ are the Lamé constants for the beam, $\alpha_{T}$ is coefficient of thermal expansion and $\theta_{0}$ is the reference temperature.
In the following sections, the forward state problems (1), (2) and (4) are expressed in an abstract form as,
$\mathcal{R}(\mathbf{u},m,d)=0\quad\mathrm{in\ }\mathcal{V}^{\prime},$ (6)
where, given the uncertain parameter $m\in\mathcal{M}$ and design parameter $d\in\mathcal{D}$, the goal is to find the state $\mathbf{u}=(\theta_{s},\theta_{f},\mathbf{u}_{s},p,\mathbf{u}_{b})\in\mathcal{U}$ such that the PDE residual operator $\mathcal{R}(\cdot)$ is minimized to zero in $\mathcal{V}^{\prime}$. Here $\mathcal{M}$ and $\mathcal{D}$ represent infinite-dimensional Hilbert spaces, and $\mathcal{V}^{\prime}$ denotes the corresponding dual space.
Weak formulation
To derive the weak formulation of the governing equations, we start with the partial differential equations (PDEs) (1) ,(2) and (4). The problem can be stated as follows: Find $\mathbf{u}=(\theta_{s},\theta_{f},\theta_{b},\mathbf{u}_{s},\mathbf{u}_{b},p)\in\mathcal{U}$, such that
$-\big{<}\phi_{s}\kappa_{s}\nabla\theta_{s},\nabla z_{s}\big{>}-\big{<}h(\theta_{s}-\theta_{f}),z_{s}\big{>}-\big{<}\phi_{s}h_{\mathrm{air}}(\theta_{s}-\theta_{\mathrm{amb}}),z_{s}\big{>}$ $=$ $0,$
$-\big{<}\phi_{f}\kappa_{f}\nabla\theta_{f},\nabla z_{f}\big{>}+\big{<}h(\theta_{f}-\theta_{f}),z_{f}\big{>}-\big{<}\phi_{f}h_{\mathrm{air}}(\theta_{f}-\theta_{\mathrm{amb}}),z_{f}\big{>}$ $=$ $0,$
$-\big{<}\kappa_{b}\nabla\theta_{b},\nabla z_{b}\big{>}-\big{<}h_{\mathrm{air}}(\theta_{b}-\theta_{\mathrm{amb}}),z_{b}\big{>}$ $=$ $0,$
$-\big{<}(\lambda+\frac{1-2\phi_{f}}{D})(\nabla\cdot\mathbf{u}_{s})\mathbf{I}+\mu\big{(}\nabla\mathbf{u}_{s}+(\nabla\mathbf{u}_{s})^{T}\big{)},\boldsymbol{w}_{u}\big{>}+\big{<}\mathbf{t},\boldsymbol{w}_{u}\big{>}$ $=$ $0,$
$-\big{<}(\lambda_{b})(\nabla\cdot\mathbf{u}_{b})\mathbf{I}+\mu_{b}\big{(}\nabla\mathbf{u}_{b}+(\nabla\mathbf{u}_{b})^{T}\big{)},\boldsymbol{w}_{b}\big{>}+\big{<}\mathbf{t},\boldsymbol{w}_{b}\big{>}$ $=$ $0,,$
[…]
$r(\mathbf{u},\mathbf{v},m,d):=\,_{\mathcal{V}}\langle\mathbf{v},\mathcal{R}(\mathbf{u},m,d)\rangle_{\mathcal{V}^{\prime}}=0\quad\forall\mathbf{v}\in\mathcal{V}$ (7)
where ,$\,{}_{\mathcal{V}}\langle\cdot,\cdot\rangle_{\mathcal{V}^{\prime}}$ represent the duality pairing between the $\mathcal{V}$ and $\mathcal{V}^{\prime}$.
Figure 2: Finite element solution of the forward model for the beam-insulator system with constant porosity $\phi_{f}=0.7$, depicting the solid temperature $\theta_{s}$, fluid temperature $\theta_{f}$, displacement magnitude $||\mathbf{u}_{s}||$, and von Mises stress $T_{\text{VM}}$.
Figure 2 presents the deterministic finite element solution of the forward model for the beam-insulator system under constant porosity $\phi_{f}=0.7$, utilizing the model parameters listed in Table 1. Thermal expansion of the thermoelastic concrete beam induces stress concentrations within the aerogel thermal break, risking mechanical failure. This can be mitigated by varying the aerogel’s porosity across the insulation, maintaining its effectiveness in interrupting heat transfer between the building interior and exterior.
Design objectives
The design objective $Q$ is defined as thermal compliance of the aerogel thermal break component over domain $\Omega$,
$Q$ $=$ $\frac{1}{2}\sum_{i=s,f}\big{<}\phi_{i}\,\kappa_{i}\,\nabla\theta_{i},\nabla\theta_{i}\big{>}+{\big{<}\phi_{i}h_{\mathrm{air}}(\theta_{i}-\theta_{\mathrm{amb}}),\theta_{i}\big{>}},$
The design problem focuses on determining the spatial distribution of the fluid volume fraction (porosity), $\phi_{f}(\mathbf{x})$, to minimize stress concentrations within the thermal break, thereby preventing mechanical failure while preserving its thermal insulation performance.
2.2 Uncertain design parameter
Due to the inherent challenges in precisely controlling the properties of additively manufactured aerogel, considerable uncertainty arises in the porosity values within the thermal break. To effectively model uncertainty, a design parameter as a spatial field $d\in[0,1]$ is introduced, employing a sigmoid mapping function, ensuring that porosity values asymptotically remain within the physically valid range, maintaining near-linear behavior within $\phi_{f}\in[90\%,10\%]$, and minimize the impact on the spatial correlation length, such that [31]
$\phi_{f}\equiv\text{sigmoid}\left(d+m\right).$ (9)
The uncertain parameter $m$ is represented as a Gaussian random field characterized by the Matérn covariance kernel. Specifically, $m$ is distributed as $\mathcal{N}(\bar{m},\mathcal{C})$ where $\bar{m}$ is the mean and $\mathcal{C}$ is the covariance operator. The sampling of Gaussian random field $m$ is equivalent to solving the elliptic equation with a robin boundary condition,
$-\gamma\nabla.(\boldsymbol{\Theta}\nabla m)+\delta m=$ $\>\dot{W}\quad\text{in}\quad\Omega$ (10)
$m+1.42\>\nabla m.\>\boldsymbol{n}=$  0onΓ $\>0\quad\text{on}\quad\Gamma$
where $\nabla$ is gradient operator,$\Delta$ is the Laplace operator and $\mathbf{n}$ is the outward unit normal vector. Here, $W$ represents the spatial White Gaussian noise with a unit variance, and the parameters $\gamma$ and $\delta$ control the variance $\sigma^{2}={1}/{4\pi\,\delta\,\gamma}$ and spatial correlation length $L_{c}=\sqrt{{8\gamma}/{\delta}}$. The factor ${\sqrt{\delta\,\gamma}}/{1.42}$ (Robin coefficient) is chosen to minimize boundary artifacts [46], and $\boldsymbol{\Theta}$ is a symmetric positive definite two-dimensional (2D) anisotropic tensor
[…]
2.3 Chance-constrained PDE-constrained optimal design under uncertainty
We consider the design problem as minimizing the design objective, thermal compliance in (LABEL:eq:qois), while accounting for uncertainty in objective through its explicit dependence on the uncertain parameter $m$, i.e., $Q(d,m)$. Extending the mean-variance optimization, e.g., [18, 19, 31], the cost functional for the proposed optimal design takes the form,
$\mathcal{J}(d)=\mathbb{E}[Q(d,m)]+\beta_{V}\,\mathbb{V}[Q(d,m)]+\beta_{R}\,R(d).$ (12)
To prevent mechanical failure of the aerogel thermal break due to stress concentration, the maximum von Mises stress $T_{\text{VM}}$ must remain below a critical threshold. This requirement is incorporated as a probabilistic (chance) constraint in the optimization problem, as $T_{\text{VM}}$ depends on uncertain parameters. However, due to the non-differentiable nature of the maximum stress, direct optimization poses significant challenges. This challenge is addressed using stress aggregation function based on p-norm [41], which is a smooth approximation of the maximum stress and enables gradient-based optimization techniques.The p-norm for the von Mises stress can be computed as
[…]
The chance constraint function can hence be written as,
$f=T_{cr}-T_{\text{pn}},$ (14)
where $T_{cr}$ is the limiting critical stress and $T_{\text{pn}}$ is the p-norm of the von Mises stress over the domain. To avoid stress concentration, we consider a chance constraint:
$P(f(m,d)\geq 0)\leq\alpha_{c},$ (15)
for a critical chance $10<\alpha_{c}<$ 1. The probability is given by
$P(f(m,d)\geq 0)$ $=$ $\mathbb{E}[\mathbb{I}_{[0,\infty]}(f(m,d))]$ (16)
[…]
where $\mathbb{I}_{[0,\infty]}(f(m,d))$ is an indicator function defined as:
$\mathbb{I}_{[0,\infty]}(f(m,d))=\begin{cases}1&\text{if }f(m,d)\geq 0\\ 0&\text{if }f(m,d)<0\\ \end{cases}$ (17)
uncertainty as,
$d_{\text{opt}}(\mathbf{x})=\min_{{d}}$ $\ \mathcal{J}({d})$ (18)
$\mathrm{s.t.\ }$ $\mathcal{R}(\boldsymbol{u},m,d)=0$
$P(f(m,d)\geq 0)\leq\alpha_{c}\quad\mathrm{in\ }\Omega$
Monte Carlo estimation
A common approach to compute the mean and variance terms in (12) is the Monte Carlo estimation for $N_{mc}$ samples,
$\mathbb{E}[Q]$ $\approx\frac{1}{N_{mc}}\sum_{i=1}^{N_{mc}}Q(m^{(i)}),$ (19)
$\mathbb{V}[Q]$ $\approx\Big{(}\frac{1}{N_{mc}}\sum_{i=1}^{N_{mc}}Q^{2}(m^{(i)})\Big{)}-\Big{(}\frac{1}{N}_{mc}\sum_{i=1}^{N_{mc}}Q(m^{(i)})\Big{)}^{2},$ (20)
where $\{m^{(i)}\}_{i=1}^{N_{mc}}$ are independent and identically distributed (i.i.d) random samples from the uncertain and spatially correlated parameter. Similarly, the chance constraint function can be approximated by
$P(f(.,d)\geq 0):=\frac{1}{M_{f}}\sum_{i=1}^{M_{f}}\mathbb{I}_{[0,\infty)]}(f(m_{i},d))$ (21)
To achieve an accurate approximation, particularly for $f_{M}$ for $\alpha$ close to 00, a large number of samples are required due to slow convergence of the Monte Carlo method. The convergence rate of Monte Carlo method is $\mathcal{O}\big{(}\frac{1}{\sqrt{N_{mc}}}\big{)}$, therefore for accurate computation of the cost functional (12), a considerable number of PDE model evaluations will be required, making the optimization under uncertainty problem computationally prohibitive.
3 Scalable framework for chance constrained optimal design
In this section, we present a computationally tractable approach by approximating the mean and variance of the design objective by means of low-order Taylor expansion [18, 19]. The quadratic Taylor approximation of $Q$ at the mean of the uncertain parameter $\bar{m}$, assuming $Q$ is sufficiently smooth and Fréchet differentiable with respect to uncertain parameter is derived as,
$Q_{\text{QUAD}}(m)$ $=$ $\sum_{l=0}^{2}\partial^{l}_{m}Q(\bar{m})\,(m-\bar{m})^{l}$ (22)
$=$ $\bar{Q}+\big{<}\bar{Q}^{m},\,m-\bar{m}\big{>}+\frac{1}{2}\big{<}\bar{Q}^{mm}\,(m-\bar{m}),\,m-\bar{m}\big{>}.$
where $\partial^{l}_{m}Q(\bar{m})$ is the $l$-th order tensor derivative at $\bar{m}$ and the objective $Q$ and its gradient and Hessian with respect to $m$, evaluated at mean $\bar{m}$, denoted as $\bar{Q}$, $\bar{Q}^{m}$ and $\bar{Q}^{mm}$, respectively. The close form for the expectation of the quadratic Taylor approximation in (22), considering $\mathbb{E}[\bar{Q}]=\bar{Q}$ and $\mathbb{E}[\big{<}\bar{Q}^{m},\,m-\bar{m}\big{>}]=0$, and Lemma 1 of [47], is obtained as
$\mathbb{E}[Q_{\text{QUAD}}]=\bar{Q}+\frac{1}{2}\mathrm{tr}\big{(}\bar{\mathcal{H}}_{c}\big{)},$ (23)
where ¯$\bar{\mathcal{H}}_{c}=\mathcal{C}\,\bar{Q}^{mm}$ denotes the covariance preconditioned Hessian computed at the mean of $m$ and $\mathrm{tr}(\cdot)$ is trace operator. Moreover, noting that variance is invariant with respect to its deterministic translation, e.g., $\mathbb{V}(m-\mathrm{const.})=\mathbb{V}(m)$), one can write,
$\mathbb{E}[(Q_{\text{QUAD}}(m)-\bar{Q})^{2}]=\mathbb{E}[(\big{<}\bar{Q}^{m},\,m-\bar{m}\big{>}+\frac{1}{2}\big{<}\bar{Q}^{mm}(m-\bar{m}),\,m-\bar{m}\big{>})^{2}],$ (24)
which, upon expansion, allows each term to be expressed as
$\mathbb{E}[(\big{<}\bar{Q}^{m},\,m-\bar{m}\big{>})^{2}]$ $=$ $\big{<}\bar{Q}^{m},\mathcal{C}\,\bar{Q}^{m}\big{>},$
$\mathbb{E}[(\big{<}\bar{Q}^{mm}(m-\bar{m}),\,m-\bar{m}\big{>})^{2}]$ $=$ $\mathrm{tr}(\bar{\mathcal{H}}_{c})^{2}+2\,\mathrm{tr}((\bar{\mathcal{H}}_{c})^{2}),$
$\mathbb{E}[\big{<}\bar{Q}^{m},\,m-\bar{m}\big{>}\,\big{<}\bar{Q}^{mm}(m-\bar{m}),\,m-\bar{m}\big{>}]$ $=$ $0$ (25)
Furthermore, the square of the first moment of $Q_{\text{QUAD}}(m)-\bar{Q}$ can be written as,
$\big{(}\mathbb{E}[Q_{\text{QUAD}}(m)-\bar{Q}]\big{)}^{2}=\big{(}\mathbb{E}[Q_{\text{QUAD}}(m)]-\mathbb{E}[\bar{Q}]\big{)}^{2}=\frac{1}{4}\big{(}\mathrm{tr}\big{(}\bar{\mathcal{H}}_{c}\big{)}\big{)}^{2}$ (26)
From (25) and (26), the variance of the quadratic Taylor approximation of $Q$ in (22) is obtained as,
$\mathbb{V}[Q_{\text{QUAD}}(m)]=\mathbb{V}[Q_{\text{QUAD}}(m)-\bar{Q}]=\big{<}\bar{Q}^{m},\mathcal{C}\,\bar{Q}^{m}\big{>}+\frac{1}{2}\mathrm{tr}((\bar{\mathcal{H}}_{c})^{2}).$ (27)
Consequently, the mean and variance of the design objective in (12) can be replaced by their corresponding quadratic Taylor approximations in (23) and (27), respectively, leading to the approximated design objective as
$\mathcal{J}_{\mathrm{QUAD}}(d)=\bar{Q}+\frac{1}{2}\mathrm{tr}\big{(}\bar{\mathcal{H}}_{c}\big{)}+\beta_{V}\,\bigg{(}\big{<}\bar{Q}^{m},\mathcal{C}\,\bar{Q}^{m}\big{>}+\frac{1}{2}\mathrm{tr}((\bar{\mathcal{H}}_{c})^{2})\bigg{)}+\beta_{R}\,R(d).$ (28)
Similar to (22), we can construct a second-order Taylor expansion of the constraint function $f$ at $\bar{m}$ as,
$f_{QUAD}(m)=\bar{f}+\big{<}\bar{f}^{m},\,m-\bar{m}\big{>}+\frac{1}{2}\big{<}\bar{f}^{mm}\,(m-\bar{m}),\,m-\bar{m}\big{>}.$ (29)
The probability $P(f(.,d)\geq 0)$ by Monte Carlo estimator (21) with the quadratic approximation yields,
$P(f(m,d))\geq 0)=\frac{1}{M_{f}}\mathbb{I}_{[0,\infty)}(f_{\text{QUAD}}(m_{i},d)).$ (30)
This approximation of the constraint function obviates the necessity for a large number of PDE solves to assess the probability utilizing sample averaging methods. In upcoming sections, we present the Lagrangian formalism to derive the $m$-gradient and $m$-Hessian, representing the gradient and Hessian of $Q$ with respect to $m$. We also introduce an efficient method for evaluating the covariance preconditioned Hessian, ¯$\bar{\mathcal{H}}_{c}$.
3.1 Analytical gradient and hessian of design objective with respect to the uncertain parameter
The Lagrangian functional, considering the adjoint variable $\mathbf{v}^{Q}$ as the Lagrange multiplier associated with the state equation with respect to the objective function $Q$, is defined as,
$\mathcal{L}(\mathbf{u},\mathbf{v}^{Q},m,d)=Q(\mathbf{u},m,d)+r(\mathbf{u},\mathbf{v}^{Q},m,d),$ (31)
where $r(\mathbf{u},\mathbf{v},m,d)$ is the weak form in (7). By imposing the condition that the first-order variation of the Lagrangian functional at $\bar{m}$ with respect to the adjoint variable $\mathbf{v}$ and the state variable $\mathbf{u}$ is equal to zero, we obtain, the state problem:
$\textit{Find $\mathbf{u}\in\mathcal{U}$, such that}\quad\langle\tilde{\mathbf{v}},\partial_{\mathbf{v}}\bar{r}\rangle=0,\quad\forall\tilde{\mathbf{v}}\in\mathcal{V},$ (32)
and adjoint problem:
$\textit{Find $\mathbf{v}\in\mathcal{V}$, such that}\quad\langle\tilde{\mathbf{u}},\partial_{\mathbf{u}}{\bar{r}^{Q}}\rangle=-\langle\tilde{\mathbf{u}},\partial_{\mathbf{u}}\bar{Q}\rangle,\quad\forall\tilde{\mathbf{u}}\in\mathcal{U}.$ (33)
In the above relations, $(\bar{\cdot})$ denotes the quantities computed at the mean of the uncertain parameter, and $(\tilde{\cdot})$ indicates directional variables. Directional variables are referred to the variables that define the direction along which perturbation is considered to compute the change in the functional to obtain the derivative of the functional. Also, here we define $r^{Q}=r(\mathbf{u},\mathbf{v}^{Q},m,d)$ for ease of notation. The gradient of $Q$ at $\bar{m}$ acting in the direction $\tilde{m}=m-\bar{m}$ is then given by,
$\langle\tilde{m},\bar{Q}^{m}\rangle=\langle\tilde{m},\partial_{m}\mathcal{L}\rangle.=\langle\tilde{m},\partial_{m}\bar{Q}+\partial_{m}{\bar{r}^{Q}}\rangle\quad\forall\tilde{m}\in\mathcal{M}$ (34)
Thus, evaluating $m$-gradient in (34) requires the solution of state problem (32) for $\mathbf{u}$ and the adjoint problem (33) for $\mathbf{v}$.
To solve for the $m$-Hessian, consider a meta-Lagrangian functional with incremental state $\hat{\mathbf{u}}$ and adjoint $\hat{\mathbf{v}}$
$\small\mathcal{L}^{H}(\mathbf{u},\mathbf{v}^{Q},m,d;\hat{\mathbf{u}}^{Q},\hat{\mathbf{v}}^{Q},\hat{m}^{Q})=\langle\hat{m}^{Q},\partial_{m}\bar{r}^{Q}+\partial_{m}\bar{Q}\rangle+\langle\hat{\mathbf{v}}^{Q},\partial_{\mathbf{v}}\bar{r}^{Q}\rangle+\langle\hat{\mathbf{u}}^{Q},\partial_{\mathbf{u}}\bar{r}^{Q}+\partial_{\mathbf{u}}\bar{Q}\rangle$ (35)
where the terms on the right-hand side are $m$-gradient, forward, and adjoint problems, respectively. Repeating the similar process of vanishing the first variation of the Lagrangian functional with respect to adjoint and state, we obtain the incremental state problem:
$\textit{Find $\hat{\mathbf{u}}\in\mathcal{U}$, such that}\quad\langle\tilde{\mathbf{v}},\partial_{\mathbf{vu}}\bar{r}^{Q}\,\hat{\mathbf{u}}^{Q}\rangle=-\langle\tilde{\mathbf{v}},\partial_{\mathbf{v}m}\bar{r}^{Q}\,\hat{m}^{Q}\rangle,\quad\forall\tilde{\mathbf{v}}\in\mathcal{V},$ (36)
and the incremental adjoint problem:
Find $\hat{\mathbf{v}}\in\mathcal{V}$, such that
$\langle\tilde{\mathbf{u}},\partial_{\mathbf{uv}}\bar{r}^{Q}\,\hat{\mathbf{v}}^{Q}\rangle=-\langle\tilde{\mathbf{u}},\partial_{\mathbf{uu}}\bar{r}^{Q}\,\hat{\mathbf{u}}+\partial_{\mathbf{uu}}\bar{Q}\,\hat{\mathbf{u}}^{Q}+\partial_{\mathbf{u}m}\bar{r}^{Q}\,\hat{m}^{Q}+\partial_{\mathbf{u}m}\bar{Q}\,\hat{m}^{Q}\rangle,\quad\forall\tilde{\mathbf{u}}\in\mathcal{U}.$ (37)
By taking the variation of $\mathcal{L}^{H}$ with respect to $m$ as zero, the Hessian $Q_{mm}(\bar{m})$ acting on $\hat{m}$ is computed as,
$\begin{split}\langle\tilde{m},\bar{Q}^{mm}\,\hat{m}^{Q}\rangle=&\langle\tilde{m},\partial_{m\mathbf{v}}\bar{r}^{Q}\,\hat{\mathbf{v}}^{Q}+\partial_{m\mathbf{u}}\bar{r}^{Q}\,\hat{\mathbf{u}}^{Q}+\partial_{m\mathbf{u}}\bar{Q}\,\hat{\mathbf{u}}^{Q}+\partial_{mm}\bar{r}^{Q}\,\hat{m}^{Q}\\ +\partial_{mm}\bar{Q}\,\hat{m}^{Q}\rangle\end{split}$ (38)
We remark that the last term on the right-hand side of (37) and the last two terms on the right-hand side of (38) are included due to the explicit dependence of the design objective (LABEL:eq:qois) on the uncertain parameter in our problem. Thus, evaluating the $m$-Hessian acting on $\hat{m}$ in (38) requires the solution of the incremental state problem (36) for $\hat{\mathbf{u}}$ and the incremental adjoint problem (37) for $\hat{\mathbf{v}}$.
Similarly, we can compute the gradient and Hessian action of $f$. The adjoint problem can be written as:
$\langle\tilde{\mathbf{u}},\partial_{\mathbf{u}}\bar{r}^{f}\rangle=-\langle\tilde{\mathbf{u}},\partial_{\mathbf{u}}\bar{f}\rangle,\quad\forall\tilde{\mathbf{u}}\in\mathcal{U}$ (39)
The gradient of $f$ at $\bar{m}$ can be defined as:
$\langle\tilde{m},\bar{f}^{m}\rangle=\langle\tilde{m},\partial_{m}\bar{f}+\partial_{m}\bar{r}^{f}\rangle\quad\forall\tilde{m}\in\mathcal{M}$ (40)
Then by solving incremental state problem,
[…]
and the incremental adjoint problem
[…]
We obtain the Hessian action of $f$ at $\tilde{m}$
$\begin{split}\langle\tilde{m},\bar{f}^{mm}\hat{m}^{f}\rangle=\langle\tilde{m},\partial_{m\mathbf{v}}\bar{r}^{f}\hat{\mathbf{v}}^{f}+\partial_{m\mathbf{u}}\bar{r}^{f}\hat{\mathbf{u}}^{f}+\partial_{m\mathbf{u}}\bar{f}\hat{\mathbf{u}}^{f}+\partial_{mm}\bar{r}^{f}\hat{m}^{f}+\partial_{mm}\bar{f}\hat{m}^{f}\rangle\quad&\\ \forall\tilde{m}\in\mathcal{M}\end{split}$ (43)
3.2 Monte Carlo with control variate
The approximation of $\mathbb{E}[Q]$ and $\mathbb{V}[Q]$ in the cost functional (12) using a quadratic Taylor series expansion may lack accuracy for the design problem,particularly for design problems where the variance of the uncertain parameter $m$ is large, as shown in [31]. To enhance accuracy, the quadratic approximation $Q_{\text{QUAD}}$ is leveraged as a control variate to reduce variance in the Monte Carlo estimators for the mean and variance. This involves correcting the moments of $Q_{\text{QUAD}}$ by applying a Monte Carlo method to estimate the mean and variance of the residual between $Q$ and its quadratic approximation $Q_{\text{QUAD}}$. We can use the same $M=M_{1}=M_{2}$ random samples for the approximation of the mean and variance. For mean of control variate approximation
$\mathbb{E}[Q_{CV}]\approx\mathbb{E}[Q]=\mathbb{E}[Q_{\text{QUAD}}]+\mathbb{E}[Q-Q_{\text{QUAD}}]:=Q(\bar{m})+\frac{1}{2}tr(\mathcal{H})$ (44)
$+\frac{1}{M}\sum_{i=1}^{M}\left(Q(m_{i})-Q(\bar{m})-\langle m_{i}-\bar{m},Q^{m}(\bar{m})\rangle-\frac{1}{2}\langle m_{i}-\bar{m},Q^{mm}(\bar{m})(m_{i}-\bar{m})\rangle\right).$
The variance can be written as
$\mathbb{V}[Q_{CV}]$ $=\mathbb{E}[{(Q_{\text{QUAD}}-Q(\bar{m}))}^{2}]+\mathbb{E}[{(Q-Q(\bar{m}))}^{2}-{(Q_{\text{QUAD}}-Q(\bar{m}))}^{2}]$ (45)
$-{(\mathbb{E}[Q_{\text{QUAD}}-Q(\bar{m})]+\mathbb{E}[(Q-Q(\bar{m}))-(Q_{\text{QUAD}}-Q(\bar{m}))])}^{2},$
which can be approximated by
$\mathbb{V}[Q_{CV}]:=\langle\mathcal{C}Q^{m}(\bar{m}),Q^{m}(\bar{m})\rangle+\frac{1}{4}{(tr(\mathcal{H}))}^{2}+\frac{1}{2}tr(\mathcal{H}^{2})$ (46)
$+\frac{1}{M}\sum_{i=1}^{M}\left({(Q(m_{i})-Q(\bar{m})}^{2}-{\left(\langle m_{i}-\bar{m},Q^{m}(\bar{m})\rangle+\frac{1}{2}\langle m_{i}-\bar{m},Q^{mm}(\bar{m})(m_{i}-\bar{m})\rangle\right)}^{2}\right)$
$-{\left(\frac{1}{2}tr(\mathcal{H})+\frac{1}{M}\sum_{i=1}^{M}\left(Q(m_{i})-\langle m_{i}-\bar{m},Q^{m}(\bar{m})\rangle-\frac{1}{2}\langle m_{i}-\bar{m},Q^{mm}(\bar{m})(m_{i}-\bar{m})\rangle\right)\right)}^{2}.$
The number of samples $M$ required to achieve the desired accuracy (and hence the computational cost) depends on the variance of the integrand and how closely the quadratic approximation matches the high-fidelity Monte Carlo estimation. A closer approximation reduces the required $M$, while a poorer match increases it. Notably, if the quadratic approximation is highly correlated with $Q$, fewer Monte Carlo samples are needed for effective variance reduction and accurate approximation.
3.3 Randomized trace estimator
The equations for the mean (23) and variance (27) of the design objectives evaluated using the quadratic approximation requires the trace of the covariance-preconditioned Hessian ¯$\bar{\mathcal{H}}c$, which typically comes out to be a large dense matrix. The traditional approach is based on Monte Carlo method, where the Hessian action is computed on a certain number of randomly chosen directions from a Gaussian distribution [18].
Despite adhering to the lower bounds on the required number of samples for accurate trace approximations, as outlined in [48], a large number of samples is still necessary. Given the large sample size needed at each optimization iteration, this approach becomes computationally prohibitive for the design under uncertainty problem addressed in this work. Here, we implement a randomized approximated eigen-decomposition for computing the trace and trace-squared using the dominant eigenvalues $\{\lambda_{n}\}_{n=1}^{N_{\mathrm{eig}}}$ of the covariance-preconditioned Hessian ¯$\bar{\mathcal{H}}_{c}$, expressed as,
$\mathrm{tr}\big{(}\bar{\mathcal{H}}_{c}\big{)}\approx\sum_{n\geq 1}^{N_{\mathrm{eig}}}\lambda_{n},\quad\mathrm{tr}\big{(}\bar{\mathcal{H}}_{c}^{2}\big{)}\approx\sum_{n\geq 1}^{N_{\mathrm{eig}}}\lambda_{n}^{2}.$ (47)
The above forms of trace estimators are justified by the symmetric positive definite property of the operator $\mathcal{H}_{c}$. It is important to note that the value of $N_{\mathrm{eig}}$ does not depend on the actual parameter dimension, but is dependent on the effective dimension of the parameter, e.g., [49, 50, 51, 52] it tends to be small when the covariance-preconditioned Hessian is low-rank or when its eigenvalues decay rapidly (see the numerical results of Section 5).
The eigenvalues $\lambda_{n}$ are obtained by solving the following generalized eigenvalue problem,
$\big{<}\zeta,\bar{Q}^{mm}\,\psi_{j}\big{>}=\lambda_{j}\big{<}\zeta,\mathcal{C}^{-1}\psi_{j}\big{>},\quad\forall\zeta\in\mathcal{M},\quad j=1,...,N_{\mathrm{eig}}$ (48)
where the eigenvectors $\psi_{j}$ exhibit orthonormality with $\mathcal{C}^{-1}$, $\big{<}\psi_{i},\mathcal{C}^{-1}\psi_{j}\big{>}=\delta_{ij},i,j=1,...,N_{\mathrm{eig}}$ and $\delta_{ij}$ denotes the Kronecker delta function. Using an $n$-degree finite element discretization, (48) is rewritten in the form of matrix as $\mathbf{M}\,\boldsymbol{\psi}=\lambda\,\mathbf{K}\,\boldsymbol{\psi}$ where $\mathbf{M},\mathbf{K}$ are $n$-by-$n$ symmetric matrices, and $\mathbf{K}$ is positive definite, $\boldsymbol{\psi}$ represents a vector of size $n$. For solving the resulting eigenproblem, we implement a randomized algorithm [53], discussed in Algorithm 1.
1
Input : Matrices $\mathbf{M},\mathbf{K}\in\mathbb{R}^{n\times n}$ obtained from finite element discretization, number of eigenpairs requested $N_{\mathrm{eig}}$, oversampling parameter $N_{\mathrm{p}}$
Output : Eigenpairs $(\lambda_{j},\psi_{j})$ for $j=1,\dots,N_{\mathrm{eig}}$
2
3Generate random sample and draw a Gaussian random matrix $\mathbf{\Omega}\in\mathbb{R}^{n\times(N_{\mathrm{eig}}+N_{\mathrm{p}})}$;
4
5Form the range matrix $\mathbf{Z}=\mathbf{K}^{-1}(\mathbf{M}\mathbf{\Omega})$;
6
7Perform QR factorization $\mathbf{Z}=\mathbf{Q}\mathbf{R}$ such that $\mathbf{Q}^{\top}\mathbf{M}\mathbf{Q}=\mathbf{I}_{N_{\mathrm{eig}}+N_{\mathrm{p}}}$;
8
9Compute the reduced matrix $\mathbf{Q}^{\top}\mathbf{K}\mathbf{Q}$;
10
11Compute the eigendecomposition $\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{\top}$, where $\boldsymbol{\Lambda}=\mathrm{diag}(\lambda_{1},\lambda_{2},\dots)$;
12
13Obtain the eigenpairs of the original problem:
$\lambda_{j}=\boldsymbol{\lambda}[j,j],\quad\psi_{j}=\mathbf{Q}\mathbf{U}[:,j],\quad j=1,\dots,N_{\mathrm{eig}}.$
Algorithm 1 Double-Pass Randomized Eigenvalue Solver
4 Gradient-based optimization
To solve the optimization problem (18), we employ the Inexact Newton Conjugate-Gradient optimization algorithm (INCG), leveraging an approximate Newton approach combined with a backtracking line search to ensure global convergence [20, 54]. In this method, the Hessian is approximated by evaluating it at the mean of the random field, while the gradient is computed precisely. The resulting Newton system is then solved approximately using a preconditioned conjugate gradient method in a matrix-free manner, allowing for efficient computation. However, applying INCG to solve (18) in a scalable manner requires special treatment of the chance constraints and efficient computation of the gradient of the cost functional with respect to high-dimensional design parameters, as discussed in this section.
4.1 Smooth approximation of the discontinuous function
The indicator function utilized in the chance constraint exhibits discontinuity at $f(m,d)=0$. To employ the gradient-based INCG, this discontinuous indicator function must be approximated with a smooth, continuous counterpart. One approach involves employing a logistic function as a smooth approximation of the indicator function [25] such that,
$\mathbb{I}_{[0,\infty]}(x)\approx l_{\omega}(x)=\frac{1}{1+e^{-2\omega x}},$ (49)
where a larger $\omega$ corresponds to a sharper transition at $x=0$ as shown in Figure 3. This smooth approximation tackles the non-differentiability of the chance constraint and makes the cost functional differentiable with respect to the design variable.
Figure 3: Approximation of the discontinuous indicator function $\mathbb{I}_{[0,\infty]}(x)$ with a smooth logistic function as given in (49) and its convergence with increasing $\omega$
4.2 Penalty Method
To impose the inequality constraint (15), a quadratic penalty method [55] is applied. The quadratic penalty function is defined as,
$S_{\gamma}(x)=\frac{\gamma}{2}{(max{\{0,x\}})}^{2}$ (50)
where $\gamma>0$ is a constant corresponding to the weight of the penalty. Using the penalty method, the chance-constrained problem can be rewritten as
$\underset{d\in[0,1]}{\text{min}}\mathcal{J}(d)+\mathcal{S}_{\gamma}(\mathbb{E}[l_{\omega}(f)]-\alpha_{c})$ (51)
This transformation simplifies the inequality-constrained problem by converting it into an unconstrained form.
4.3 Design gradients via adjoint
The Lagrangian formalism is employed to derive the gradient of the quadratic approximation of the design objective $\mathcal{J}_{\mathrm{QUAD}}(d)$ with respect to the design parameter, referred to as the d-gradient. This gradient provides critical information to the INCG optimizer, in order to facilitate gradient-based optimization processes. By substituting the trace estimators (47) into (28) and converting it into unconstrained problem using quadratic penalty (50) and smooth approximation (49), we have
$\mathcal{J}_{\mathrm{QUAD}}(d)$ $=\bigg{(}\bar{Q}+\frac{1}{2}\sum_{j\geq 1}^{N_{\mathrm{eig}}}\lambda_{j}^{Q}\bigg{)}+\beta_{V}\bigg{(}\big{<}\bar{Q}^{m},\mathcal{C}\,\bar{Q}^{m}\big{>}+\frac{1}{2}\sum_{j\geq 1}^{N_{\mathrm{eig}}}\left({\lambda_{j}^{Q}}\right)^{2}\bigg{)}+\beta_{R}\,R(d)$ (52)
$+S_{\gamma}(\mathbb{E}[l_{\omega}(f_{\text{QUAD}}(m,d))]-\alpha_{c})$
To evaluate the $d$-gradient, we consider a quadratic meta-Lagrangian functional
$\mathcal{L}_{\mathrm{QUAD}}$ $(d,\mathbf{u},\mathbf{v}^{*},\{{\lambda_{j}^{Q}}\},\{{\psi_{j}}^{Q}\},\{{\hat{\mathbf{u}}_{j}^{Q}}\},\{{\hat{\mathbf{v}}_{j}}^{Q}\},$ (53)
$\left(\mathbf{u}^{Q}\right)^{*},\left(\mathbf{v}^{Q}\right)^{*},\{\left(\lambda_{j}^{Q}\right)^{*}\},\{\left(\psi_{j}^{Q}\right)^{*}\},\{\left(\hat{\mathbf{u}}_{j}^{Q}\right)^{*}\},\{\left(\hat{\mathbf{v}}_{j}^{Q}\right)^{*}\},$
$\{{\lambda_{j}^{f}}\},\{{\psi_{j}}^{f}\},\{{\hat{\mathbf{u}}_{j}}^{f}\},\{{\hat{\mathbf{v}}_{j}}^{f}\},\left(\mathbf{u}^{f}\right)^{*},\left(\mathbf{v}^{f}\right)^{*},\{\left(\lambda_{j}^{f}\right)^{*}\},\{\left(\psi_{j}^{f}\right)^{*}\},\{\left(\hat{\mathbf{u}}_{j}^{f}\right)^{*}\},\{\left(\hat{\mathbf{v}}_{j}^{f}\right)^{*}\})$
$=\mathcal{J}_{\mathrm{QUAD}}(d)+\langle\mathbf{v}^{*},\partial_{\mathbf{v}}\bar{r}^{Q}\rangle$
$+\langle\left(\mathbf{u}^{Q}\right)^{*},\partial_{\mathbf{u}}\bar{r}^{Q}+\partial_{\mathbf{u}}\bar{Q}\rangle$
$+\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\langle\left(\psi_{j}^{Q}\right)^{*},(\bar{Q}^{mm}-{\lambda_{j}^{Q}}\mathcal{C}^{-1})\psi_{j}^{Q}\rangle$
$+\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\left(\lambda_{j}^{Q}\right)^{*}(\langle{\psi_{j}}^{Q},\mathcal{C}^{-1}{\psi_{j}}^{Q}\rangle-1)+\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\langle{\hat{\mathbf{v}}_{j}}^{{Q}^{*}},\partial_{\mathbf{vu}}\bar{r}_{Q}\,{\hat{\mathbf{u}}_{j}}^{Q}+\partial_{\mathbf{v}m}\bar{r}_{Q}\,{\psi_{j}}^{Q}\rangle$
$+\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\langle\left(\hat{\mathbf{u}}_{j}^{Q}\right)^{*},\partial_{\mathbf{uv}}\bar{r}_{Q}\,{\hat{\mathbf{v}}_{j}}^{Q}+\partial_{\mathbf{uu}}\bar{r}_{Q}\,\left(\mathbf{u}^{Q}\right)^{*}+\partial_{\mathbf{uu}}\bar{Q}\,{\hat{\mathbf{u}}_{j}}^{Q}+\partial_{\mathbf{u}m}\bar{Q}\,{\psi_{j}}^{Q}+\partial_{\mathbf{u}m}\bar{r}_{Q}\,{\psi_{j}}^{Q}\rangle$
$+\langle\left(\mathbf{u}^{f}\right)^{*},\partial_{\mathbf{u}}\bar{r}_{f}+\partial_{\mathbf{u}}\bar{f}\rangle$
$+\sum^{N_{\mathrm{eig}}^{f}}_{j=1}\langle\left(\psi_{j}^{f}\right)^{*},(\bar{f}^{mm}-{\lambda_{j}^{f}}\mathcal{C}^{-1})\psi_{j}^{f}\rangle$
$+\sum^{N_{\mathrm{eig}}^{f}}_{j=1}\left(\lambda_{j}^{f}\right)^{*}(\langle{\psi_{j}}^{f},\mathcal{C}^{-1}{\psi_{j}}^{f}\rangle-1)+\sum^{N_{\mathrm{eig}}^{f}}_{j=1}\langle\left(\hat{\mathbf{v}}_{j}^{f}\right)^{*},\partial_{\mathbf{vu}}\bar{r}_{f}\,{\hat{\mathbf{u}}_{j}}^{f}+\partial_{\mathbf{v}m}\bar{r}_{f}\,{\psi_{j}}^{f}\rangle$
$+\sum^{N_{\mathrm{eig}}^{f}}_{j=1}\langle\left(\hat{\mathbf{u}}_{j}^{f}\right)^{*},\partial_{\mathbf{uv}}\bar{r}_{f}\,{\hat{\mathbf{v}}_{j}}^{f}+\partial_{\mathbf{uu}}\bar{r}_{f}\,\left(\mathbf{u}^{f}\right)^{*}+\partial_{\mathbf{uu}}\bar{f}\,{\hat{\mathbf{u}}_{j}}^{f}+\partial_{\mathbf{u}m}\bar{f}\,{\psi_{j}}^{f}+\partial_{\mathbf{u}m}\bar{r}_{f}\,{\psi_{j}}^{f}\rangle$
where $(\cdot^{*})$ represents the adjoint variables, $(\hat{\cdot})$ denotes the incremental variables. The terms on the right hand side of (53), correspond to the following: the quadratic approximation of the design objective (28), the state (32) and adjoint (33) problems for design objective $Q$, the generalized eigenvalue problems (48) for design objective $Q$, the orthonormality constraint of the eigenvectors for design objective $Q$, the incremental state problem (36) of design objective $Q$ for eigenvector directions, the incremental adjoint problem (37) of design objective $Q$ for the directions of the eigenvectors, the adjoint problem (39) for chance constraint $f$, the generalized eigenvalue problems (48) for chance constraint $f$, the orthonormality constraint of the eigenvectors for chance constraint $f$, the incremental state problem (41) of chance constraint $f$ and the incremental adjoint problem (42) of chance constraint for the directions of the eigenvectors.
To compute the design gradient, we need to calculate five sets of functional derivatives to solve for the corresponding ten sets of adjoint variables
$\left(\mathbf{u}^{Q}\right)^{*},\left(\mathbf{u}^{f}\right)^{*},\left(\mathbf{v}^{Q}\right)^{*},\left(\mathbf{v}^{f}\right)^{*},\left(\psi_{j}^{Q}\right)^{*},\left(\psi_{j}^{f}\right)^{*},\left(\hat{\mathbf{u}}^{Q}_{j}\right)^{*},\left(\hat{\mathbf{u}^{f}_{j}}\right)^{*},\left(\hat{\mathbf{v}^{Q}_{j}}\right)^{*},\left(\hat{\mathbf{v}^{f}_{j}}\right)^{*}\}$ in (53), while the adjoint variable set for the eigenvalue set $\{{\lambda_{j}^{Q}}^{*},{\lambda_{j}^{f}}^{*}\}$ is not required due to its independency of the design parameter $d$.
Setting the first variation of (53) with respect to the eigenvalue set $\{\lambda_{j}\}$ to zero isolates the terms corresponding to the quadratic approximation of the design objective and the orthonormality constraint of the eigenvectors, resulting in:
$\{{\psi^{Q}_{j}}^{*}\}=\bigg{(}\frac{1}{2}+\beta_{V}\,\lambda_{j}^{Q}\bigg{)}\,\{\psi_{j}^{Q}\},\quad j=1,\cdots,N_{\mathrm{eig}}^{Q}.$ (54)
Setting the first variations of (53) to zero with respect to $\mathbf{v}^{Q}_{j}$ and $\mathbf{u}^{Q}_{j}$ results in the incremental state problem:
$\small\textit{Find $\hat{\mathbf{u}}^{Q}\in\mathcal{U}$, such that}\quad\big{<}\tilde{\mathbf{v}},\partial_{\mathbf{vu}}\bar{r}^{Q}\,\left(\mathbf{u}^{Q}_{j}\right)^{*}\big{>}=-\big{<}\tilde{\mathbf{v}},\partial_{\mathbf{v}m}\bar{r}^{Q}\,\left(\psi^{Q}_{j}\right)^{*}\big{>},\quad\forall\tilde{\mathbf{v}}\in\mathcal{V}.$ (55)
and the incremental adjoint problem:
$\footnotesize\begin{split}\textit{Find $\hat{\mathbf{v}}^{Q}\in\mathcal{V}$, such that}&\quad\\ \big{<}\tilde{\mathbf{u}},\partial_{\mathbf{uv}}\bar{r}^{Q}\,\left(\mathbf{v}^{Q}_{j}\right)^{*}\big{>}=-\big{<}\tilde{\mathbf{u}},\partial_{\mathbf{uu}}\bar{r}^{Q}\,{\left(\hat{\mathbf{u}}^{Q}_{j}\right)}^{*}+\partial_{\mathbf{uu}}\bar{Q}\,&{\left(\hat{\mathbf{u}}^{Q}_{j}\right)}^{*}+\partial_{\mathbf{u}m}\bar{r}^{Q}\,\left(\psi^{Q}_{j}\right)^{*}\big{>},\quad\forall\tilde{\mathbf{u}}\in\mathcal{U},\end{split}$ (56)
similar to (36) and (37), respectively. In combination with (54), we can formulate the adjoint variables for the incremental state and adjoint problems individually for each eigenvector as
$\{\left(\hat{\mathbf{u}}\right)^{*}_{j}\}$ $=\bigg{(}\frac{1}{2}+\beta_{V}\,\lambda^{Q}_{j}\bigg{)}\,\{\hat{\mathbf{u}}^{Q}_{j}\},\quad j=1,\cdots,N_{\mathrm{eig}}^{Q},$ (57)
$\{\hat{\mathbf{v}}^{*}_{j}\}$ $=\bigg{(}\frac{1}{2}+\beta_{V}\,\lambda^{Q}_{j}\bigg{)}\,\{\hat{\mathbf{v}}^{Q}_{j}\},\quad j=1,\cdots,N_{\mathrm{eig}}^{Q}.$ (58)
By vanishing the first variation of the Lagrangian functional (53) with respect to the adjoint or the state, we obtain the linear state problem: Find ${\mathbf{u}^{Q}}^{*}\in\mathcal{U}$, such that
$\langle\tilde{\mathbf{v}},\partial_{\mathbf{vu}}\bar{r}^{Q}\,\left(\mathbf{u}^{Q}\right)^{*}\rangle=$ $-2\beta_{V}\langle\tilde{\mathbf{v}},\partial_{\mathbf{v}m}\bar{r}^{Q}\,(\mathcal{C}\partial_{m}\bar{r}^{Q})\rangle$ (59)
$-\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\langle\tilde{\mathbf{v}},\partial_{\mathbf{v}mu}\bar{r}^{Q}\,\hat{\mathbf{u}}^{Q}_{j}\,\left(\psi^{Q}_{j}\right)^{*}+\partial_{\mathbf{v}mm}\bar{r}^{Q}\,\psi^{Q}_{j}\,\left(\psi^{Q}_{j}\right)^{*}\rangle$
$-\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\langle\tilde{\mathbf{v}},\partial_{\mathbf{vuu}}\bar{r}^{Q}\,\hat{\mathbf{u}}^{Q}_{j}\,\left(\hat{\mathbf{u}}^{Q}_{j}\right)^{*}+\partial_{\mathbf{vu}m}\bar{r}^{Q}\,\psi^{Q}_{j}\,\left(\hat{\mathbf{u}}^{Q}_{j}\right)^{*}\rangle,\quad\forall\tilde{\mathbf{v}}\in\mathcal{V},$
and linear adjoint problem: Find $\left(\mathbf{v}^{Q}\right)^{*}\in\mathcal{V}$, such that
$\langle\tilde{\mathbf{u}},\partial_{\mathbf{uv}}\bar{r}^{Q}\,\left(\mathbf{v}^{Q}\right)^{*}\rangle=$ $-\langle\tilde{\mathbf{u}},\partial_{\mathbf{u}}\bar{Q}\rangle-2\beta_{V}\langle\tilde{\mathbf{u}},\partial_{\mathbf{u}m}\bar{r}^{Q}\,(\mathcal{C}\partial_{m}\bar{r}^{Q})\rangle$ (60)
$-\langle\tilde{\mathbf{u}},\partial_{\mathbf{uu}}\bar{r}^{Q}\,\left(\mathbf{u}^{Q}\right)^{*}+\partial_{\mathbf{uu}}\bar{Q}\,\left(\mathbf{u}^{Q}\right)^{*}\rangle$
$-\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\langle\tilde{\mathbf{u}},\partial_{\mathbf{u}m\mathbf{v}}\bar{r}^{Q}\,\hat{\mathbf{v}}_{j}\,\left(\psi^{Q}_{j}\right)^{*}$
$+\partial_{\mathbf{u}m\mathbf{u}}\bar{r}^{Q}\,\hat{\mathbf{u}}_{j}\,\left(\psi^{Q}_{j}\right)^{*}+\partial_{\mathbf{u}mm}\bar{r}^{Q}\,\psi^{Q}_{j}\,\left(\psi^{Q}_{j}\right)^{*}\rangle$
$-\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\langle\tilde{\mathbf{u}},\partial_{\mathbf{uvu}}\bar{r}^{Q}\,\hat{\mathbf{u}}^{Q}_{j}\,\left(\hat{\mathbf{v}}^{Q}_{j}\right)^{*}+\partial_{\mathbf{uv}m}\bar{r}^{Q}\,\psi^{Q}_{j}\,\left(\hat{\mathbf{v}}^{Q}_{j}\right)^{*}\rangle$
$-\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\langle\tilde{\mathbf{u}},\partial_{\mathbf{uuv}}\bar{r}^{Q}\,\hat{\mathbf{v}}^{Q}_{j}\,\left(\hat{\mathbf{u}}^{Q}_{j}\right)^{*}+\partial_{\mathbf{uuu}}\bar{r}^{Q}\,\hat{\mathbf{u}}^{Q}_{j}\,\left(\hat{\mathbf{u}}^{Q}_{j}\right)^{*}$
$+\partial_{\mathbf{uuu}}\bar{Q}\,\hat{\mathbf{u}}^{Q}_{j}\,\left(\hat{\mathbf{u}}^{Q}_{j}\right)^{*}+\partial_{\mathbf{uuu}m}\bar{r}^{Q}\,\psi^{Q}_{j}\,\left(\hat{\mathbf{u}}^{Q}_{j}\right)^{*}\rangle,$
$\quad\forall\tilde{\mathbf{u}}\in\mathcal{U}.$
By vanishing the first variation of (53) with respect to the eigenvalue set $\{\lambda_{j}^{f}\}$, we get
$\{\left(\psi^{f}_{j}\right)^{*}\}=\bigg{(}\frac{1}{2}+\beta_{V}\,\lambda^{f}_{j}\bigg{)}\,\{\psi^{f}_{j}\},\quad j=1,\cdots,N^{f}_{\mathrm{eig}}.$ (61)
Equating to zero of the first variations of (53) with respect to $\mathbf{v}_{j}^{f}$ and $\mathbf{u}_{j}^{f}$ results in the incremental state problem:
$\textit{Find $\hat{\mathbf{u}}^{f}\in\mathcal{U}$, such that}\quad\big{<}\tilde{\mathbf{v}},\partial_{\mathbf{vu}}\bar{r}^{f}\,\mathbf{u}^{f*}_{j}\big{>}=-\big{<}\tilde{\mathbf{v}},\partial_{\mathbf{v}m}\bar{r}^{f}\,\psi^{{f}^{*}}_{j}\big{>},\quad\forall\tilde{\mathbf{v}}\in\mathcal{V}.$ (62)
and linear adjoint problem:
Find $\left(\mathbf{v}^{f}\right)^{*}\in\mathcal{V}$, such that (63)
$\langle\tilde{\mathbf{v}},\partial_{\mathbf{uv}}\bar{r}^{f}\,\left(\mathbf{u}^{f}\right)^{*}\rangle=$ $-\sum_{n=1}^{N_{eig}^{f}}\langle\tilde{\mathbf{v}},\partial_{\mathbf{v}m\mathbf{u}}\bar{r}^{f}\hat{\mathbf{u}}_{n}^{f}{\left(\psi_{n}^{f}\right)}^{*}+\partial_{\mathbf{v}mm}\bar{r}^{f}\psi_{n}^{f}{\left(\psi_{n}^{f}\right)}^{*}\rangle$
$-\sum_{n=1}^{N_{eig}^{f}}\langle\tilde{\mathbf{v}},\partial_{\mathbf{vuu}}\bar{r}^{f}\hat{\mathbf{u}}_{n}^{f}{\left(\hat{\mathbf{u}}_{n}^{f}\right)}^{*}+\partial_{\mathbf{vu}m}\bar{r}^{f}\psi_{n}^{f}{\left(\hat{\mathbf{u}}_{n}^{f}\right)}^{*}\rangle$
$-\langle\tilde{\mathbf{v}},\partial_{\mathbf{v}m}\bar{r}^{f}m^{f}\rangle$
$\quad\forall\tilde{\mathbf{u}}\in\mathcal{U}.$
Finally, with driving all the adjoint variables, the $d$-gradient is derived as
$\langle\tilde{d},D_{d}\mathcal{J}_{\mathrm{QUAD}}(d)\rangle=\langle\tilde{d},\partial_{d}\mathcal{L}_{\mathrm{QUAD}}\rangle.$ (64)
The methodology for calculating the design gradient is detailed in Algorithm 2. The computational cost required to evaluate the quadratic approximation of the cost functional, denoted as $\mathcal{J}_{\mathrm{QUAD}}(d)$ in (52), includes solving one potentially nonlinear state partial differential equation (PDE), one linear adjoint problem, and $22(N_{\mathrm{eig}}^{Q}+N_{o})+2(N_{\mathrm{eig}}^{f}+N_{o})+$ 2 linear PDEs, as prescribed by the double-pass randomized algorithm presented in Algorithm 1. Furthermore, obtaining the $d$-gradient involves solving $2(N_{\mathrm{eig}}^{Q}+N_{\mathrm{eig}}^{f}+3)$ additional linear PDEs.
1
Input : Design objective (12) and weak form of the PDF (7).
Output : Gradient of the quadratic approximation of the design objective $\mathcal{J}_{\mathrm{QUAD}}(d)$ with respect to the design parameter $z$.
2
3Solve for $m$-gradient (34);
4
5Solve for $m$-Hessian (38);
6
7Solve for $m_{f}$-gradient ;
8
9Solve for $m_{f}$-Hessian ;
10
11Solve for generalized eigenpairs with Algorithm 1;
12
13Compute $\mathcal{J}_{\mathrm{QUAD}}$ by (52);
14
15Solve the linear state problem (59);
16
17Solve the linear adjoint problem (63);
18
19Obtain the design gradient (64)
Algorithm 2 Analytical design gradient
5 Numerical Results
This section presents numerical experiments utilizing the proposed design under uncertainty framework for beam-insulation scenario in Figure 1 in both 2D and 3D. The objective is to achieve thermal insulation and mechanical stability avoiding stress concentration while mitigating uncertainty in the design process. The analyses presented in this section, investigate the accuracy and convergence of quadratic and Monte Carlo approximation of design objective and explores the effect of different limiting critical stress, critical chance, and uncertain parameter variance and spatial correlation on the optimal design, spatial pattern of $d_{\text{opt}}$, as well as the probability distributions of the design objectives $Q$. The parameters of the forward model are summarized in Table 1. The computational implementation of the framework relies on several open-source libraries like FEniCS [56] for the finite element solution of the forward model, SOUPy [57] for the quadratic approximation of the design objective and implementation of chance constraint, and hIPPYLib [54] for the Gaussian random field, trace estimator and INCG optimizer.
Parameter Value
$\kappa_{s}$ 0.477 $W/mK$
$\kappa_{f}$ 0.085 $W/mK$
$\kappa_{b}$ 5 $W/mK$
$\alpha_{T}$ $10^{-5}\>(1/K)$
$h$ 81059 $W/m^{2}K$
D 0.25 ×$\times 10^{-8}Pa$
$\lambda$ 6.77 $GPa$
$\mu$ 3.38 $GPa$
$\lambda_{b}$ 17.3 $GPa$
$\mu_{b}$ 11.5 $GPa$
Table 1: List of the forward model parameters and their values used in the design problem.
5.1 Spatially-correlated uncertain parameter
Refer to caption  Refer to caption  Refer to caption  Refer to caption
Refer to caption  Refer to caption  Refer to caption  Refer to caption
[…]
(a)                                              (b)
Refer to caption            Refer to caption
(c)                                              (d)
Figure 5: The effect of boundary on the Gaussian random field in (11) for the scenario used in Figure 7: (a) marginal variance of the anisotropic uncertain parameter $m$, and point-wise correlation structure centered at (b) point (0.5,0.75) (c) point (0.25,0.75) (d) point (0.25,0.25). Coordinates $(x,y)$ are defined with respect to the bottom left corner of the beam-insulator domaim. The uncertain parameter $m$ is characterized by a mean value of $\bar{m}=0$, a variance of $\sigma^{2}=0.5^{2}$ and a correlation length of $L_{c}=0.25$.
[…]
5.2 Quadratic and Monte Carlo approximations
To assess the accuracy and efficiency of the quadratic approximation and Monte Carlo estimations of mean and variance of design objective with a control variate on the optimal design solution, we perform comparative analyses. As the sample size and number of dominant eigenvalues increase, both methods converge to similar mean and variance values, consistent with standard Monte Carlo estimations, as shown in Figure 6. For large sample sizes, the estimated mean and variance from the standard Monte Carlo method (Figure 6(a)) and the Monte Carlo method with control variate (Figure 6(b)) are nearly identical. However, the quadratic approximation as a control variate achieves lower error bounds with significantly fewer samples, as evident in the box plots. Furthermore, with only $N_{tr}=25$ trace estimators, the quadratic approximation closely matches the Monte Carlo results (Figure 6(c)), leading to significant savings in computational cost without loss in accuracy.
Refer to caption  Refer to caption
Refer to caption  Refer to caption
Figure 6: Convergence plots corresponding to scenario in Figure 7 for the mean and variance of the design objective $Q$ : (a) standard Monte Carlo method (b) Monte Carlo with control variate and (c) Quadratic approximation. The uncertain parameter $m$ has the correlation length $L_{c}=0.05$ and a variance of $\sigma^{2}={0.25}^{2}$.
Figures 7 and 8 compare the effects of different variances and correlation lengths of the uncertain parameter $m$ on the optimal design $d_{\text{opt}}$ indicating the spatial distribution of aerogel porosity and the probability distributions of the design objective $Q$ representing the thermal insulation performance of the thermal break. The optimal design is evaluated using standard Monte Carlo, Monte Carlo with control variate, and a quadratic Taylor approximation of the mean and variance of $Q$. In Figure 7 (low variance and short correlation length), the morphologies of $d_{\text{opt}}$ and probability distributions of $Q$ are nearly identical across all three methods. However, in Figure 8 (high variance and long correlation length), the quadratic approximation predicts different optimal design pattern with higher $Q$, while Monte Carlo with control variate yields results closely matching standard Monte Carlo but with significantly fewer samples (PDF solved).
(a)                            (b)                            (c)
Figure 7: The top row shows the optimal design obtained for $\sigma=0.25$ for the three different approaches: (a) Sampled Average approximation (500 samples) (b) Monte Carlo correction (100 samples) and (c) Quadratic approximation ($N_{tr}=25)$. The bottom row shows the QoI distribution of the design objective $Q$ for the three different cases above. The uncertain parameter $m$ has a correlation length $L_{c}=0.05$ for all three cases.
Interestingly, while the spatial pattern of the optimal design in Figure 7 appears nearly symmetric, the higher uncertainty in Figure 8 results in optimal designs that consistently position stronger material (lower aerogel porosity) along the horizontal axis, across all three approximation methods. This pattern reflects the anisotropy of the uncertain parameter, modeled with a higher correlation length along the layer deposition direction (see Figure 4). Also, the optimal designs at mean $\tilde{m}=0$ are much similar in Figure 7 as compared to Figure 8 where there is a deviation in the design patterns, which shows that there is a trade-off favoring the reduction of the uncertainty over mean performance for our design formulation (12). These results demonstrate that the proposed optimization under uncertainty framework effectively captures spatial uncertainty, producing more reliable design solutions.
Figure 8: The top row shows the optimal design obtained for $\sigma=0.5$ for the three different approaches: (a) standard Monte Carlo using 500 samples, (b) Monte Carlo with control variate using 100 samples, and (c) Quadratic approximation using $N_{tr}=25$. The bottom row shows the QoI distribution of the design objective $Q$ for the three different cases above. The uncertain parameter $m$ has a correlation length of $L_{c}=0.25$ for all three cases.
5.3 Scalability with respect to number of design parameters
We assess the scalability of the algorithm for PDE-constrained optimization under uncertainty with respect to the design dimension, determined by the number of nodes in the spatial discretization of $d$, and governed by two primary factors: Firstly, the eigenvalue decay rate of the covariance-preconditioned Hessian, estimated via the randomized trace estimator, and secondly the convergence behavior of the INCG optimizer. Figure 9 illustrates the behavior of the chance constraint function in (14) with the number of steps in the continuation scheme, the gradient norm 64 and the absolute values of the generalized eigenvalues 48 for the last step of chance convergence discussed in Algorithm 3. These metrics are presented for two design parameter dimensions, corresponding to the optimal design problem depicted in Figure 7 (c).
(a)                                   (b)                                   (c)
Figure 9: Optimization solution corresponding to Figure 7 (c) for different dimensions of design parameter: (a) Convergence of the chance $P(f(m,d)\geq 0)$ with the number of steps, (b) convergence of the norm of the design gradient with number of iterations of INCG optimizer, (c) eigenvalue decay of the covariance-preconditioned Hessian
These results demonstrate that the quadratic approximation enables rapid decay of the critical chance, with the gradient norm converging below the tolerance threshold within 100 iterations for both mesh resolutions. The minimal dependence of convergence on mesh resolution highlights that the gradient-based INCG optimizer is effectively independent of the number of uncertain parameters. Additionally, Figure 9(c) shows a rapid eigenvalue decay, unaffected by the dimension of uncertain parameters, supporting the scalability of the randomized eigensolver used for trace approximation.
1st                     2nd                     10th                     20th
Refer to caption  Refer to caption  Refer to caption  Refer to caption
40th                   60th                   80th                   100th
Figure 10: Eigenfunctions of the covariance-preconditioned Hessian of the design objective corresponding to eigenvalue decay for the scenario in Figure 7 (c).
Figure 10 depicts the eigenfunctions of the covariance-preconditioned Hessian for the design objective. Eigenvectors associated with larger eigenvalues capture broader spatial scales of the design parameter, while those linked to smaller eigenvalues become increasingly oscillatory and less significant. Together, these dimension-independent factors suggest that leveraging quadratic Taylor approximations and stochastic optimization techniques enables a scalable approach for PDE-constrained optimal design under uncertainty with the enforcement of chance constraints, regardless of the dimensionality of the uncertain parameters.
5.4 Optimal design: effect of chance constraint
(a)                              (b)                              (c)
Figure 11: The top row shows the optimal design results obtained at the mean of the uncertain sample $\tilde{m}=0$ for different values of critical chance: (a) $\alpha_{c}=0.05$ (b) $\alpha_{c}=0.08$ and (c) $\alpha_{c}=0.1$. The bottom row represents the von Mises stress for the corresponding optimal designs in the top row. In all results $T_{cr}=22.5\>\text{MPa},L_{c}=0.25,\sigma=0.25,\beta_{V}=0.1$ and $\beta_{R}=1\times 10^{-5}$.
(a)                              (b)                              (c)
Figure 12: The top row shows the optimal design results obtained at the mean of the uncertain sample $\tilde{m}=0$ for different values of limiting critical stress: (a) $T_{cr}=20\>\text{MPa}$ (b) $T_{cr}=22.5\>\text{MPa}$ and (c) $T_{cr}=25\>\text{MPa}$. The bottom row represents the von Mises stress for the corresponding optimal designs in the top row. In all results $\alpha_{c}=0.05,L_{c}=0.25,\sigma=0.25,\beta_{V}=0.1$ and $\beta_{R}=1\times 10^{-5}$.
The chance constraint function defined in (14) incorporates parameters such as $\alpha_{c}$ and $T_{cr}$, which govern the enforcement of constrain to mitigate stress concentrations within the domain. In particular, the probability of chance $P(f(m,d)\geq 0)$ is limited by a threshold $\alpha_{c}$, where the function $f$ is the difference between the limiting critical stress $T_{cr}$ and the p-norm of the von Mises stress $T_{\text{pn}}$. To examine the influence of $\alpha_{c}$ on optimal design, a parametric study was conducted with constant parameters, as shown in Figure 11. The critical stress limit was set at $T_{cr}=22.5\>\text{MPa}$, while $\alpha_{c}$ was varied among 0.05, 0.08, and 0.1. This figure presents the optimal design, and von Mises stress distribution evaluated at the mean of the uncertain parameter $\tilde{m}=0$. The results indicate that increasing critical chance $\alpha_{c}$ expands the regions of mechanically stronger material (highlighted in red), effectively reducing localized stresses across the component. However, this improvement in mechanical stability comes at the expense of insulation performance of the thermal break.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 13: Porosity and states of the scenario at the optimal design shown in Figure 11 (b) obtained at the mean(first column) of the uncertain parameter $\tilde{m}=0$, three i.i.d. samples (second, third and fourth column) of $m$. The rows show the porosity $\phi_{f}$, solid displacement magnitude $||\mathbf{u}_{s}||$, magnitude of effective solid stress $||\mathbf{T}_{s}||$, , solid temperature $\theta_{s}$, fluid temperature $\theta_{f}$ and fluid pressure $p$.
The limiting critical stress, $T_{cr}$, defines the maximum permissible von Mises stress within the beam structure. Figure 12 demonstrates the influence of $T_{cr}$ on the optimal designs, evaluated at the mean of the uncertain parameter $\tilde{m}=0$, along with the corresponding von Mises stress distributions. The findings indicate that increasing the critical stress $T_{cr}$ results in the allocation of stronger insulating material (depicted in blue) across a larger region of the domain. This enhances the insulation performance but compromises the mechanical strength of the component. Thus, a higher value of $T_{cr}$ improves insulation efficiency at the expense of reduced mechanical stability. Figure 13 shows the state parameters and porosity field evaluated at the mean of the uncertain parameter $\tilde{m}=0$ and three samples for the optimal design in Figure 12 (b). The states include the solid displacement magnitude $||\mathbf{u}_{s}||$, magnitude of effective solid stress $||\mathbf{T}_{s}||$, , solid temperature $\theta_{s}$, fluid temperature $\theta_{f}$ and fluid pressure $p$.
5.5 Three-dimensional design of beam-insulator system
Refer to caption Refer to caption
(a) (b)
Refer to caption Refer to caption
(c) (d)
Refer to caption Refer to caption
(e) (f)
Figure 14: Convergence of critical chance $\alpha_{c}$ with the results obtained for optimal design at mean $\tilde{m}=0$ for steps in adaptive optimization discussed in Algorithm 3 from step 1 to step 5 shown in (b)-(f). The uncertain parameter $m$ has a correlation length of $L_{c}=0.25$ and a variance of $\sigma^{2}={0.1}^{2}$,the critical chance is $\alpha_{c}=0.1$, and the critical stress considered is $T_{cr}=15\>\text{MPa}$.
We demonstrate the efficiency and scalability of the proposed optimization under uncertainty framework by applying it to a three-dimensional design problem. Specifically, we consider the 3D thermal break scenario shown in Figure 1, with a thickness of 0.5. In this case, a uniform circular load is also applied to the middle of the top surface, simulating the influence of additional building envelope assemblies. The uncertain parameter $m$ is modeled with a horizontal correlation length of $L_{c}=0.25$ and a variance of $\sigma^{2}={0.1}^{2}$. For the chance constraint, we set a critical chance of $\alpha_{c}=0.1$ and a critical stress of $T_{cr}=15\>\text{MPa}$. The finite element discretization of the 3D domain resulted in 389,044 design parameters.
Refer to caption Refer to caption
(a) (b)
Figure 15: Final design obtained after five steps of the continuation scheme, achieving convergence of the critical chance $\alpha_{c}$. The 3D domain’s finite element discretization, resulted in 389044 design parameters.
Figure 14 illustrates the optimal designs obtained at each step of the continuation scheme outlined in Algorithm 3. Regions shown in red correspond to material with low porosity ($\phi_{f}=0.1$). As the continuation scheme progresses, the red regions gradually decrease, reducing mechanical support in the domain and increasing the p-norm of von Mises stress. Consequently, the probability in (14) approaches the critical chance threshold of $\alpha_{c}=0.1$. By step 5, the final design, shown in Figure 15, emerges, featuring a column-like structure of low-porosity material $\phi_{f}=0.1$, providing support aligned with the direction of the applied external load.
6 Conclusions
This study introduces an efficient and scalable computational framework for chance-constrained optimal design under high-dimensional uncertainty governed by partial differential equations (PDEs). The framework optimizes the spatial distribution of a spatially correlated and uncertain porosity field in additively manufactured aerogel thermal breaks to maximize thermal insulation performance while mitigating stress concentrations. Two key strategies are implemented within the framework. The first utilizes a quadratic Taylor expansion of the design objective with respect to the uncertain parameters to efficiently estimate the mean and variance in the cost function, as well as the chance constraints. The second employs the quadratic approximation as a control variate to enhance the efficiency of Monte Carlo estimations. The framework incorporates a randomized algorithm to solve the generalized eigenvalue problem, enabling efficient computation of the trace of the covariance-preconditioned Hessian. Scalability with respect of high-dimensional uncertain parameter is achieved through the rapid decay of eigenvalues and the dimension-independent convergence of the optimization iterations.
In future work, we aim to enhance the current design framework under uncertainty by incorporating uncertainties in the forward model parameters. These uncertainties can be rigorously quantified using Bayesian inference methods informed by experimental measurements of aerogel properties. Additionally, our present results demonstrate effective control over interface thickness between materials using Tikhonov regularization. To further improve control for a broader range of design scenarios, we plan to integrate phase-field regularization, as proposed in [31]. This enhancement would be implemented through a continuation scheme, which, while potentially increasing computational demands, offers improved precision in interface control. Finally, despite the efficiency of the existing framework, its application to large-scale engineering systems remains computationally challenging due to the extensive requirement for repeated PDE solves within the optimization loop. To overcome this limitation, we propose incorporating a Bayesian neural operator surrogate model for the multiphase PDE system, following the construction methodology outlined in [58]. This surrogate model would ensure reliable predictions by accurately capturing the relationship between joint uncertainties and design parameters. Furthermore, it would enable efficient approximation of both the PDE state and its sensitivities with respect to the design parameters, facilitating scalable and computationally feasible gradient-based optimization for complex systems.
This study addresses the computational challenges of PDE-constrained optimization under high-dimensional uncertainty by leveraging intrinsic dimensionality and incorporating chance constraints to prevent stress concentrations. The framework enhances optimal design reliability, offering a robust and scalable approach while paving the way for future advancements to refine and extend its capabilities.
Acknowledgments
The authors gratefully acknowledge the financial support received from the U.S. National Science Foundation (NSF) CAREER Award CMMI-2143662. The authors express their appreciation for the helpful discussions with Dr. Peng Chen, from Georgia Institute of Technology, regarding the problem formulation and software libraries. Additionally, the authors would like to acknowledge the support provided by the Center for Computational Research at the University at Buffalo.
Appendix A Adaptive optimization
Input : Initial design parameter $d_{0}$, smoothing parameter $\omega_{0}$, penalty parameter $\gamma_{0}$, power parameters $\sigma_{\omega}$ and $\sigma_{\gamma}$, $\hat{f}=f_{\text{QUAD}}(d)$
Output : Design gradient $d_{opt}$
1
2while $||d_{k}-d_{k-1}||\leq\mathcal{\epsilon_{\text{out}}}$ or $k<k_{\text{max}}$; $d_{k+1}=\text{Inexact Newton Conjugate-Gradient}(d_{k},\mathcal{J}(d),\nabla_{d}\mathcal{J}(d),\epsilon_{\text{in}})$.; Evaluate approximate chance $\hat{f}_{k+1}$ at $d_{k+1}$.; Update $\omega_{k+1}=\sigma_{\omega}\omega_{k}$, $\gamma_{k+1}=\sigma_{\gamma}\gamma_{k}$;
3end while
4Obtain the design gradient $d_{opt}$;
Algorithm 3 Adaptive Optimization
The gradient-based scheme uses a continuation scheme that gradually increases the smoothing parameter $\omega$ and penalty parameter $\gamma$ to achieve convergence [25]. An initial value of the smoothing parameter $\omega_{0}$ and the penalty parameter $\gamma_{0}$ is defined for the continuation scheme. These parameters are updated at the end of each iteration using scaling parameters $\sigma_{\omega}$ and $\sigma_{\gamma}$ for smoothing and penalty parameters, respectively. For $k^{th}$ iteration , the updated smoothing parameter is $\omega_{k+1}=\sigma_{\omega}\omega_{k}$ and the updated penalty parameter is $\gamma_{k+1}=\sigma_{\gamma}\gamma_{k}$. In an outer loop, we update the parameters $\omega$ and $\gamma$, whereas an Inexact Newton Conjugate-Gradient (INCG) is applied in the inner loop to solve the optimization problem (51). The termination criterion for the outer loop is if the maximum number of iterations $k_{\text{max}}$ is reached. For each iteration, the approximated chance is computed through the quadratic approximation $f_{\text{QUAD}}$ of the chance function as given in (14). The output is the optimal value of design parameter $d_{\text{opt}}$. The algorithm for this adaptive optimization is given in Algorithm 1.

### v2

@@
-Chance Constrained PDE-Constrained Optimal Design Strategies Under High-Dimensional Uncertainty
+Chance-Constrained Optimal Design of Porous Thermal Insulation Systems Under Spatially Correlated Uncertainty
 Abstract
-This study presents an advanced computational framework for the optimal design of thermal insulation components in buildings, utilizing silica aerogel porous materials.
-The framework aims to achieve superior thermal insulation while maintaining structural integrity of the component under stress concentrations.
-A multiphase continuum model is employed to simulate the thermomechanical behavior of the insulation system, governed by a set of coupled partial differential equations (PDEs).
-The design process explicitly accounts for spatially correlated uncertainties in the design parameters, particularly the spatially varying aerogel porosity, resulting in a high-dimensional, PDE-constrained optimization under uncertainty.
-The optimization problem is formulated using a risk-averse cost functional to balance insulation performance with uncertainty mitigation, incorporating statistical moments of the design objective.
-Additionally, chance constraints are imposed to limit the probability of stress exceeding critical thresholds.
-To address the challenges arising from high-dimensional parameters, the optimization leverages a second-order Taylor expansion of both the design objective and the chance constraint functions, combined with a low-rank approximation of the Hessian matrix for efficient evaluation of the generalized eigenvalue problem.
-This approach supports scalable computations, either directly or as a variance-reduction control variate for Monte Carlo estimations.
-Combined with a gradient-based optimization approach, we achieve a scalable solution algorithm with dimension-independent computational costs in terms of number of PDE solved.
-Two- and three-dimensional numerical experiments on the design of thermal breaks in buildings showcase the features of the proposed design under uncertainty framework with respect to accuracy, scalability, and computational cost.
+This paper presents a computationally efficient method for the optimal design of silica aerogel porous material systems, balancing thermal insulation performance with mechanical stability under stress concentrations.
+The proposed approach explicitly accounts for additive manufacturing uncertainties by modeling material porosity as a spatially correlated stochastic field within a multiphase finite element formulation.
+A risk-averse objective function, incorporating statistical moments of the design objective, is employed in conjunction with chance constraints that enforce mechanical stability by restricting the probability of exceeding critical stress thresholds.
+To mitigate the prohibitively high computational cost associated with the large-dimensional uncertainty space and Monte Carlo estimations of the objective function’s statistical moments, a second-order Taylor expansion is utilized as a control variate.
+Furthermore, a continuation-based smoothing strategy is introduced to address the non-differentiability of the chance constraints, ensuring compatibility with gradient-based optimization.
+The resulting framework achieves computational scalability, remaining agnostic to the dimensionality of the stochastic design space.
+The effectiveness of the method is demonstrated through numerical experiments on two- and three-dimensional thermal break systems for building insulation.
+The results highlight the framework’s capability to solve large-scale, chance-constrained optimal design problems governed by finite element models with uncertain design parameter spaces reaching dimensions in the hundreds of thousands.
 keywords:
-Thermal insulation , Optimal design , PDE-constrained optimization , chance constraint , high-dimensional parameters , uncertainty quantification
-journal: International Journal for Numerical Methods in Engineering –
+Thermal insulation , Optimal design , chance constraint optimization , high-dimensional parameters , uncertainty quantification
+journal: Journal of Risk and Uncertainty in Engineering Systems –
 […]
@@
 1 Introduction
-Silica aerogels have emerged as a premier class of porous materials for high-performance thermal insulation, widely utilized across diverse engineering applications due to their exceptionally low thermal conductivity [1, 2, 3].
-Among these applications, aerogel-based insulation components for building envelopes hold significant promise, offering the potential to enhance energy efficiency, reduce carbon emissions, and promote environmental sustainability.
-However, the adoption of silica aerogels in such systems is hindered by their inherently low mechanical strength and susceptibility to failure under localized stress concentrations.
-Recent advancements in extrusion-based additive manufacturing may be able to address these limitations by enabling precise spatial control over material deposition, thereby enhancing the mechanical stability of components.
-These developments have paved the way for integrating silica aerogels into lightweight thermal break insulation elements [2, 4], which effectively interrupt heat transfer pathways and mitigate the adverse effects of thermal bridging at assembly interfaces [5, 6].
-Nevertheless, significant uncertainties remain regarding the thermal and mechanical performance of additively manufactured aerogel components, posing challenges to their adoption as reliable and sustainable solutions for thermal break applications in building envelopes.
-In recent years, significant efforts have been dedicated to developing computational methods for topology optimization governed by partial differential equations (PDEs) (e.g., [7, 8, 9, 10, 11, 12, 13, 14, 15]).
-However, the optimal design of additively manufactured aerogel thermal breaks poses two major computational challenges.
-The first challenge arises from uncertainties in design parameters, stemming from variations in material properties, external loads, and environmental conditions.
-In particular, the layer-by-layer manufacturing process of aerogel components introduces spatially correlated uncertainties along the layer deposition direction, necessitating the representation of these uncertainties as spatially varying random fields.
-When discretized, the resulting high-dimensional parameter space makes the optimization problem computationally prohibitive due to the “curse of dimensionality” and the need for repeated evaluations of nonlinear PDEs governing material behavior.
-The second challenge involves incorporating stress constraints into the optimization process, which introduces nonlinearity and numerical instability.
-Stress gradients often exhibit discontinuities, further complicating the numerical implementation.
-To mitigate these issues, global stress constraints that aggregate stress measures (e.g., von Mises stress) across the design domain are commonly employed, significantly reducing the number of constraints.
-However, selecting an efficient stress aggregation function with appropriate parameters is crucial to achieving computational efficiency while ensuring the detection of localized stress concentrations.
-Furthermore, stress measures are influenced by uncertainties in material properties and operating conditions, necessitating the incorporation of randomness into stress constraints during the optimization process.
-To address uncertainty in optimization, one approach is to utilize a risk-averse formulation of the cost functional, that takes into account the statistical moments of the objective, e.g., [16, 17, 18, 19, 20, 21, 22, 11, 23, 24].
-However, these methods face significant computational challenges when evaluating the statistical properties of objectives or conditional probabilities, especially in the context of PDE-constrained optimization within high-dimensional parameter spaces.
-Various approximations have been proposed to address these challenges.
-Ghattas et al.
-[25, 20, 19, 18] utilized Taylor expansions for scalable, dimension-independent solutions with significant computational savings.
-This linearization, previously applied in robust design [26, 27] and topology optimization [28], was extended by Kriegesmann et al.
-[21, 29, 30] with a first-order second-moment method for efficient variance gradient evaluation.
-Tan et al.
-[31] extended the Taylor approximation method to a multi-objective PDE-constrained optimization framework for multi-material systems under uncertainty.
-Their approach incorporates phase-field regularization to effectively manage spatial sparsity and the material interface.
-Chen and Royset [32] provided further analysis of approximation-based algorithms for PDE-constrained optimization under uncertainty.
-In the context of stress constraints, several studies have investigated the integration of these constraints into optimization frameworks for diverse materials and manufacturing processes [33, 34, 35, 36, 37, 14, 38, 39, 40, 41].
-Significant efforts have also focused on incorporating uncertainty into stress-constrained optimization [42], addressing factors such as variable loading and heat sources.
-For instance, Silva et al.
-[35] developed a method to account for uncertainties in boundary conditions during manufacturing.
-Similarly, Oh et al.
-[43] proposed a topology optimization approach that incorporates uncertainties in loading positions.
-Building on previous work, this study presents a scalable computational framework for risk-averse PDE-constrained optimization under spatially correlated and high-dimensional uncertainty, while incorporating chance constraints.
-The framework is applied to the design of thermal breaks in building envelope, aiming to maximize thermal insulation performance and ensure mechanical stability by mitigating stress concentrations through chance constraints.
-The forward model is based on a thermomechanical continuum mixture theory for porous materials.
-The PDE model parameters, such as solid and fluid conductivities and solid elastic modulus, have been determined through model calibration in prior studies.
-Material porosity, modeled as a spatially correlated random field, serves as the design parameter.
-Discretization of this field transforms the problem into a high-dimensional optimization under uncertainty To achieve target performance levels while mitigating the uncertainty, the framework employs a risk-averse formulation that integrates the mean and variance of the design objective.
-The computational burden of Monte Carlo moment estimation is mitigated using second-order Taylor approximations for the design objective and chance constraints.
-An inexact Newton-Conjugate Gradient algorithm with convergence properties independent of problem dimension is utilized to ensure computational scalability—measured by the number of PDE solutions—regardless of the dimensionality of the design parameter space.
-The effectiveness and computational cost of the framework are explored through two- and three-dimensional numerical examples, including a case study of a beam-thermal break component in the building envelope.
-Compared to our earlier work in [31], the key contributions of this study are as follows: (1) Incorporation of probabilistic constraints in the optimization framework to mitigate stress concentrations.
-(2) Utilization of second-order Taylor approximations as a control variate for Monte Carlo estimation of the design objective, enabling more accurate optimal designs in high-variance uncertainty scenarios.
-(3) Implementation of an inexact Newton-Conjugate Gradient algorithm, facilitating parallelization and accelerating PDE-constrained optimization calculations.
-(4) Extension of the numerical framework to three-dimensional design of insulation components.
-The rest of the paper is organized as follows: Section 2 presents the optimal design under uncertainty problem, including the thermomechanical forward model, the modeling of spatially correlated uncertainty in the design parameter, and the formulation of mean-variance optimization.
-Section 3 is dedicated to the approximation methods, encompassing Taylor approximation and the Monte Carlo estimate with control variate of the moments of the design objective.
-Section 4 describes the gradient-based optimization method and explains the computation of the gradient of the cost functional concerning the design parameter.
-Section 5 presents the numerical experiments on the optimal design of thermal breaks, followed by the conclusions and future works in Section 6.
+The development of high-performance insulation materials is critical for improving energy efficiency and reducing the environmental impact of buildings.
+Silica aerogels, renowned for their ultra-low thermal conductivity [1, 2, 3], offer exceptional insulating properties but are highly susceptible to mechanical failure due to stress-concentration.
+Advances in additive manufacturing have facilitated precise control over the spatial distribution of material properties, potentially enabling mechanically robust and lightweight insulation systems.
+However, the inherent uncertainties in the thermal and mechanical performance of these engineered materials pose significant challenges to their reliable integration into construction applications, necessitating robust design strategies.
+While finite element-based topology optimization techniques have advanced considerably [4, 5, 6, 7, 8, 9, 10, 11, 12], design of additively manufactured aerogel components presents two primary challenges.
+The first challenge arises from uncertainties in spatially varying design parameters, particularly those induced by the layer-by-layer fabrication process.
+This manufacturing variability introduces spatially correlated stochastic fluctuations along the deposition direction, leading to a high-dimensional design space when discretized using finite element methods.
+Consequently, the computational cost becomes prohibitive due to the necessity of solving multiple nonlinear finite element solution of the porous material response.
+The second challenge involves incorporating stress constraints into the optimization framework.
+These constraints introduce significant nonlinearity and discontinuities, particularly when addressing localized stress concentrations.
+Global stress constraints, often formulated using aggregated measures such as von Mises stress, require carefully chosen aggregation functions to maintain both computational efficiency and sensitivity to critical stress regions.
+Furthermore, the presence of uncertainty necessitates a stochastic formulation of stress constraints, integrating probabilistic constrained into the optimal design formulation.
+To address the first challenge and ensure robustness in optimal design under uncertainty, risk-averse optimization methods incorporates statistical moments, such as mean and variance, into the objective function [13, 14, 15, 16, 17, 8, 18, 19, 20].
+However, directly integrating these moments into finite element-based optimization with high-dimensional parameters imposes significant computational costs.
+Various approximation techniques have been developed to alleviate this burden.
+Notably, Ghattas et al.
+[21, 14, 13] demonstrated that Taylor series expansions of statistical moments provide scalable and computationally efficient solutions for optimization problems governed by finite element models.
+Kriegesmann et al.
+[16, 22, 23] introduced first-order second-moment methods for efficient gradient-based variance estimation.
+More recently, Tan et al.
+[24] extended Taylor-based approximations to multi-objective optimization, incorporating phase-field regularization to enhance spatial sparsity control.
+To address the second challenge, several studies have explored the integration of stress constraints into optimization frameworks across various materials and manufacturing processes [25, 26, 27, 28, 29, 11, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39].
+Efforts have also focused on incorporating uncertainty into stress-constrained optimization [40], addressing factors such as variable loading and heat sources.
+For example, Silva et al.
+[27] accounted for uncertainties in boundary conditions during manufacturing, while Oh et al.
+[41] incorporated uncertainties in loading positions in topology optimization.
+Mart’ınez and Ortigosa [42] developed a risk-averse approach for fail-safe structures, treating collapse as an uncertainty source rather than a deterministic event.
+Liu et al.
+[43] introduced interval reliability-based topology optimization, considering multi-source uncertainties in dynamic responses.
+We present a novel computational framework for the optimal design of aerogel-based insulation systems under spatially varying stochastic design parameters, incorporating chance constraints to ensure mechanical stability.
+The thermomechanical property of the porous silica aerogels is characterized by a multiphase finite element model, where material porosity is represented as a spatially correlated stochastic field.
+The objective is to optimize the spatial distribution of material porosity to maximize thermal insulation while mitigating stress concentrations of the insulation component.
+To alleviate the computational burden associated with Monte Carlo-based moment estimation in the mean-variance cost functional, we employ second-order Taylor approximations to efficiently approximate both the design objective and chance constraints.
+Building on our previous work in [24], this study introduces the following key contributions: (1) Explicit integration of probabilistic stress constraints to directly regulate stress concentrations within the optimization framework, ensuring mechanical robustness.
+(2) Rather than directly employing the Taylor expansion of the moments of the design objective, this approximation is leveraged as a control variate in Monte Carlo estimation, substantially enhancing accuracy in high-variance uncertainty scenarios.
+(3) Implementation of an inexact Newton-Conjugate Gradient algorithm, enabling scalable optimization for large-scale problems while maintaining convergence guarantees.
+The effectiveness and computational efficiency of the proposed approach are demonstrated through two- and three-dimensional numerical experiments of a beam-thermal break component within a building envelope.
+The remainder of the paper involves presenting the thermomechanical model for silica aerogels, incorporating spatially correlated uncertainties and the mean-variance cost functional for optimal design in Section 2.
+Section 3 presents the Taylor approximations and Monte Carlo estimators with control variates.
+Section 4 details the integration of chance constraints within the gradient-based optimization framework.
+Numerical experiments demonstrating the method’s effectiveness are provided in Section 5, followed by conclusions and directions for future research in Section 6.
 2 Problem formulation: Design of aerogel thermal break under uncertainty
-This study investigates the optimal design of building insulation components using silica aerogel materials.
-Figure 1 illustrates the schematic representation of a building thermal break, consisting of a concrete beam integrated with an aerogel insulation system, along with the thermal and mechanical boundary conditions employed in the numerical experiments.
-Concrete beams significantly influence the thermal performance of building envelopes due to their high thermal conductivity, which creates preferential pathways for heat transfer.
-The inclusion of silica aerogel thermal breaks disrupts this conduction, effectively mitigating thermal bridging from the interior to the exterior of the building envelope.
-However, thermal expansion of the concrete beam can induce stress concentrations within the silica aerogel insulation component, potentially leading to mechanical failure and compromising the thermal break’s effectiveness.
-In this section, we introduce the problem of optimal design under uncertainty, leveraging a forward model of aerogel and concrete materials governed by a system of partial differential equations (PDEs).
-We then describe how the uncertainty in design parameters are modeled, as well as formulation of design problem.
+Figure 1 illustrates a building insulation system involving a concrete beam with an integrated silica aerogel thermal break.
+Concrete’s high thermal conductivity promotes heat transfer through building envelopes, forming thermal bridges, while thermal breaks mitigate this effect.
+However, thermal expansion of the concrete beam can generate stress concentrations in the aerogel, risking mechanical failure and reducing insulation effectiveness.
+This section formulates the optimal design problem under uncertainty, incorporating thermomechanical models governed by partial differential equations (PDEs) and detailing the modeling of uncertainty in design parameters.
 Figure 1: The domain of a beam-insulator system used for numerical experiments, indicating both the mechanical (left) and thermal (right) boundary conditions.
-2.1 Forward thermomechanical model and design objectives
-For silica aerogel materials, we utilize a thermomechanical model based on continuum mixture theory [44].
-The model accounts for both an incompressible solid aerogel phase and a compressible fluid (gaseous) phase within the material.
-According to the mixture theory, each phase independently adheres to its own motion, balance laws, internal energy, and entropy, while collectively following the overall balance laws governing the entire mixture.
-As a result, a system of four partial differential equations is derived, which governs the heat transfer in the solid and fluid phases,and the deformation in the solid phase and the evolution of pore pressure.
-For our problem of designing thermal breaks, we focus on steady-state and static conditions, which simplify the governing equations to the following PDEs,
-$\begin{cases}-\nabla\cdot(\phi_{s}\,\kappa_{s}\,\nabla\theta_{s})+h(\theta_{s}-\theta_{f})=0\quad&\text{in}\quad\Omega,\\ -\nabla\cdot(\phi_{f}\,\kappa_{f}\,\nabla\theta_{f})-h(\theta_{s}-\theta_{f})=0\quad&\text{in}\quad\Omega,\\ -\phi_{s}\,\kappa_{s}\,\nabla\theta_{s}=h_{\mathrm{air}}(\theta_{s}-\theta_{\mathrm{amb}})\quad&\text{on}\quad\Gamma_{1}\cup\Gamma_{4},\\ -\phi_{f}\,\kappa_{f}\,\nabla\theta_{f}=h_{\mathrm{air}}(\theta_{f}-\theta_{\mathrm{amb}})\quad&\text{on}\quad\Gamma_{1}\cup\Gamma_{4},\\ \end{cases}$ (1)
-$\begin{cases}D\,p+(\nabla\cdot\mathbf{u}_{s})=0\quad&\text{in}\quad\Omega,\\ -\nabla\cdot\mathbf{T}^{\prime}_{s}-(2\,\phi_{f}-1)\,\nabla p=0\quad&\text{in}\quad\Omega,\\ \mathbf{u}_{s}=\bar{\mathbf{u}}_{s}\quad&\text{on}\quad\Gamma_{1}\cup\Gamma_{3},\\ \nabla\mathbf{u}_{s}.\mathbf{n}=0\quad&\text{on}\quad\Gamma_{2}\end{cases}$ (2)
-where model parameters are $\theta=(\kappa_{s},\kappa_{f},D,\mathbf{E}_{s})$ and states are $u=(\theta_{s},\theta_{f},\mathbf{u_{s}},p)$.
-The state variables $\theta_{s}$, $\theta_{f}$, $\mathbf{u}_{s}$, and $p$ represent the solid and fluid temperatures, solid displacement, and fluid pressure, respectively, defined over the domain $\Omega$.
-Here, $\phi_{f}(\mathbf{x})$ represents the volume fraction of fluid (porosity), and $\phi_{s}(\mathbf{x})$ represents the volume fraction of solid, which can alternatively be written as $1-\phi_{f}(\mathbf{x})$.
-The solid displacement is subject to fixed values $\bar{\mathbf{u}}_{s}$ at the boundaries $\Gamma_{1}$ and $\Gamma_{3}$.
-A roller boundary condition is implemented at $\Gamma_{2}$.
-At $\Gamma_{1}$ and $\Gamma_{4}$, we consider convective heat transfer with a coefficient $h_{\mathrm{air}}$ and ambient temperature $\theta_{\mathrm{amb}}$.
-A Neumann type boundary condition with no heat flux is implemented at $\Gamma_{2}$.
-In the above equations, $\Omega\subset\mathbb{R}^{d}$ is an open and bounded domain with Lipschitz boundary $\partial\Omega$, and $\Gamma_{1},\Gamma_{3},\Gamma\subset\partial\Omega$ denote the Dirichlet and $\Gamma_{2}\subset\partial\Omega$ indicate the Neumann portions of $\partial\Omega$ such that $\Gamma_{1}\cup\Gamma_{2}\cup\Gamma_{4}=\partial\Omega$, $\Gamma_{1}\cap\Gamma_{2}=\emptyset$, $\Gamma_{3}\cap\Gamma_{4}=\Gamma_{3}$, and $\Gamma_{1}\cap\Gamma_{4}=\emptyset$.
-In (2), the effective solid stress is
-$\mathbf{T}^{\prime}_{s}=2\,\mu\,\mathbf{E}_{s}+\lambda\,\mathrm{tr}(\mathbf{E}_{s})\,\mathbf{I}\quad\text{with}\quad\mathbf{E}_{s}=\frac{1}{2}\big{(}\nabla\mathbf{u}_{s}+(\nabla\mathbf{u}_{s})^{T}\big{)},$ (3)
-where $\mathbf{E}_{s}$ is the solid strain and $\lambda$ and $\mu$ are the Lamé constants.
-Determining the model parameters $\kappa_{s}$, $\kappa_{f}$, $h$, $K$, $C$, and $\mu$ from experimental data of silica aerogel is demonstrated in [45] and assumed to be known.
-for the concrete beam we consider the following thermoelastic model,
-$\begin{cases}-\nabla\cdot(\,\kappa_{b}\,\nabla\theta_{b})=0\quad&\text{in}\quad\Omega_{b},\\ \kappa_{b}\,\nabla\theta_{b}=h_{\mathrm{air}}(\theta_{b}-\theta_{\mathrm{amb}})\quad&\text{on}\quad\Gamma_{3},\\ \nabla.\mathbf{T}_{b}=0&\text{in}\quad\Omega_{b}\\ \end{cases}$ (4)
-where $\kappa_{b}$ is thermal conductivity of beam, $\theta_{b}$ represents temperature of beam and $\mathbf{T}_{b}$ is the beam stress. In (4), the stress in beam is
-$\mathbf{T}_{b}=2\mu_{b}\mathbf{E}_{b}+\lambda_{b}\mathrm{tr}(\mathbf{E}_{b})\mathbf{I}-\alpha_{T}(3\lambda_{b}+2\mu_{b})(\theta_{b}-\theta_{0})\quad\text{with}\quad\mathbf{E}_{b}=\frac{1}{2}\big{(}\nabla\mathbf{u}_{b}+(\nabla\mathbf{u}_{b})^{T}\big{)}$ (5)
-where $\mathbf{E}_{b}$ is strain in the beam, and $\lambda_{b}$ and $\mu_{b}$ are the Lamé constants for the beam, $\alpha_{T}$ is coefficient of thermal expansion and $\theta_{0}$ is the reference temperature.
-In the following sections, the forward state problems (1), (2) and (4) are expressed in an abstract form as,
-$\mathcal{R}(\mathbf{u},m,d)=0\quad\mathrm{in\ }\mathcal{V}^{\prime},$ (6)
-where, given the uncertain parameter $m\in\mathcal{M}$ and design parameter $d\in\mathcal{D}$, the goal is to find the state $\mathbf{u}=(\theta_{s},\theta_{f},\mathbf{u}_{s},p,\mathbf{u}_{b})\in\mathcal{U}$ such that the PDE residual operator $\mathcal{R}(\cdot)$ is minimized to zero in $\mathcal{V}^{\prime}$.
-Here $\mathcal{M}$ and $\mathcal{D}$ represent infinite-dimensional Hilbert spaces, and $\mathcal{V}^{\prime}$ denotes the corresponding dual space.
-Weak formulation
-To derive the weak formulation of the governing equations, we start with the partial differential equations (PDEs) (1) ,(2) and (4).
-The problem can be stated as follows: Find $\mathbf{u}=(\theta_{s},\theta_{f},\theta_{b},\mathbf{u}_{s},\mathbf{u}_{b},p)\in\mathcal{U}$, such that
-$-\big{<}\phi_{s}\kappa_{s}\nabla\theta_{s},\nabla z_{s}\big{>}-\big{<}h(\theta_{s}-\theta_{f}),z_{s}\big{>}-\big{<}\phi_{s}h_{\mathrm{air}}(\theta_{s}-\theta_{\mathrm{amb}}),z_{s}\big{>}$ $=$ $0,$
-$-\big{<}\phi_{f}\kappa_{f}\nabla\theta_{f},\nabla z_{f}\big{>}+\big{<}h(\theta_{f}-\theta_{f}),z_{f}\big{>}-\big{<}\phi_{f}h_{\mathrm{air}}(\theta_{f}-\theta_{\mathrm{amb}}),z_{f}\big{>}$ $=$ $0,$
-$-\big{<}\kappa_{b}\nabla\theta_{b},\nabla z_{b}\big{>}-\big{<}h_{\mathrm{air}}(\theta_{b}-\theta_{\mathrm{amb}}),z_{b}\big{>}$ $=$ $0,$
+The solid displacement are fixed at the boundaries $\Gamma_{1}$ and $\Gamma_{3}$ and roller boundary conditions are implemented at $\Gamma_{2}$.
+At $\Gamma_{1}$ and $\Gamma_{4}$, we consider convective heat transfer cold ambient temperature in the building exterior and hot ambient temperature in the building interior.
+no heat flux boundary is implemented at $\Gamma_{2}$.
+2.1 Thermomechanical finite element models and design objectives
+For silica aerogel materials, we adopt a multiphase model based on continuum mixture theory [44], capturing the interactions between an incompressible solid aerogel phase and a compressible fluid (gaseous) phase.
+The governing equations consist of a system of PDEs that describe heat transfer in both phases, solid-phase deformation, and pore pressure evolution, over the domain $\Omega$ and boundaries $\Gamma_{1}$, $\Gamma_{2}$, $\Gamma_{3}$, and $\Gamma_{4}$, as
+$\begin{cases}-\nabla\cdot(\phi_{s}\,\kappa_{s}\,\nabla\theta_{s})+h(\theta_{s}-\theta_{f})=0\quad&\text{in}\quad\Omega,\\ -\nabla\cdot(\phi_{f}\,\kappa_{f}\,\nabla\theta_{f})-h(\theta_{s}-\theta_{f})=0\quad&\text{in}\quad\Omega,\\ -\phi_{s}\,\kappa_{s}\,\nabla\theta_{s}=h_{\mathrm{conv}}(\theta_{s}-\theta_{\mathrm{a}})\quad&\text{on}\quad\Gamma_{1}\cup\Gamma_{4},\\ -\phi_{f}\,\kappa_{f}\,\nabla\theta_{f}=h_{\mathrm{conv}}(\theta_{f}-\theta_{\mathrm{a}})\quad&\text{on}\quad\Gamma_{1}\cup\Gamma_{4},\\ \end{cases}$ (1)
+$\begin{cases}D\,p+(\nabla\cdot\mathbf{u}_{s})=0\quad&\text{in}\quad\Omega,\\ -\nabla\cdot\mathbf{T}^{\prime}_{s}-(2\,\phi_{f}-1)\,\nabla p=0\quad&\text{in}\quad\Omega,\\ \mathbf{u}_{s}=\mathbf{0}\quad&\text{on}\quad\Gamma_{1}\cup\Gamma_{3},\\ \nabla\mathbf{u}_{s}\cdot\mathbf{n}=0\quad&\text{on}\quad\Gamma_{2}\end{cases}$ (2)
+where $\nabla$ is gradient operator, $\mathbf{n}$ is the outward unit normal vector, and the effective solid stress is $\mathbf{T}^{\prime}_{s}=2\,\mu\,\mathbf{E}_{s}+\lambda\,\mathrm{tr}(\mathbf{E}_{s})\,\mathbf{I}$ with $\mathbf{E}_{s}(\mathbf{u}_{s})$ being the solid strain.
+The model parameters are solid conductivity $\kappa_{s}$, fluid conductivity $\kappa_{f}$, fluid compressibility $D$, the Lamé constants $\lambda$ and $\mu$, and convective heat coefficient $h_{\mathrm{conv}}$ with respect to hot or cold ambient temperature $\theta_{\mathrm{a}}$, and are assumed to be known for silica aerogel from previous studies.
+The state variables are $\theta_{s}$, $\theta_{f}$, $\mathbf{u}_{s}$, and $p$ represent the solid and fluid temperatures, solid displacement, and fluid pressure, respectively.
+Additionally, $\phi_{s}$ and $\phi_{f}$ are the volume fractions of solid and fluid (porosity) phases.
+The concrete beam is assumed to follow a thermoelastic model,
+$\begin{cases}-\nabla\cdot(\,\kappa_{b}\,\nabla\theta_{b})=0\quad&\text{in}\quad\Omega_{b},\\ -\kappa_{b}\,\nabla\theta_{b}=h_{\mathrm{conv}}(\theta_{b}-\theta_{\mathrm{a}})\quad&\text{on}\quad\Gamma_{3},\\ \nabla\cdot\mathbf{T}_{b}=0&\text{in}\quad\Omega_{b}\\ \end{cases}$ (3)
+where $\theta_{b}$ represents temperature of beam and Cauchy stress is $\mathbf{T}_{b}=2\mu_{b}\mathbf{E}_{b}+\lambda_{b}\mathrm{tr}(\mathbf{E}_{b})\mathbf{I}-\alpha_{T}(3\lambda_{b}+2\mu_{b})(\theta_{b}-\theta_{0})$ with $\mathbf{E}_{b}$ is strain in the beam.
+In (3), $\kappa_{b}$ is thermal conductivity of beam, $\lambda_{b}$ and $\mu_{b}$ are the Lamé constants for the beam, $\alpha_{T}$ is coefficient of thermal expansion and $\theta_{0}$ is the reference temperature.
+Finite element solution of the above model requires the weak formulation of the governing equations as:
+Find $\mathbf{u}=(\theta_{s},\theta_{f},\theta_{b},\mathbf{u}_{s},\mathbf{u}_{b},p)$, such that
+$-\big{<}\phi_{s}\kappa_{s}\nabla\theta_{s},\nabla z_{s}\big{>}-\big{<}h(\theta_{s}-\theta_{f}),z_{s}\big{>}-\big{<}\phi_{s}h_{\mathrm{conv}}(\theta_{s}-\theta_{\mathrm{a}}),z_{s}\big{>}$ $=$ $0,$
+$-\big{<}\phi_{f}\kappa_{f}\nabla\theta_{f},\nabla z_{f}\big{>}+\big{<}h(\theta_{f}-\theta_{f}),z_{f}\big{>}-\big{<}\phi_{f}h_{\mathrm{conv}}(\theta_{f}-\theta_{\mathrm{a}}),z_{f}\big{>}$ $=$ $0,$
+$-\big{<}\kappa_{b}\nabla\theta_{b},\nabla z_{b}\big{>}-\big{<}h_{\mathrm{conv}}(\theta_{b}-\theta_{\mathrm{a}}),z_{b}\big{>}$ $=$ $0,$
 $-\big{<}(\lambda+\frac{1-2\phi_{f}}{D})(\nabla\cdot\mathbf{u}_{s})\mathbf{I}+\mu\big{(}\nabla\mathbf{u}_{s}+(\nabla\mathbf{u}_{s})^{T}\big{)},\boldsymbol{w}_{u}\big{>}+\big{<}\mathbf{t},\boldsymbol{w}_{u}\big{>}$ $=$ $0,$
-$-\big{<}(\lambda_{b})(\nabla\cdot\mathbf{u}_{b})\mathbf{I}+\mu_{b}\big{(}\nabla\mathbf{u}_{b}+(\nabla\mathbf{u}_{b})^{T}\big{)},\boldsymbol{w}_{b}\big{>}+\big{<}\mathbf{t},\boldsymbol{w}_{b}\big{>}$ $=$ $0,,$
+$-\big{<}(\lambda_{b})(\nabla\cdot\mathbf{u}_{b})\mathbf{I}+\mu_{b}\big{(}\nabla\mathbf{u}_{b}+(\nabla\mathbf{u}_{b})^{T}\big{)},\boldsymbol{w}_{b}\big{>}+\big{<}\mathbf{t},\boldsymbol{w}_{b}\big{>}$ $=$ $0,$ (4)
 […]
-$r(\mathbf{u},\mathbf{v},m,d):=\,_{\mathcal{V}}\langle\mathbf{v},\mathcal{R}(\mathbf{u},m,d)\rangle_{\mathcal{V}^{\prime}}=0\quad\forall\mathbf{v}\in\mathcal{V}$ (7)
-where ,$\,{}_{\mathcal{V}}\langle\cdot,\cdot\rangle_{\mathcal{V}^{\prime}}$ represent the duality pairing between the $\mathcal{V}$ and $\mathcal{V}^{\prime}$.
-Figure 2: Finite element solution of the forward model for the beam-insulator system with constant porosity $\phi_{f}=0.7$, depicting the solid temperature $\theta_{s}$, fluid temperature $\theta_{f}$, displacement magnitude $||\mathbf{u}_{s}||$, and von Mises stress $T_{\text{VM}}$.
+Figure 2: Finite element solution of the thermomechanical (forward) model for the beam-insulator system with constant porosity $\phi_{f}=0.7$, depicting the solid temperature $\theta_{s}$, fluid temperature $\theta_{f}$, displacement magnitude $||\mathbf{u}_{s}||$, and von Mises stress $T_{\text{VM}}$.
 Figure 2 presents the deterministic finite element solution of the forward model for the beam-insulator system under constant porosity $\phi_{f}=0.7$, utilizing the model parameters listed in Table 1.
-Thermal expansion of the thermoelastic concrete beam induces stress concentrations within the aerogel thermal break, risking mechanical failure.
-This can be mitigated by varying the aerogel’s porosity across the insulation, maintaining its effectiveness in interrupting heat transfer between the building interior and exterior.
-Design objectives
-The design objective $Q$ is defined as thermal compliance of the aerogel thermal break component over domain $\Omega$,
-$Q$ $=$ $\frac{1}{2}\sum_{i=s,f}\big{<}\phi_{i}\,\kappa_{i}\,\nabla\theta_{i},\nabla\theta_{i}\big{>}+{\big{<}\phi_{i}h_{\mathrm{air}}(\theta_{i}-\theta_{\mathrm{amb}}),\theta_{i}\big{>}},$
-The design problem focuses on determining the spatial distribution of the fluid volume fraction (porosity), $\phi_{f}(\mathbf{x})$, to minimize stress concentrations within the thermal break, thereby preventing mechanical failure while preserving its thermal insulation performance.
-2.2 Uncertain design parameter
-Due to the inherent challenges in precisely controlling the properties of additively manufactured aerogel, considerable uncertainty arises in the porosity values within the thermal break.
-To effectively model uncertainty, a design parameter as a spatial field $d\in[0,1]$ is introduced, employing a sigmoid mapping function, ensuring that porosity values asymptotically remain within the physically valid range, maintaining near-linear behavior within $\phi_{f}\in[90\%,10\%]$, and minimize the impact on the spatial correlation length, such that [31]
-$\phi_{f}\equiv\text{sigmoid}\left(d+m\right).$ (9)
-The uncertain parameter $m$ is represented as a Gaussian random field characterized by the Matérn covariance kernel.
-Specifically, $m$ is distributed as $\mathcal{N}(\bar{m},\mathcal{C})$ where $\bar{m}$ is the mean and $\mathcal{C}$ is the covariance operator.
+As shown in this figure, thermal expansion of the thermoelastic concrete beam induces stress concentrations within the aerogel thermal break.
+We define the design objective $Q$ as thermal compliance of the thermal break,
+$Q=\frac{1}{2}\sum_{i=s,f}\big{<}\phi_{i}\,\kappa_{i}\,\nabla\theta_{i},\nabla\theta_{i}\big{>}+{\big{<}\phi_{i}h_{\mathrm{conv}}(\theta_{i}-\theta_{\mathrm{a}}),\theta_{i}\big{>}},$ (5)
+The design goal is determining the spatial distribution of the fluid volume fraction (porosity), $\phi_{f}(\mathbf{x})$, to minimize stress concentrations within the thermal break, thereby preventing mechanical failure while preserving its thermal insulation performance.
+2.2 Spatially correlated uncertain design parameter
+Due to the inherent challenges in precisely controlling the properties of additively manufactured aerogel, considerable uncertainty arises in the porosity values within the thermal break arising from the layer-by-layer deposition of aerogel ink.
+To effectively model such uncertainty, first we define uncertain parameter $m$ as a Gaussian random field characterized by the Matérn covariance kernel.
+Specifically, $m$ is distributed as $\mathcal{N}(\bar{m},\mathcal{K})$ where $\bar{m}$ is the mean and $\mathcal{K}$ is the covariance operator.
 The sampling of Gaussian random field $m$ is equivalent to solving the elliptic equation with a robin boundary condition,
@@
 $m+1.42\>\nabla m.\>\boldsymbol{n}=$  0onΓ $\>0\quad\text{on}\quad\Gamma$
-where $\nabla$ is gradient operator,$\Delta$ is the Laplace operator and $\mathbf{n}$ is the outward unit normal vector.
-Here, $W$ represents the spatial White Gaussian noise with a unit variance, and the parameters $\gamma$ and $\delta$ control the variance $\sigma^{2}={1}/{4\pi\,\delta\,\gamma}$ and spatial correlation length $L_{c}=\sqrt{{8\gamma}/{\delta}}$.
-The factor ${\sqrt{\delta\,\gamma}}/{1.42}$ (Robin coefficient) is chosen to minimize boundary artifacts [46], and $\boldsymbol{\Theta}$ is a symmetric positive definite two-dimensional (2D) anisotropic tensor
-[…]
-[…]
+Here, $W$ represents the spatial White Gaussian noise with a unit variance, and the parameters $\gamma$ and $\delta$ control the variance $\sigma^{2}={1}/{4\pi\,\delta\,\gamma}$ and spatial correlation length $L_{CR}=\sqrt{{8\gamma}/{\delta}}$.
+The factor ${\sqrt{\delta\,\gamma}}/{1.42}$ is chosen to minimize boundary artifacts [45], and $\boldsymbol{\Theta}$ is an anisotropic tensor.
+Then we define the design parameter $d$ as a spatial field, employing a sigmoid mapping function $\phi_{f}\equiv\text{sigmoid}\left(d+m\right)$, ensuring that porosity values asymptotically remain within the physically valid range, maintaining near-linear behavior within $\phi_{f}\in[90\%,10\%]$ [24].
+For simplicity, for the rest of this paper, we express the forward state problems (1), (2) and (3) in an abstract form $\mathcal{R}(\mathbf{u},m,d)=0$, and the weak form (4) is expressed as $r(\mathbf{u},\mathbf{v},m,d)$, where $\mathbf{u}$ is the state variables, $\mathbf{v}$ is adjoint variables and $m$ and $d$ are uncertain and design parameters.
 2.3 Chance-constrained PDE-constrained optimal design under uncertainty
-We consider the design problem as minimizing the design objective, thermal compliance in (LABEL:eq:qois), while accounting for uncertainty in objective through its explicit dependence on the uncertain parameter $m$, i.e., $Q(d,m)$.
-Extending the mean-variance optimization, e.g., [18, 19, 31], the cost functional for the proposed optimal design takes the form,
-$\mathcal{J}(d)=\mathbb{E}[Q(d,m)]+\beta_{V}\,\mathbb{V}[Q(d,m)]+\beta_{R}\,R(d).$ (12)
-To prevent mechanical failure of the aerogel thermal break due to stress concentration, the maximum von Mises stress $T_{\text{VM}}$ must remain below a critical threshold.
+We formulate the design problem as minimizing the thermal compliance in (5), which explicitly depends on the uncertain parameter, $Q(d,m)$.
+The mean-variance cost functional, e.g., [13, 14, 24], is given by:
+$\mathcal{J}(d)=\mathbb{E}[Q(d,m)]+\beta_{V}\,\mathbb{V}[Q(d,m)]+\beta_{RG}\,R_{p}(d).$ (7)
+where $R_{p}(d)$ is the regularization term, while $\beta_{V}$ and $\beta_{RG}$ represent the coefficients assigned to the variance and the regularization component, respectively.
+To prevent mechanical failure of the aerogel thermal break due to stress concentration, the von Mises stress $T_{\text{VM}}$ must remain below a critical threshold.
 This requirement is incorporated as a probabilistic (chance) constraint in the optimization problem, as $T_{\text{VM}}$ depends on uncertain parameters.
 However, due to the non-differentiable nature of the maximum stress, direct optimization poses significant challenges.
-This challenge is addressed using stress aggregation function based on p-norm [41], which is a smooth approximation of the maximum stress and enables gradient-based optimization techniques.The p-norm for the von Mises stress can be computed as
+This challenge is addressed using stress aggregation function based on p-norm [33], which is a smooth approximation of the maximum stress and enables gradient-based optimization techniques.
+The p-norm for the von Mises stress over the domain can be computed as
 […]
 The chance constraint function can hence be written as,
-$f=T_{cr}-T_{\text{pn}},$ (14)
-where $T_{cr}$ is the limiting critical stress and $T_{\text{pn}}$ is the p-norm of the von Mises stress over the domain. To avoid stress concentration, we consider a chance constraint:
-$P(f(m,d)\geq 0)\leq\alpha_{c},$ (15)
-for a critical chance $10<\alpha_{c}<$ 1. The probability is given by
-$P(f(m,d)\geq 0)$ $=$ $\mathbb{E}[\mathbb{I}_{[0,\infty]}(f(m,d))]$ (16)
-[…]
-where $\mathbb{I}_{[0,\infty]}(f(m,d))$ is an indicator function defined as:
-$\mathbb{I}_{[0,\infty]}(f(m,d))=\begin{cases}1&\text{if }f(m,d)\geq 0\\ 0&\text{if }f(m,d)<0\\ \end{cases}$ (17)
-uncertainty as,
-$d_{\text{opt}}(\mathbf{x})=\min_{{d}}$ $\ \mathcal{J}({d})$ (18)
+$f=T_{cr}-T_{\text{pn}}.$ (9)
+In above relation $T_{cr}$ is the limiting critical stress, $\mathbf{T}_{0}$ is stress deviator tensor, and $\mathrm{tr}(\cdot)$ is the trace.
+To avoid stress concentration, we consider a chance constraint:
+$P(f(m,d)\geq 0)=\mathbb{E}[\chi_{[0,\infty)}(f(m,d))]\leq\alpha_{c},$ (10)
+for a critical chance $10<\alpha_{c}<$ 1, where $\chi_{[0,\infty)}(f(m,d))$ is an indicator function defined as:
+$\chi_{[0,\infty)}(f(m,d))=\begin{cases}1&\text{if }f(m,d)\geq 0\\ 0&\text{if }f(m,d)<0\\ \end{cases}$ (11)
+Thus, the risk-averse optimal design problem under uncertainty can be formulated as a PDE-constrained optimization problem, expressed as:
+$d_{\text{opm}}=\min_{{d}}$ $\ \mathcal{J}({d})$ (12)
 $\mathrm{s.t.\ }$ $\mathcal{R}(\boldsymbol{u},m,d)=0$
-$P(f(m,d)\geq 0)\leq\alpha_{c}\quad\mathrm{in\ }\Omega$
-Monte Carlo estimation
-A common approach to compute the mean and variance terms in (12) is the Monte Carlo estimation for $N_{mc}$ samples,
+$P(f(m,d)\geq 0)\leq\alpha_{c}$
+3 Scalable framework for chance constrained optimal design
+A conventional method for calculating the mean and variance components in (7) and the probability constrained in (10) is the Monte Carlo estimation for $N_{mc}$ samples,
 $\mathbb{E}[Q]$ $\approx\frac{1}{N_{mc}}\sum_{i=1}^{N_{mc}}Q(m^{(i)}),$ (13)
 $\mathbb{V}[Q]$ $\approx\Big{(}\frac{1}{N_{mc}}\sum_{i=1}^{N_{mc}}Q^{2}(m^{(i)})\Big{)}-\Big{(}\frac{1}{N}_{mc}\sum_{i=1}^{N_{mc}}Q(m^{(i)})\Big{)}^{2},$ (14)
-where $\{m^{(i)}\}_{i=1}^{N_{mc}}$ are independent and identically distributed (i.i.d) random samples from the uncertain and spatially correlated parameter.
-Similarly, the chance constraint function can be approximated by
-$P(f(.,d)\geq 0):=\frac{1}{M_{f}}\sum_{i=1}^{M_{f}}\mathbb{I}_{[0,\infty)]}(f(m_{i},d))$ (21)
-To achieve an accurate approximation, particularly for $f_{M}$ for $\alpha$ close to 00, a large number of samples are required due to slow convergence of the Monte Carlo method.
-The convergence rate of Monte Carlo method is $\mathcal{O}\big{(}\frac{1}{\sqrt{N_{mc}}}\big{)}$, therefore for accurate computation of the cost functional (12), a considerable number of PDE model evaluations will be required, making the optimization under uncertainty problem computationally prohibitive.
-3 Scalable framework for chance constrained optimal design
-In this section, we present a computationally tractable approach by approximating the mean and variance of the design objective by means of low-order Taylor expansion [18, 19].
-The quadratic Taylor approximation of $Q$ at the mean of the uncertain parameter $\bar{m}$, assuming $Q$ is sufficiently smooth and Fréchet differentiable with respect to uncertain parameter is derived as,
-$Q_{\text{QUAD}}(m)$ $=$ $\sum_{l=0}^{2}\partial^{l}_{m}Q(\bar{m})\,(m-\bar{m})^{l}$ (22)
-$=$ $\bar{Q}+\big{<}\bar{Q}^{m},\,m-\bar{m}\big{>}+\frac{1}{2}\big{<}\bar{Q}^{mm}\,(m-\bar{m}),\,m-\bar{m}\big{>}.$
-where $\partial^{l}_{m}Q(\bar{m})$ is the $l$-th order tensor derivative at $\bar{m}$ and the objective $Q$ and its gradient and Hessian with respect to $m$, evaluated at mean $\bar{m}$, denoted as $\bar{Q}$, $\bar{Q}^{m}$ and $\bar{Q}^{mm}$, respectively.
-The close form for the expectation of the quadratic Taylor approximation in (22), considering $\mathbb{E}[\bar{Q}]=\bar{Q}$ and $\mathbb{E}[\big{<}\bar{Q}^{m},\,m-\bar{m}\big{>}]=0$, and Lemma 1 of [47], is obtained as
-$\mathbb{E}[Q_{\text{QUAD}}]=\bar{Q}+\frac{1}{2}\mathrm{tr}\big{(}\bar{\mathcal{H}}_{c}\big{)},$ (23)
-where ¯$\bar{\mathcal{H}}_{c}=\mathcal{C}\,\bar{Q}^{mm}$ denotes the covariance preconditioned Hessian computed at the mean of $m$ and $\mathrm{tr}(\cdot)$ is trace operator.
-Moreover, noting that variance is invariant with respect to its deterministic translation, e.g., $\mathbb{V}(m-\mathrm{const.})=\mathbb{V}(m)$), one can write,
-$\mathbb{E}[(Q_{\text{QUAD}}(m)-\bar{Q})^{2}]=\mathbb{E}[(\big{<}\bar{Q}^{m},\,m-\bar{m}\big{>}+\frac{1}{2}\big{<}\bar{Q}^{mm}(m-\bar{m}),\,m-\bar{m}\big{>})^{2}],$ (24)
-which, upon expansion, allows each term to be expressed as
-$\mathbb{E}[(\big{<}\bar{Q}^{m},\,m-\bar{m}\big{>})^{2}]$ $=$ $\big{<}\bar{Q}^{m},\mathcal{C}\,\bar{Q}^{m}\big{>},$
-$\mathbb{E}[(\big{<}\bar{Q}^{mm}(m-\bar{m}),\,m-\bar{m}\big{>})^{2}]$ $=$ $\mathrm{tr}(\bar{\mathcal{H}}_{c})^{2}+2\,\mathrm{tr}((\bar{\mathcal{H}}_{c})^{2}),$
-$\mathbb{E}[\big{<}\bar{Q}^{m},\,m-\bar{m}\big{>}\,\big{<}\bar{Q}^{mm}(m-\bar{m}),\,m-\bar{m}\big{>}]$ $=$ $0$ (25)
-Furthermore, the square of the first moment of $Q_{\text{QUAD}}(m)-\bar{Q}$ can be written as,
-$\big{(}\mathbb{E}[Q_{\text{QUAD}}(m)-\bar{Q}]\big{)}^{2}=\big{(}\mathbb{E}[Q_{\text{QUAD}}(m)]-\mathbb{E}[\bar{Q}]\big{)}^{2}=\frac{1}{4}\big{(}\mathrm{tr}\big{(}\bar{\mathcal{H}}_{c}\big{)}\big{)}^{2}$ (26)
-From (25) and (26), the variance of the quadratic Taylor approximation of $Q$ in (22) is obtained as,
-$\mathbb{V}[Q_{\text{QUAD}}(m)]=\mathbb{V}[Q_{\text{QUAD}}(m)-\bar{Q}]=\big{<}\bar{Q}^{m},\mathcal{C}\,\bar{Q}^{m}\big{>}+\frac{1}{2}\mathrm{tr}((\bar{\mathcal{H}}_{c})^{2}).$ (27)
-Consequently, the mean and variance of the design objective in (12) can be replaced by their corresponding quadratic Taylor approximations in (23) and (27), respectively, leading to the approximated design objective as
-$\mathcal{J}_{\mathrm{QUAD}}(d)=\bar{Q}+\frac{1}{2}\mathrm{tr}\big{(}\bar{\mathcal{H}}_{c}\big{)}+\beta_{V}\,\bigg{(}\big{<}\bar{Q}^{m},\mathcal{C}\,\bar{Q}^{m}\big{>}+\frac{1}{2}\mathrm{tr}((\bar{\mathcal{H}}_{c})^{2})\bigg{)}+\beta_{R}\,R(d).$ (28)
-Similar to (22), we can construct a second-order Taylor expansion of the constraint function $f$ at $\bar{m}$ as,
-$f_{QUAD}(m)=\bar{f}+\big{<}\bar{f}^{m},\,m-\bar{m}\big{>}+\frac{1}{2}\big{<}\bar{f}^{mm}\,(m-\bar{m}),\,m-\bar{m}\big{>}.$ (29)
+$P(f(.,d)\geq 0)$ $\approx\frac{1}{N_{mc}}\sum_{i=1}^{N_{mc}}\chi_{[0,\infty)]}(f(m^{(i)},d))$ (15)
+where $\{m^{(i)}\}_{i=1}^{N_{mc}}$ are random samples from $m$.
+Achieving an accurate approximation requires a large number of finite element model evaluations due to the slow convergence of the Monte Carlo method, making optimization under uncertainty computationally prohibitive.
+To address this, we introduce a computationally efficient approach that approximates the mean and variance of the design objective using a quadratic Taylor expansion, leveraging this approximation as a control variate to accelerate Monte Carlo sampling.
+3.1 Quadratic Taylor approximation of the design objective and chance constraint
+The quadratic Taylor approximation of $Q$ at the mean of the uncertain parameter $\bar{m}$ is given by [24, 13, 14],
+$Q_{\text{QUAD}}(m)=\bar{Q}+\big{<}\bar{Q}^{m},\,m-\bar{m}\big{>}+\frac{1}{2}\big{<}\bar{Q}^{mm}\,(m-\bar{m}),\,m-\bar{m}\big{>}.$ (16)
+Here, the objective $Q$ and its gradient and Hessian with respect to $m$, evaluated at mean $\bar{m}$, are denoted as $\bar{Q}$, $\bar{Q}^{m}$ and $\bar{Q}^{mm}$, respectively.
+The expectation and variance of $Q_{\text{QUAD}}$ are derived in [24] such that,
+$\mathbb{E}[Q_{\text{QUAD}}]=\bar{Q}+\frac{1}{2}\mathrm{tr}\big{(}\bar{\mathcal{H}}_{k}\big{)},\quad\mathbb{V}[Q_{\text{QUAD}}(m)]=\big{<}\bar{Q}^{m},\mathcal{K}\,\bar{Q}^{m}\big{>}+\frac{1}{2}\mathrm{tr}((\bar{\mathcal{H}}_{k})^{2}),$ (17)
+where ¯$\bar{\mathcal{H}}_{k}=\mathcal{K}\,\bar{Q}^{mm}$ represents the preconditioned Hessian for covariance at the mean of the uncertain parameter $m$ and $\mathcal{K}$ is the covariance operator.
+The design objective in (7) can then be approximated as
+$\mathcal{J}_{\mathrm{QUAD}}(d)=\bar{Q}+\frac{1}{2}\mathrm{tr}\big{(}\bar{\mathcal{H}}_{k}\big{)}+\beta_{V}\,\bigg{(}\big{<}\bar{Q}^{m},\mathcal{K}\,\bar{Q}^{m}\big{>}+\frac{1}{2}\mathrm{tr}((\bar{\mathcal{H}}_{k})^{2})\bigg{)}+\beta_{RG}\,R_{p}(d).$ (18)
+We can also construct a second-order Taylor expansion of the chance constraint function $f$ at $\bar{m}$ as,
+$f_{QUAD}(m)=\bar{f}+\big{<}\bar{f}^{m},\,m-\bar{m}\big{>}+\frac{1}{2}\big{<}\bar{f}^{mm}\,(m-\bar{m}),\,m-\bar{m}\big{>},$ (19)
+where $\bar{f}$, $\bar{f}^{m}$ and $\bar{f}^{mm}$ are the constraint function $f$ and it gradient and Hessian evaluated at $\bar{m}$.
 The probability $P(f(.,d)\geq 0)$ by Monte Carlo estimator (15) with the quadratic approximation yields,
-$P(f(m,d))\geq 0)=\frac{1}{M_{f}}\mathbb{I}_{[0,\infty)}(f_{\text{QUAD}}(m_{i},d)).$ (30)
-This approximation of the constraint function obviates the necessity for a large number of PDE solves to assess the probability utilizing sample averaging methods.
-In upcoming sections, we present the Lagrangian formalism to derive the $m$-gradient and $m$-Hessian, representing the gradient and Hessian of $Q$ with respect to $m$.
-We also introduce an efficient method for evaluating the covariance preconditioned Hessian, ¯$\bar{\mathcal{H}}_{c}$.
-3.1 Analytical gradient and hessian of design objective with respect to the uncertain parameter
-The Lagrangian functional, considering the adjoint variable $\mathbf{v}^{Q}$ as the Lagrange multiplier associated with the state equation with respect to the objective function $Q$, is defined as,
-$\mathcal{L}(\mathbf{u},\mathbf{v}^{Q},m,d)=Q(\mathbf{u},m,d)+r(\mathbf{u},\mathbf{v}^{Q},m,d),$ (31)
-where $r(\mathbf{u},\mathbf{v},m,d)$ is the weak form in (7).
-By imposing the condition that the first-order variation of the Lagrangian functional at $\bar{m}$ with respect to the adjoint variable $\mathbf{v}$ and the state variable $\mathbf{u}$ is equal to zero, we obtain, the state problem:
-$\textit{Find $\mathbf{u}\in\mathcal{U}$, such that}\quad\langle\tilde{\mathbf{v}},\partial_{\mathbf{v}}\bar{r}\rangle=0,\quad\forall\tilde{\mathbf{v}}\in\mathcal{V},$ (32)
-and adjoint problem:
-$\textit{Find $\mathbf{v}\in\mathcal{V}$, such that}\quad\langle\tilde{\mathbf{u}},\partial_{\mathbf{u}}{\bar{r}^{Q}}\rangle=-\langle\tilde{\mathbf{u}},\partial_{\mathbf{u}}\bar{Q}\rangle,\quad\forall\tilde{\mathbf{u}}\in\mathcal{U}.$ (33)
-In the above relations, $(\bar{\cdot})$ denotes the quantities computed at the mean of the uncertain parameter, and $(\tilde{\cdot})$ indicates directional variables.
-Directional variables are referred to the variables that define the direction along which perturbation is considered to compute the change in the functional to obtain the derivative of the functional.
-Also, here we define $r^{Q}=r(\mathbf{u},\mathbf{v}^{Q},m,d)$ for ease of notation.
-The gradient of $Q$ at $\bar{m}$ acting in the direction $\tilde{m}=m-\bar{m}$ is then given by,
-$\langle\tilde{m},\bar{Q}^{m}\rangle=\langle\tilde{m},\partial_{m}\mathcal{L}\rangle.=\langle\tilde{m},\partial_{m}\bar{Q}+\partial_{m}{\bar{r}^{Q}}\rangle\quad\forall\tilde{m}\in\mathcal{M}$ (34)
-Thus, evaluating $m$-gradient in (34) requires the solution of state problem (32) for $\mathbf{u}$ and the adjoint problem (33) for $\mathbf{v}$.
-To solve for the $m$-Hessian, consider a meta-Lagrangian functional with incremental state $\hat{\mathbf{u}}$ and adjoint $\hat{\mathbf{v}}$
-$\small\mathcal{L}^{H}(\mathbf{u},\mathbf{v}^{Q},m,d;\hat{\mathbf{u}}^{Q},\hat{\mathbf{v}}^{Q},\hat{m}^{Q})=\langle\hat{m}^{Q},\partial_{m}\bar{r}^{Q}+\partial_{m}\bar{Q}\rangle+\langle\hat{\mathbf{v}}^{Q},\partial_{\mathbf{v}}\bar{r}^{Q}\rangle+\langle\hat{\mathbf{u}}^{Q},\partial_{\mathbf{u}}\bar{r}^{Q}+\partial_{\mathbf{u}}\bar{Q}\rangle$ (35)
-where the terms on the right-hand side are $m$-gradient, forward, and adjoint problems, respectively.
-Repeating the similar process of vanishing the first variation of the Lagrangian functional with respect to adjoint and state, we obtain the incremental state problem:
-$\textit{Find $\hat{\mathbf{u}}\in\mathcal{U}$, such that}\quad\langle\tilde{\mathbf{v}},\partial_{\mathbf{vu}}\bar{r}^{Q}\,\hat{\mathbf{u}}^{Q}\rangle=-\langle\tilde{\mathbf{v}},\partial_{\mathbf{v}m}\bar{r}^{Q}\,\hat{m}^{Q}\rangle,\quad\forall\tilde{\mathbf{v}}\in\mathcal{V},$ (36)
-and the incremental adjoint problem:
-Find $\hat{\mathbf{v}}\in\mathcal{V}$, such that
-$\langle\tilde{\mathbf{u}},\partial_{\mathbf{uv}}\bar{r}^{Q}\,\hat{\mathbf{v}}^{Q}\rangle=-\langle\tilde{\mathbf{u}},\partial_{\mathbf{uu}}\bar{r}^{Q}\,\hat{\mathbf{u}}+\partial_{\mathbf{uu}}\bar{Q}\,\hat{\mathbf{u}}^{Q}+\partial_{\mathbf{u}m}\bar{r}^{Q}\,\hat{m}^{Q}+\partial_{\mathbf{u}m}\bar{Q}\,\hat{m}^{Q}\rangle,\quad\forall\tilde{\mathbf{u}}\in\mathcal{U}.$ (37)
-By taking the variation of $\mathcal{L}^{H}$ with respect to $m$ as zero, the Hessian $Q_{mm}(\bar{m})$ acting on $\hat{m}$ is computed as,
-$\begin{split}\langle\tilde{m},\bar{Q}^{mm}\,\hat{m}^{Q}\rangle=&\langle\tilde{m},\partial_{m\mathbf{v}}\bar{r}^{Q}\,\hat{\mathbf{v}}^{Q}+\partial_{m\mathbf{u}}\bar{r}^{Q}\,\hat{\mathbf{u}}^{Q}+\partial_{m\mathbf{u}}\bar{Q}\,\hat{\mathbf{u}}^{Q}+\partial_{mm}\bar{r}^{Q}\,\hat{m}^{Q}\\ +\partial_{mm}\bar{Q}\,\hat{m}^{Q}\rangle\end{split}$ (38)
-We remark that the last term on the right-hand side of (37) and the last two terms on the right-hand side of (38) are included due to the explicit dependence of the design objective (LABEL:eq:qois) on the uncertain parameter in our problem.
-Thus, evaluating the $m$-Hessian acting on $\hat{m}$ in (38) requires the solution of the incremental state problem (36) for $\hat{\mathbf{u}}$ and the incremental adjoint problem (37) for $\hat{\mathbf{v}}$.
-Similarly, we can compute the gradient and Hessian action of $f$. The adjoint problem can be written as:
-$\langle\tilde{\mathbf{u}},\partial_{\mathbf{u}}\bar{r}^{f}\rangle=-\langle\tilde{\mathbf{u}},\partial_{\mathbf{u}}\bar{f}\rangle,\quad\forall\tilde{\mathbf{u}}\in\mathcal{U}$ (39)
-The gradient of $f$ at $\bar{m}$ can be defined as:
-$\langle\tilde{m},\bar{f}^{m}\rangle=\langle\tilde{m},\partial_{m}\bar{f}+\partial_{m}\bar{r}^{f}\rangle\quad\forall\tilde{m}\in\mathcal{M}$ (40)
-Then by solving incremental state problem,
-[…]
-and the incremental adjoint problem
-[…]
-We obtain the Hessian action of $f$ at $\tilde{m}$
-$\begin{split}\langle\tilde{m},\bar{f}^{mm}\hat{m}^{f}\rangle=\langle\tilde{m},\partial_{m\mathbf{v}}\bar{r}^{f}\hat{\mathbf{v}}^{f}+\partial_{m\mathbf{u}}\bar{r}^{f}\hat{\mathbf{u}}^{f}+\partial_{m\mathbf{u}}\bar{f}\hat{\mathbf{u}}^{f}+\partial_{mm}\bar{r}^{f}\hat{m}^{f}+\partial_{mm}\bar{f}\hat{m}^{f}\rangle\quad&\\ \forall\tilde{m}\in\mathcal{M}\end{split}$ (43)
+$P(f(m,d))\geq 0)=\frac{1}{N_{mc}}\sum_{i=1}^{N_{mc}}\chi_{[0,\infty)}(f_{\text{QUAD}}(m_{i},d)).$ (20)
+In order to compute the second-order Taylor expansion (19), we need to compute the term $\big{<}\bar{f}^{mm}\,(m-\bar{m}),\,m-\bar{m}\big{>}$, which can be approximated through a low-rank approximation as,
+$\big{<}\bar{f}^{mm}(m-\bar{m}),m-\bar{m}\big{>}=\sum_{i=1}^{N_{\mathrm{eig}}^{f}}\lambda_{i}^{f}\big{|}\big{<}m-\bar{m},\mathcal{K}^{-1}\psi_{i}^{f}\big{>}\big{|}^{2},$ (21)
+where $\lambda_{i}^{f}$ and $\psi_{i}^{f}$ are the eigenvalues and eigenfunctions corresponding to the solution of a generalized eigenvalue problem.
+The generalized eigenvalue problem corresponding to the second-order Taylor expansion can be formulated as,
+$\mathcal{K}^{1/2}\>\bar{f}^{mm}\>\mathcal{K}^{1/2}\>\psi_{i}^{f}=\lambda_{i}^{f}\psi_{i}^{f},\quad i=1,2,\dots,N_{\mathrm{eig}}^{f},$ (22)
+where $\lambda_{1}^{f},\lambda_{2}^{f},...,\lambda_{N_{\mathrm{eig}}^{f}}^{f}$ are the $N_{\mathrm{eig}}^{f}$ largest eigenvalues.
+Computing the approximations of the mean, variance, and constraint function requires a Lagrangian-based formalism to derive the $m$-gradient, $m^{f}$-gradient, $m$-Hessian and $m^{f}$-Hessian, which represent the gradients and Hessians of the design objective $Q$ and constraint function $f$ with respect to $m$, as detailed in Section A.
+Finally, the the mean and variance in (17) requires the trace of the covariance-preconditioned Hessian ¯$\bar{\mathcal{H}}_{k}$.
+Here, we leverage a randomized trace estimator using the dominant eigenvalues $\{\lambda^{Q}_{n}\}_{n=1}^{N^{Q}_{\mathrm{eig}}}$ of ¯$\bar{\mathcal{H}}_{k}$,
+$\mathrm{tr}\big{(}\bar{\mathcal{H}}_{k}\big{)}\approx\sum_{n\geq 1}^{N^{Q}_{\mathrm{eig}}}\lambda^{Q}_{n},\quad\mathrm{tr}\big{(}\bar{\mathcal{H}}_{k}^{2}\big{)}\approx\sum_{n\geq 1}^{N^{Q}_{\mathrm{eig}}}{(\lambda^{Q}_{n})}^{2}.$ (23)
+When eigenvalues decay rapidly, $N^{Q}_{\mathrm{eig}}$ only depend on the intrinsic low dimensionality of the otherwise high-dimensional parameters.
+The eigenvalues $\lambda^{Q}_{n}$ are obtained by solving the generalized eigenvalue problem with an input of oversampling factor $N_{o}$, via double-pass randomized eigenvalue solver (see [24, 14] for more details).
 3.2 Monte Carlo with control variate
 The approximation of $\mathbb{E}[Q]$ and $\mathbb{V}[Q]$ in the cost functional (7) using a quadratic Taylor series expansion may lack accuracy for the design problem, particularly for design problems where the variance of the uncertain parameter $m$ is large, as shown in [24].
-To enhance accuracy, the quadratic approximation $Q_{\text{QUAD}}$ is leveraged as a control variate to reduce variance in the Monte Carlo estimators for the mean and variance.
-This involves correcting the moments of $Q_{\text{QUAD}}$ by applying a Monte Carlo method to estimate the mean and variance of the residual between $Q$ and its quadratic approximation $Q_{\text{QUAD}}$.
-We can use the same $M=M_{1}=M_{2}$ random samples for the approximation of the mean and variance.
+To enhance accuracy, the $Q_{\text{QUAD}}$ is leveraged as a control variate to reduce variance in the sampling estimations of the mean and variance.
+In particular, A Monte Carlo method is employed to correct the moments of $Q_{\text{QUAD}}$ by computing the mean and variance of the residual between $Q$ and its quadratic approximation $Q_{\text{QUAD}}$.
 For mean of control variate approximation
-$\mathbb{E}[Q_{CV}]\approx\mathbb{E}[Q]=\mathbb{E}[Q_{\text{QUAD}}]+\mathbb{E}[Q-Q_{\text{QUAD}}]:=Q(\bar{m})+\frac{1}{2}tr(\mathcal{H})$ (44)
-$+\frac{1}{M}\sum_{i=1}^{M}\left(Q(m_{i})-Q(\bar{m})-\langle m_{i}-\bar{m},Q^{m}(\bar{m})\rangle-\frac{1}{2}\langle m_{i}-\bar{m},Q^{mm}(\bar{m})(m_{i}-\bar{m})\rangle\right).$
+$\mathbb{E}[Q_{CV}]\approx\mathbb{E}[Q]=\mathbb{E}[Q_{\text{QUAD}}]+\mathbb{E}[Q-Q_{\text{QUAD}}]:=Q(\bar{m})+\frac{1}{2}tr(\bar{\mathcal{H}}_{k})$ (24)
+$+\frac{1}{N_{mc}}\sum_{i=1}^{N_{mc}}\left(Q(m^{(i)})-Q(\bar{m})-\langle m^{(i)}-\bar{m},Q^{m}(\bar{m})\rangle-\frac{1}{2}\langle m^{(i)}-\bar{m},Q^{mm}(\bar{m})(m^{(i)}-\bar{m})\rangle\right).$
 The variance can be written as
@@
 which can be approximated by
-$\mathbb{V}[Q_{CV}]:=\langle\mathcal{C}Q^{m}(\bar{m}),Q^{m}(\bar{m})\rangle+\frac{1}{4}{(tr(\mathcal{H}))}^{2}+\frac{1}{2}tr(\mathcal{H}^{2})$ (46)
-$+\frac{1}{M}\sum_{i=1}^{M}\left({(Q(m_{i})-Q(\bar{m})}^{2}-{\left(\langle m_{i}-\bar{m},Q^{m}(\bar{m})\rangle+\frac{1}{2}\langle m_{i}-\bar{m},Q^{mm}(\bar{m})(m_{i}-\bar{m})\rangle\right)}^{2}\right)$
-$-{\left(\frac{1}{2}tr(\mathcal{H})+\frac{1}{M}\sum_{i=1}^{M}\left(Q(m_{i})-\langle m_{i}-\bar{m},Q^{m}(\bar{m})\rangle-\frac{1}{2}\langle m_{i}-\bar{m},Q^{mm}(\bar{m})(m_{i}-\bar{m})\rangle\right)\right)}^{2}.$
-The number of samples $M$ required to achieve the desired accuracy (and hence the computational cost) depends on the variance of the integrand and how closely the quadratic approximation matches the high-fidelity Monte Carlo estimation.
-A closer approximation reduces the required $M$, while a poorer match increases it.
+$\mathbb{V}[Q_{CV}]:=\langle\mathcal{K}Q^{m}(\bar{m}),Q^{m}(\bar{m})\rangle+\frac{1}{4}{(tr({\bar{\mathcal{H}}_{k}}))}^{2}+\frac{1}{2}tr({\bar{\mathcal{H}}_{k}}^{2})$ (26)
+$+\frac{1}{N_{mc}}\sum_{i=1}^{N_{mc}}\left({(Q(m^{(i)})-Q(\bar{m})}^{2}-{\left(\langle m^{(i)}-\bar{m},Q^{m}(\bar{m})\rangle+\frac{1}{2}\langle m^{(i)}-\bar{m},Q^{mm}(\bar{m})(m^{(i)}-\bar{m})\rangle\right)}^{2}\right)$
+$-{\left(\frac{1}{2}tr({\bar{\mathcal{H}}_{k}})+\frac{1}{N_{mc}}\sum_{i=1}^{N_{mc}}\left(Q(m^{(i)})-\langle m^{(i)}-\bar{m},Q^{m}(\bar{m})\rangle-\frac{1}{2}\langle m^{(i)}-\bar{m},Q^{mm}(\bar{m})(m^{(i)}-\bar{m})\rangle\right)\right)}^{2}.$
+The number of samples $N_{mc}$ required to achieve the desired accuracy (and hence the computational cost) depends on the variance of the integrand and how closely the quadratic approximation matches the high-fidelity Monte Carlo estimation.
 Notably, if the quadratic approximation is highly correlated with $Q$, fewer Monte Carlo samples are needed for effective variance reduction and accurate approximation.
-3.3 Randomized trace estimator
-The equations for the mean (23) and variance (27) of the design objectives evaluated using the quadratic approximation requires the trace of the covariance-preconditioned Hessian ¯$\bar{\mathcal{H}}c$, which typically comes out to be a large dense matrix.
-The traditional approach is based on Monte Carlo method, where the Hessian action is computed on a certain number of randomly chosen directions from a Gaussian distribution [18].
-Despite adhering to the lower bounds on the required number of samples for accurate trace approximations, as outlined in [48], a large number of samples is still necessary.
-Given the large sample size needed at each optimization iteration, this approach becomes computationally prohibitive for the design under uncertainty problem addressed in this work.
-Here, we implement a randomized approximated eigen-decomposition for computing the trace and trace-squared using the dominant eigenvalues $\{\lambda_{n}\}_{n=1}^{N_{\mathrm{eig}}}$ of the covariance-preconditioned Hessian ¯$\bar{\mathcal{H}}_{c}$, expressed as,
-$\mathrm{tr}\big{(}\bar{\mathcal{H}}_{c}\big{)}\approx\sum_{n\geq 1}^{N_{\mathrm{eig}}}\lambda_{n},\quad\mathrm{tr}\big{(}\bar{\mathcal{H}}_{c}^{2}\big{)}\approx\sum_{n\geq 1}^{N_{\mathrm{eig}}}\lambda_{n}^{2}.$ (47)
-The above forms of trace estimators are justified by the symmetric positive definite property of the operator $\mathcal{H}_{c}$.
-It is important to note that the value of $N_{\mathrm{eig}}$ does not depend on the actual parameter dimension, but is dependent on the effective dimension of the parameter, e.g., [49, 50, 51, 52] it tends to be small when the covariance-preconditioned Hessian is low-rank or when its eigenvalues decay rapidly (see the numerical results of Section 5).
-The eigenvalues $\lambda_{n}$ are obtained by solving the following generalized eigenvalue problem,
-$\big{<}\zeta,\bar{Q}^{mm}\,\psi_{j}\big{>}=\lambda_{j}\big{<}\zeta,\mathcal{C}^{-1}\psi_{j}\big{>},\quad\forall\zeta\in\mathcal{M},\quad j=1,...,N_{\mathrm{eig}}$ (48)
-where the eigenvectors $\psi_{j}$ exhibit orthonormality with $\mathcal{C}^{-1}$, $\big{<}\psi_{i},\mathcal{C}^{-1}\psi_{j}\big{>}=\delta_{ij},i,j=1,...,N_{\mathrm{eig}}$ and $\delta_{ij}$ denotes the Kronecker delta function.
-Using an $n$-degree finite element discretization, (48) is rewritten in the form of matrix as $\mathbf{M}\,\boldsymbol{\psi}=\lambda\,\mathbf{K}\,\boldsymbol{\psi}$ where $\mathbf{M},\mathbf{K}$ are $n$-by-$n$ symmetric matrices, and $\mathbf{K}$ is positive definite, $\boldsymbol{\psi}$ represents a vector of size $n$.
-For solving the resulting eigenproblem, we implement a randomized algorithm [53], discussed in Algorithm 1.
-1
-Input : Matrices $\mathbf{M},\mathbf{K}\in\mathbb{R}^{n\times n}$ obtained from finite element discretization, number of eigenpairs requested $N_{\mathrm{eig}}$, oversampling parameter $N_{\mathrm{p}}$
-Output : Eigenpairs $(\lambda_{j},\psi_{j})$ for $j=1,\dots,N_{\mathrm{eig}}$
-2
-3Generate random sample and draw a Gaussian random matrix $\mathbf{\Omega}\in\mathbb{R}^{n\times(N_{\mathrm{eig}}+N_{\mathrm{p}})}$;
-4
-5Form the range matrix $\mathbf{Z}=\mathbf{K}^{-1}(\mathbf{M}\mathbf{\Omega})$;
-6
-7Perform QR factorization $\mathbf{Z}=\mathbf{Q}\mathbf{R}$ such that $\mathbf{Q}^{\top}\mathbf{M}\mathbf{Q}=\mathbf{I}_{N_{\mathrm{eig}}+N_{\mathrm{p}}}$;
-8
-9Compute the reduced matrix $\mathbf{Q}^{\top}\mathbf{K}\mathbf{Q}$;
-10
-11Compute the eigendecomposition $\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{\top}$, where $\boldsymbol{\Lambda}=\mathrm{diag}(\lambda_{1},\lambda_{2},\dots)$;
-12
-13Obtain the eigenpairs of the original problem:
-$\lambda_{j}=\boldsymbol{\lambda}[j,j],\quad\psi_{j}=\mathbf{Q}\mathbf{U}[:,j],\quad j=1,\dots,N_{\mathrm{eig}}.$
-Algorithm 1 Double-Pass Randomized Eigenvalue Solver
 4 Gradient-based optimization
@@
 However, applying INCG to solve (12) in a scalable manner requires special treatment of the chance constraints and efficient computation of the gradient of the cost functional with respect to high-dimensional design parameters, as discussed in this section.
-4.1 Smooth approximation of the discontinuous function
-The indicator function utilized in the chance constraint exhibits discontinuity at $f(m,d)=0$.
-To employ the gradient-based INCG, this discontinuous indicator function must be approximated with a smooth, continuous counterpart.
-One approach involves employing a logistic function as a smooth approximation of the indicator function [25] such that,
-$\mathbb{I}_{[0,\infty]}(x)\approx l_{\omega}(x)=\frac{1}{1+e^{-2\omega x}},$ (49)
-where a larger $\omega$ corresponds to a sharper transition at $x=0$ as shown in Figure 3.
-This smooth approximation tackles the non-differentiability of the chance constraint and makes the cost functional differentiable with respect to the design variable.
-Figure 3: Approximation of the discontinuous indicator function $\mathbb{I}_{[0,\infty]}(x)$ with a smooth logistic function as given in (49) and its convergence with increasing $\omega$
+4.1 Continuous approximation of the discontinuous function
+The indicator function utilized in the chance constraint is discontinuous at $f(m,d)=0$.
+To employ the gradient-based INCG, this discontinuous indicator function must be approximated with a continuous function.
+A common approach involves using a sigmoid function to approximate the indicator function, as proposed by [21], given by
+$\chi_{[0,\infty)}(x)\approx l_{\omega}(x)=\frac{1}{1+e^{-2\omega x}},$ (27)
+where the parameter $\omega$ controls the sharpness of the transition around $x=0$, as illustrated in Figure 3.
+A higher value of $\omega$ results in a steeper transition.
+This smooth approximation addresses the non-differentiability of the chance constraint, enabling the cost functional to be differentiable with respect to the design variable.
+The value of $\omega$ is gradually increased through a continuation scheme discussed in Algorithm 2.
+This gradual scaling of $\omega$ helps in avoiding local optima and prevents numerical instabilities.
+Figure 3: Approximation of the discontinuous indicator function $\chi_{[0,\infty)}(x)$ with a sigmoid function as given in (27) and its convergence with increasing $\omega$
 4.2 Penalty Method
-To impose the inequality constraint (15), a quadratic penalty method [55] is applied. The quadratic penalty function is defined as,
-$S_{\gamma}(x)=\frac{\gamma}{2}{(max{\{0,x\}})}^{2}$ (50)
-where $\gamma>0$ is a constant corresponding to the weight of the penalty. Using the penalty method, the chance-constrained problem can be rewritten as
-$\underset{d\in[0,1]}{\text{min}}\mathcal{J}(d)+\mathcal{S}_{\gamma}(\mathbb{E}[l_{\omega}(f)]-\alpha_{c})$ (51)
-This transformation simplifies the inequality-constrained problem by converting it into an unconstrained form.
-4.3 Design gradients via adjoint
-The Lagrangian formalism is employed to derive the gradient of the quadratic approximation of the design objective $\mathcal{J}_{\mathrm{QUAD}}(d)$ with respect to the design parameter, referred to as the d-gradient.
-This gradient provides critical information to the INCG optimizer, in order to facilitate gradient-based optimization processes.
-By substituting the trace estimators (47) into (28) and converting it into unconstrained problem using quadratic penalty (50) and smooth approximation (49), we have
-$\mathcal{J}_{\mathrm{QUAD}}(d)$ $=\bigg{(}\bar{Q}+\frac{1}{2}\sum_{j\geq 1}^{N_{\mathrm{eig}}}\lambda_{j}^{Q}\bigg{)}+\beta_{V}\bigg{(}\big{<}\bar{Q}^{m},\mathcal{C}\,\bar{Q}^{m}\big{>}+\frac{1}{2}\sum_{j\geq 1}^{N_{\mathrm{eig}}}\left({\lambda_{j}^{Q}}\right)^{2}\bigg{)}+\beta_{R}\,R(d)$ (52)
-$+S_{\gamma}(\mathbb{E}[l_{\omega}(f_{\text{QUAD}}(m,d))]-\alpha_{c})$
-To evaluate the $d$-gradient, we consider a quadratic meta-Lagrangian functional
-$\mathcal{L}_{\mathrm{QUAD}}$ $(d,\mathbf{u},\mathbf{v}^{*},\{{\lambda_{j}^{Q}}\},\{{\psi_{j}}^{Q}\},\{{\hat{\mathbf{u}}_{j}^{Q}}\},\{{\hat{\mathbf{v}}_{j}}^{Q}\},$ (53)
-$\left(\mathbf{u}^{Q}\right)^{*},\left(\mathbf{v}^{Q}\right)^{*},\{\left(\lambda_{j}^{Q}\right)^{*}\},\{\left(\psi_{j}^{Q}\right)^{*}\},\{\left(\hat{\mathbf{u}}_{j}^{Q}\right)^{*}\},\{\left(\hat{\mathbf{v}}_{j}^{Q}\right)^{*}\},$
-$\{{\lambda_{j}^{f}}\},\{{\psi_{j}}^{f}\},\{{\hat{\mathbf{u}}_{j}}^{f}\},\{{\hat{\mathbf{v}}_{j}}^{f}\},\left(\mathbf{u}^{f}\right)^{*},\left(\mathbf{v}^{f}\right)^{*},\{\left(\lambda_{j}^{f}\right)^{*}\},\{\left(\psi_{j}^{f}\right)^{*}\},\{\left(\hat{\mathbf{u}}_{j}^{f}\right)^{*}\},\{\left(\hat{\mathbf{v}}_{j}^{f}\right)^{*}\})$
-$=\mathcal{J}_{\mathrm{QUAD}}(d)+\langle\mathbf{v}^{*},\partial_{\mathbf{v}}\bar{r}^{Q}\rangle$
-$+\langle\left(\mathbf{u}^{Q}\right)^{*},\partial_{\mathbf{u}}\bar{r}^{Q}+\partial_{\mathbf{u}}\bar{Q}\rangle$
-$+\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\langle\left(\psi_{j}^{Q}\right)^{*},(\bar{Q}^{mm}-{\lambda_{j}^{Q}}\mathcal{C}^{-1})\psi_{j}^{Q}\rangle$
-$+\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\left(\lambda_{j}^{Q}\right)^{*}(\langle{\psi_{j}}^{Q},\mathcal{C}^{-1}{\psi_{j}}^{Q}\rangle-1)+\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\langle{\hat{\mathbf{v}}_{j}}^{{Q}^{*}},\partial_{\mathbf{vu}}\bar{r}_{Q}\,{\hat{\mathbf{u}}_{j}}^{Q}+\partial_{\mathbf{v}m}\bar{r}_{Q}\,{\psi_{j}}^{Q}\rangle$
-$+\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\langle\left(\hat{\mathbf{u}}_{j}^{Q}\right)^{*},\partial_{\mathbf{uv}}\bar{r}_{Q}\,{\hat{\mathbf{v}}_{j}}^{Q}+\partial_{\mathbf{uu}}\bar{r}_{Q}\,\left(\mathbf{u}^{Q}\right)^{*}+\partial_{\mathbf{uu}}\bar{Q}\,{\hat{\mathbf{u}}_{j}}^{Q}+\partial_{\mathbf{u}m}\bar{Q}\,{\psi_{j}}^{Q}+\partial_{\mathbf{u}m}\bar{r}_{Q}\,{\psi_{j}}^{Q}\rangle$
-$+\langle\left(\mathbf{u}^{f}\right)^{*},\partial_{\mathbf{u}}\bar{r}_{f}+\partial_{\mathbf{u}}\bar{f}\rangle$
-$+\sum^{N_{\mathrm{eig}}^{f}}_{j=1}\langle\left(\psi_{j}^{f}\right)^{*},(\bar{f}^{mm}-{\lambda_{j}^{f}}\mathcal{C}^{-1})\psi_{j}^{f}\rangle$
-$+\sum^{N_{\mathrm{eig}}^{f}}_{j=1}\left(\lambda_{j}^{f}\right)^{*}(\langle{\psi_{j}}^{f},\mathcal{C}^{-1}{\psi_{j}}^{f}\rangle-1)+\sum^{N_{\mathrm{eig}}^{f}}_{j=1}\langle\left(\hat{\mathbf{v}}_{j}^{f}\right)^{*},\partial_{\mathbf{vu}}\bar{r}_{f}\,{\hat{\mathbf{u}}_{j}}^{f}+\partial_{\mathbf{v}m}\bar{r}_{f}\,{\psi_{j}}^{f}\rangle$
-$+\sum^{N_{\mathrm{eig}}^{f}}_{j=1}\langle\left(\hat{\mathbf{u}}_{j}^{f}\right)^{*},\partial_{\mathbf{uv}}\bar{r}_{f}\,{\hat{\mathbf{v}}_{j}}^{f}+\partial_{\mathbf{uu}}\bar{r}_{f}\,\left(\mathbf{u}^{f}\right)^{*}+\partial_{\mathbf{uu}}\bar{f}\,{\hat{\mathbf{u}}_{j}}^{f}+\partial_{\mathbf{u}m}\bar{f}\,{\psi_{j}}^{f}+\partial_{\mathbf{u}m}\bar{r}_{f}\,{\psi_{j}}^{f}\rangle$
-where $(\cdot^{*})$ represents the adjoint variables, $(\hat{\cdot})$ denotes the incremental variables.
-The terms on the right hand side of (53), correspond to the following: the quadratic approximation of the design objective (28), the state (32) and adjoint (33) problems for design objective $Q$, the generalized eigenvalue problems (48) for design objective $Q$, the orthonormality constraint of the eigenvectors for design objective $Q$, the incremental state problem (36) of design objective $Q$ for eigenvector directions, the incremental adjoint problem (37) of design objective $Q$ for the directions of the eigenvectors, the adjoint problem (39) for chance constraint $f$, the generalized eigenvalue problems (48) for chance constraint $f$, the orthonormality constraint of the eigenvectors for chance constraint $f$, the incremental state problem (41) of chance constraint $f$ and the incremental adjoint problem (42) of chance constraint for the directions of the eigenvectors.
-To compute the design gradient, we need to calculate five sets of functional derivatives to solve for the corresponding ten sets of adjoint variables
-$\left(\mathbf{u}^{Q}\right)^{*},\left(\mathbf{u}^{f}\right)^{*},\left(\mathbf{v}^{Q}\right)^{*},\left(\mathbf{v}^{f}\right)^{*},\left(\psi_{j}^{Q}\right)^{*},\left(\psi_{j}^{f}\right)^{*},\left(\hat{\mathbf{u}}^{Q}_{j}\right)^{*},\left(\hat{\mathbf{u}^{f}_{j}}\right)^{*},\left(\hat{\mathbf{v}^{Q}_{j}}\right)^{*},\left(\hat{\mathbf{v}^{f}_{j}}\right)^{*}\}$ in (53), while the adjoint variable set for the eigenvalue set $\{{\lambda_{j}^{Q}}^{*},{\lambda_{j}^{f}}^{*}\}$ is not required due to its independency of the design parameter $d$.
-Setting the first variation of (53) with respect to the eigenvalue set $\{\lambda_{j}\}$ to zero isolates the terms corresponding to the quadratic approximation of the design objective and the orthonormality constraint of the eigenvectors, resulting in:
-$\{{\psi^{Q}_{j}}^{*}\}=\bigg{(}\frac{1}{2}+\beta_{V}\,\lambda_{j}^{Q}\bigg{)}\,\{\psi_{j}^{Q}\},\quad j=1,\cdots,N_{\mathrm{eig}}^{Q}.$ (54)
-Setting the first variations of (53) to zero with respect to $\mathbf{v}^{Q}_{j}$ and $\mathbf{u}^{Q}_{j}$ results in the incremental state problem:
-$\small\textit{Find $\hat{\mathbf{u}}^{Q}\in\mathcal{U}$, such that}\quad\big{<}\tilde{\mathbf{v}},\partial_{\mathbf{vu}}\bar{r}^{Q}\,\left(\mathbf{u}^{Q}_{j}\right)^{*}\big{>}=-\big{<}\tilde{\mathbf{v}},\partial_{\mathbf{v}m}\bar{r}^{Q}\,\left(\psi^{Q}_{j}\right)^{*}\big{>},\quad\forall\tilde{\mathbf{v}}\in\mathcal{V}.$ (55)
-and the incremental adjoint problem:
-$\footnotesize\begin{split}\textit{Find $\hat{\mathbf{v}}^{Q}\in\mathcal{V}$, such that}&\quad\\ \big{<}\tilde{\mathbf{u}},\partial_{\mathbf{uv}}\bar{r}^{Q}\,\left(\mathbf{v}^{Q}_{j}\right)^{*}\big{>}=-\big{<}\tilde{\mathbf{u}},\partial_{\mathbf{uu}}\bar{r}^{Q}\,{\left(\hat{\mathbf{u}}^{Q}_{j}\right)}^{*}+\partial_{\mathbf{uu}}\bar{Q}\,&{\left(\hat{\mathbf{u}}^{Q}_{j}\right)}^{*}+\partial_{\mathbf{u}m}\bar{r}^{Q}\,\left(\psi^{Q}_{j}\right)^{*}\big{>},\quad\forall\tilde{\mathbf{u}}\in\mathcal{U},\end{split}$ (56)
-similar to (36) and (37), respectively. In combination with (54), we can formulate the adjoint variables for the incremental state and adjoint problems individually for each eigenvector as
-$\{\left(\hat{\mathbf{u}}\right)^{*}_{j}\}$ $=\bigg{(}\frac{1}{2}+\beta_{V}\,\lambda^{Q}_{j}\bigg{)}\,\{\hat{\mathbf{u}}^{Q}_{j}\},\quad j=1,\cdots,N_{\mathrm{eig}}^{Q},$ (57)
-$\{\hat{\mathbf{v}}^{*}_{j}\}$ $=\bigg{(}\frac{1}{2}+\beta_{V}\,\lambda^{Q}_{j}\bigg{)}\,\{\hat{\mathbf{v}}^{Q}_{j}\},\quad j=1,\cdots,N_{\mathrm{eig}}^{Q}.$ (58)
-By vanishing the first variation of the Lagrangian functional (53) with respect to the adjoint or the state, we obtain the linear state problem: Find ${\mathbf{u}^{Q}}^{*}\in\mathcal{U}$, such that
-$\langle\tilde{\mathbf{v}},\partial_{\mathbf{vu}}\bar{r}^{Q}\,\left(\mathbf{u}^{Q}\right)^{*}\rangle=$ $-2\beta_{V}\langle\tilde{\mathbf{v}},\partial_{\mathbf{v}m}\bar{r}^{Q}\,(\mathcal{C}\partial_{m}\bar{r}^{Q})\rangle$ (59)
-$-\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\langle\tilde{\mathbf{v}},\partial_{\mathbf{v}mu}\bar{r}^{Q}\,\hat{\mathbf{u}}^{Q}_{j}\,\left(\psi^{Q}_{j}\right)^{*}+\partial_{\mathbf{v}mm}\bar{r}^{Q}\,\psi^{Q}_{j}\,\left(\psi^{Q}_{j}\right)^{*}\rangle$
-$-\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\langle\tilde{\mathbf{v}},\partial_{\mathbf{vuu}}\bar{r}^{Q}\,\hat{\mathbf{u}}^{Q}_{j}\,\left(\hat{\mathbf{u}}^{Q}_{j}\right)^{*}+\partial_{\mathbf{vu}m}\bar{r}^{Q}\,\psi^{Q}_{j}\,\left(\hat{\mathbf{u}}^{Q}_{j}\right)^{*}\rangle,\quad\forall\tilde{\mathbf{v}}\in\mathcal{V},$
-and linear adjoint problem: Find $\left(\mathbf{v}^{Q}\right)^{*}\in\mathcal{V}$, such that
-$\langle\tilde{\mathbf{u}},\partial_{\mathbf{uv}}\bar{r}^{Q}\,\left(\mathbf{v}^{Q}\right)^{*}\rangle=$ $-\langle\tilde{\mathbf{u}},\partial_{\mathbf{u}}\bar{Q}\rangle-2\beta_{V}\langle\tilde{\mathbf{u}},\partial_{\mathbf{u}m}\bar{r}^{Q}\,(\mathcal{C}\partial_{m}\bar{r}^{Q})\rangle$ (60)
-$-\langle\tilde{\mathbf{u}},\partial_{\mathbf{uu}}\bar{r}^{Q}\,\left(\mathbf{u}^{Q}\right)^{*}+\partial_{\mathbf{uu}}\bar{Q}\,\left(\mathbf{u}^{Q}\right)^{*}\rangle$
-$-\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\langle\tilde{\mathbf{u}},\partial_{\mathbf{u}m\mathbf{v}}\bar{r}^{Q}\,\hat{\mathbf{v}}_{j}\,\left(\psi^{Q}_{j}\right)^{*}$
-$+\partial_{\mathbf{u}m\mathbf{u}}\bar{r}^{Q}\,\hat{\mathbf{u}}_{j}\,\left(\psi^{Q}_{j}\right)^{*}+\partial_{\mathbf{u}mm}\bar{r}^{Q}\,\psi^{Q}_{j}\,\left(\psi^{Q}_{j}\right)^{*}\rangle$
-$-\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\langle\tilde{\mathbf{u}},\partial_{\mathbf{uvu}}\bar{r}^{Q}\,\hat{\mathbf{u}}^{Q}_{j}\,\left(\hat{\mathbf{v}}^{Q}_{j}\right)^{*}+\partial_{\mathbf{uv}m}\bar{r}^{Q}\,\psi^{Q}_{j}\,\left(\hat{\mathbf{v}}^{Q}_{j}\right)^{*}\rangle$
-$-\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\langle\tilde{\mathbf{u}},\partial_{\mathbf{uuv}}\bar{r}^{Q}\,\hat{\mathbf{v}}^{Q}_{j}\,\left(\hat{\mathbf{u}}^{Q}_{j}\right)^{*}+\partial_{\mathbf{uuu}}\bar{r}^{Q}\,\hat{\mathbf{u}}^{Q}_{j}\,\left(\hat{\mathbf{u}}^{Q}_{j}\right)^{*}$
-$+\partial_{\mathbf{uuu}}\bar{Q}\,\hat{\mathbf{u}}^{Q}_{j}\,\left(\hat{\mathbf{u}}^{Q}_{j}\right)^{*}+\partial_{\mathbf{uuu}m}\bar{r}^{Q}\,\psi^{Q}_{j}\,\left(\hat{\mathbf{u}}^{Q}_{j}\right)^{*}\rangle,$
-$\quad\forall\tilde{\mathbf{u}}\in\mathcal{U}.$
-By vanishing the first variation of (53) with respect to the eigenvalue set $\{\lambda_{j}^{f}\}$, we get
-$\{\left(\psi^{f}_{j}\right)^{*}\}=\bigg{(}\frac{1}{2}+\beta_{V}\,\lambda^{f}_{j}\bigg{)}\,\{\psi^{f}_{j}\},\quad j=1,\cdots,N^{f}_{\mathrm{eig}}.$ (61)
-Equating to zero of the first variations of (53) with respect to $\mathbf{v}_{j}^{f}$ and $\mathbf{u}_{j}^{f}$ results in the incremental state problem:
-$\textit{Find $\hat{\mathbf{u}}^{f}\in\mathcal{U}$, such that}\quad\big{<}\tilde{\mathbf{v}},\partial_{\mathbf{vu}}\bar{r}^{f}\,\mathbf{u}^{f*}_{j}\big{>}=-\big{<}\tilde{\mathbf{v}},\partial_{\mathbf{v}m}\bar{r}^{f}\,\psi^{{f}^{*}}_{j}\big{>},\quad\forall\tilde{\mathbf{v}}\in\mathcal{V}.$ (62)
-and linear adjoint problem:
-Find $\left(\mathbf{v}^{f}\right)^{*}\in\mathcal{V}$, such that (63)
-$\langle\tilde{\mathbf{v}},\partial_{\mathbf{uv}}\bar{r}^{f}\,\left(\mathbf{u}^{f}\right)^{*}\rangle=$ $-\sum_{n=1}^{N_{eig}^{f}}\langle\tilde{\mathbf{v}},\partial_{\mathbf{v}m\mathbf{u}}\bar{r}^{f}\hat{\mathbf{u}}_{n}^{f}{\left(\psi_{n}^{f}\right)}^{*}+\partial_{\mathbf{v}mm}\bar{r}^{f}\psi_{n}^{f}{\left(\psi_{n}^{f}\right)}^{*}\rangle$
-$-\sum_{n=1}^{N_{eig}^{f}}\langle\tilde{\mathbf{v}},\partial_{\mathbf{vuu}}\bar{r}^{f}\hat{\mathbf{u}}_{n}^{f}{\left(\hat{\mathbf{u}}_{n}^{f}\right)}^{*}+\partial_{\mathbf{vu}m}\bar{r}^{f}\psi_{n}^{f}{\left(\hat{\mathbf{u}}_{n}^{f}\right)}^{*}\rangle$
-$-\langle\tilde{\mathbf{v}},\partial_{\mathbf{v}m}\bar{r}^{f}m^{f}\rangle$
-$\quad\forall\tilde{\mathbf{u}}\in\mathcal{U}.$
-Finally, with driving all the adjoint variables, the $d$-gradient is derived as
-$\langle\tilde{d},D_{d}\mathcal{J}_{\mathrm{QUAD}}(d)\rangle=\langle\tilde{d},\partial_{d}\mathcal{L}_{\mathrm{QUAD}}\rangle.$ (64)
-The methodology for calculating the design gradient is detailed in Algorithm 2.
-The computational cost required to evaluate the quadratic approximation of the cost functional, denoted as $\mathcal{J}_{\mathrm{QUAD}}(d)$ in (52), includes solving one potentially nonlinear state partial differential equation (PDE), one linear adjoint problem, and $22(N_{\mathrm{eig}}^{Q}+N_{o})+2(N_{\mathrm{eig}}^{f}+N_{o})+$ 2 linear PDEs, as prescribed by the double-pass randomized algorithm presented in Algorithm 1.
-Furthermore, obtaining the $d$-gradient involves solving $2(N_{\mathrm{eig}}^{Q}+N_{\mathrm{eig}}^{f}+3)$ additional linear PDEs.
-1
-Input : Design objective (12) and weak form of the PDF (7).
-Output : Gradient of the quadratic approximation of the design objective $\mathcal{J}_{\mathrm{QUAD}}(d)$ with respect to the design parameter $z$.
-2
-3Solve for $m$-gradient (34);
-4
-5Solve for $m$-Hessian (38);
-6
-7Solve for $m_{f}$-gradient ;
-8
-9Solve for $m_{f}$-Hessian ;
-10
-11Solve for generalized eigenpairs with Algorithm 1;
-12
-13Compute $\mathcal{J}_{\mathrm{QUAD}}$ by (52);
-14
-15Solve the linear state problem (59);
-16
-17Solve the linear adjoint problem (63);
-18
-19Obtain the design gradient (64)
-Algorithm 2 Analytical design gradient
+In order to address the inequality constraint in (10), we convert the original chance-constrained optimization problem into the following unconstrained problem,
+$\underset{d\in[0,1]}{\text{min}}\mathcal{J}(d)+\tau_{\gamma}(\mathbb{E}[l_{\omega}(f)]-\alpha_{c}).$ (28)
+Here, $\tau_{\gamma}$ is a quadratic penalty method as outlined in [47].The quadratic penalty function is given by,
+$\tau_{\gamma}(x)=\frac{\gamma}{2}{(max{\{0,x\}})}^{2},$ (29)
+where $\gamma>0$ is a constant which determines the penalty weight.
+In this manner, the inequality constraint is replaced with a penalization term, that results in the transformation of the constrained problem into an unconstrained form.
+As $\gamma$ increases as per the continuation scheme in Algorithm 2, the penalized solution approaches the feasible region more strictly, making the unconstrained optimization problem effectively mimic the constrained form.
+4.3 Design gradient
+By substituting the trace estimators (23) into (18) and converting it into unconstrained problem using quadratic penalty (29) and continuous approximation (27), we have
+$\mathcal{J}_{\mathrm{QUAD}}(d)$ $=\bigg{(}\bar{Q}+\frac{1}{2}\sum_{j\geq 1}^{N_{\mathrm{eig}}}\lambda_{j}^{Q}\bigg{)}+\beta_{V}\bigg{(}\big{<}\bar{Q}^{m},\mathcal{K}\,\bar{Q}^{m}\big{>}+\frac{1}{2}\sum_{j\geq 1}^{N_{\mathrm{eig}}}\left({\lambda_{j}^{Q}}\right)^{2}\bigg{)}+\beta_{RG}\,R_{p}(d)$ (30)
+$+\tau_{\gamma}(\mathbb{E}[l_{\omega}(f_{\text{QUAD}}(m,d))]-\alpha_{c})$
+Implementing the INCG optimizer requires computing the gradient of the quadratic approximation of the design objective, $\mathcal{J}_{\mathrm{QUAD}}(d)$, with respect to the design parameter, referred to as the $d$-gradient.
+The Lagrangian formalism of the approximated cost, along with variational calculus principles for obtaining the necessary Lagrange multipliers, and an algorithm for computing the $d$-gradient are presented in Section B.
+The computational cost required to evaluate the quadratic approximation of the cost functional, denoted as $\mathcal{J}_{\mathrm{QUAD}}(d)$ in (30), includes solving one potentially nonlinear state PDE, one linear adjoint problem, and $22(N_{\mathrm{eig}}^{Q}+N_{o})+2(N_{\mathrm{eig}}^{f}+N_{o})+$ 2 linear PDEs, as prescribed by the double-pass randomized algorithm (see [24, 14]).
+Furthermore, obtaining the $d$-gradient involves solving $32(N_{\mathrm{eig}}^{Q}+N_{\mathrm{eig}}^{f})+$ 3 additional linear PDEs.
+The design gradient is needed for obtaining the optimal design through a continuation scheme as discussed in Algorithm 2.
 5 Numerical Results
-This section presents numerical experiments utilizing the proposed design under uncertainty framework for beam-insulation scenario in Figure 1 in both 2D and 3D.
-The objective is to achieve thermal insulation and mechanical stability avoiding stress concentration while mitigating uncertainty in the design process.
-The analyses presented in this section, investigate the accuracy and convergence of quadratic and Monte Carlo approximation of design objective and explores the effect of different limiting critical stress, critical chance, and uncertain parameter variance and spatial correlation on the optimal design, spatial pattern of $d_{\text{opt}}$, as well as the probability distributions of the design objectives $Q$.
-The parameters of the forward model are summarized in Table 1.
-The computational implementation of the framework relies on several open-source libraries like FEniCS [56] for the finite element solution of the forward model, SOUPy [57] for the quadratic approximation of the design objective and implementation of chance constraint, and hIPPYLib [54] for the Gaussian random field, trace estimator and INCG optimizer.
 Parameter Value
@@
 Table 1: List of the forward model parameters and their values used in the design problem.
+This section presents numerical experiments using the proposed framework for design under uncertainty for beam-insulation scenario in Figure 1 in both 2D and 3D.
+The objective is to achieve thermal insulation and mechanical stability avoiding stress concentration while mitigating uncertainty in the design process.
+The analyses presented in this section, investigate the accuracy and convergence of quadratic and Monte Carlo approximation of design objective and explores the effect of different limiting critical stress, critical chance, and uncertain parameter variance and spatial correlation on the optimal design, spatial pattern of $d_{\text{opt}}$, as well as the probability distributions of the design objectives $Q$.
+The parameters of the thermomechanical model are summarized in Table 1.
+The computational implementation of the framework relies on several open-source libraries like FEniCS [48] for the finite element solution of the thermomechanical model, SOUPy [49] for the quadratic approximation of the design objective and implementation of chance constraint, and hIPPYLib [46] for the Gaussian random field, trace estimator and INCG optimizer [^1] The codes and links to the dependent libraries for this work can be found at: https://github.com/pce-lab/InsulationDesignUncertainty.
+[…]
 5.1 Spatially-correlated uncertain parameter
-Refer to caption  Refer to caption  Refer to caption  Refer to caption
-Refer to caption  Refer to caption  Refer to caption  Refer to caption
-[…]
-[…]
-(a)                                              (b)
-Refer to caption            Refer to caption
-(c)                                              (d)
-Figure 5: The effect of boundary on the Gaussian random field in (11) for the scenario used in Figure 7: (a) marginal variance of the anisotropic uncertain parameter $m$, and point-wise correlation structure centered at (b) point (0.5,0.75) (c) point (0.25,0.75) (d) point (0.25,0.25).
-Coordinates $(x,y)$ are defined with respect to the bottom left corner of the beam-insulator domaim.
-The uncertain parameter $m$ is characterized by a mean value of $\bar{m}=0$, a variance of $\sigma^{2}=0.5^{2}$ and a correlation length of $L_{c}=0.25$.
 […]
 5.2 Quadratic and Monte Carlo approximations
-To assess the accuracy and efficiency of the quadratic approximation and Monte Carlo estimations of mean and variance of design objective with a control variate on the optimal design solution, we perform comparative analyses.
-As the sample size and number of dominant eigenvalues increase, both methods converge to similar mean and variance values, consistent with standard Monte Carlo estimations, as shown in Figure 6.
-For large sample sizes, the estimated mean and variance from the standard Monte Carlo method (Figure 6(a)) and the Monte Carlo method with control variate (Figure 6(b)) are nearly identical.
-However, the quadratic approximation as a control variate achieves lower error bounds with significantly fewer samples, as evident in the box plots.
-Furthermore, with only $N_{tr}=25$ trace estimators, the quadratic approximation closely matches the Monte Carlo results (Figure 6(c)), leading to significant savings in computational cost without loss in accuracy.
 Refer to caption  Refer to caption
@@
 Figure 5: Convergence plots corresponding to scenario in Figure 6 for the mean and variance of the design objective $Q$ : (a) standard Monte Carlo method (b) Monte Carlo with control variate and (c) Quadratic approximation.
-The uncertain parameter $m$ has the correlation length $L_{c}=0.05$ and a variance of $\sigma^{2}={0.25}^{2}$.
-Figures 7 and 8 compare the effects of different variances and correlation lengths of the uncertain parameter $m$ on the optimal design $d_{\text{opt}}$ indicating the spatial distribution of aerogel porosity and the probability distributions of the design objective $Q$ representing the thermal insulation performance of the thermal break.
+The uncertain parameter $m$ has the correlation length $L_{CR}=0.25$ and a variance of $\sigma^{2}={0.5}^{2}$.
+(a)                            (b)                            (c)
+Figure 6: The top row shows the optimal design obtained for $\sigma=0.25$ for the three different approaches: (a) Sampled Average approximation (500 samples) (b) Monte Carlo correction (100 samples) and (c) Quadratic approximation ($N_{eig}^{Q}=25$).
+The bottom row shows the QoI distribution of the design objective $Q$ for the three different cases above.
+The uncertain parameter $m$ has a correlation length $L_{CR}=0.05$ for all three cases.
+Figure 7: The top row shows the optimal design obtained for $\sigma=0.5$ for the three different approaches: (a) standard Monte Carlo using 500 samples, (b) Monte Carlo with control variate using 100 samples, and (c) Quadratic approximation using $N_{eig}^{Q}=25$.
+The bottom row shows the QoI distribution of the design objective $Q$ for the three different cases above.
+The uncertain parameter $m$ has a correlation length of $L_{CR}=0.25$ for all three cases.
+To evaluate the accuracy and efficiency of the quadratic approximation and Monte Carlo estimations with a control variate in optimal design, we perform comparative analyses.
+Figure 5 shows that, as the sample size and number of dominant eigenvalues increase, both methods converge to similar mean and variance values, aligning with standard Monte Carlo estimations For large sample sizes, the estimated mean and variance from standard Monte Carlo (Figure 5(a)) and Monte Carlo with a control variate (Figure 5(b)) are nearly identical.
+However, the quadratic approximation as a control variate achieves lower error bounds with significantly fewer samples, as shown in the box plots.
+Notably, with only $N_{eig}^{Q}=25$ trace estimators, the quadratic approximation closely matches Monte Carlo results (Figure 5(c)), substantially reducing computational cost while maintaining accuracy.
+Figures 6 and 7 compare the effects of different variances and correlation lengths of the uncertain parameter $m$ on the optimal design $d_{\text{opt}}$ indicating the spatial distribution of aerogel porosity and the probability distributions of the design objective $Q$ representing the thermal insulation performance of the thermal break.
 The optimal design is evaluated using standard Monte Carlo, Monte Carlo with control variate, and a quadratic Taylor approximation of the mean and variance of $Q$.
 In Figure 6 (low variance and short correlation length), the morphologies of $d_{\text{opt}}$ and probability distributions of $Q$ are nearly identical across all three methods.
-However, in Figure 8 (high variance and long correlation length), the quadratic approximation predicts different optimal design pattern with higher $Q$, while Monte Carlo with control variate yields results closely matching standard Monte Carlo but with significantly fewer samples (PDF solved).
-(a)                            (b)                            (c)
-Figure 7: The top row shows the optimal design obtained for $\sigma=0.25$ for the three different approaches: (a) Sampled Average approximation (500 samples) (b) Monte Carlo correction (100 samples) and (c) Quadratic approximation ($N_{tr}=25)$.
-The bottom row shows the QoI distribution of the design objective $Q$ for the three different cases above.
-The uncertain parameter $m$ has a correlation length $L_{c}=0.05$ for all three cases.
+However, in Figure 7 (high variance and long correlation length), the quadratic approximation predicts different optimal design pattern with higher $Q$, while Monte Carlo with control variate yields results closely matching standard Monte Carlo but with significantly fewer samples (PDEs solved).
 Interestingly, while the spatial pattern of the optimal design in Figure 6 appears nearly symmetric, the higher uncertainty in Figure 7 results in optimal designs that consistently position stronger material (lower aerogel porosity) along the horizontal axis, across all three approximation methods.
 This pattern reflects the anisotropy of the uncertain parameter, modeled with a higher correlation length along the layer deposition direction (see Figure 4).
-Also, the optimal designs at mean $\tilde{m}=0$ are much similar in Figure 7 as compared to Figure 8 where there is a deviation in the design patterns, which shows that there is a trade-off favoring the reduction of the uncertainty over mean performance for our design formulation (12).
+Also, the optimal designs at mean $\bar{m}=0$ are much similar in Figure 6 as compared to Figure 7 where there is a deviation in the design patterns, which shows that there is a trade-off favoring the reduction of the uncertainty over mean performance for our design formulation (7).
 These results demonstrate that the proposed optimization under uncertainty framework effectively captures spatial uncertainty, producing more reliable design solutions.
-Figure 8: The top row shows the optimal design obtained for $\sigma=0.5$ for the three different approaches: (a) standard Monte Carlo using 500 samples, (b) Monte Carlo with control variate using 100 samples, and (c) Quadratic approximation using $N_{tr}=25$.
-The bottom row shows the QoI distribution of the design objective $Q$ for the three different cases above.
-The uncertain parameter $m$ has a correlation length of $L_{c}=0.25$ for all three cases.
 5.3 Scalability with respect to number of design parameters
-We assess the scalability of the algorithm for PDE-constrained optimization under uncertainty with respect to the design dimension, determined by the number of nodes in the spatial discretization of $d$, and governed by two primary factors: Firstly, the eigenvalue decay rate of the covariance-preconditioned Hessian, estimated via the randomized trace estimator, and secondly the convergence behavior of the INCG optimizer.
-Figure 9 illustrates the behavior of the chance constraint function in (14) with the number of steps in the continuation scheme, the gradient norm 64 and the absolute values of the generalized eigenvalues 48 for the last step of chance convergence discussed in Algorithm 3.
-These metrics are presented for two design parameter dimensions, corresponding to the optimal design problem depicted in Figure 7 (c).
 (a)                                   (b)                                   (c)
 Figure 8: Optimization solution corresponding to Figure 6 (c) for different dimensions of design parameter: (a) Convergence of the chance $P(f(m,d)\geq 0)$ with the number of steps, (b) convergence of the norm of the design gradient with number of iterations of INCG optimizer, (c) eigenvalue decay of the covariance-preconditioned Hessian
+We assess the scalability of the algorithm for optimization under uncertainty with respect to the design dimension, determined by the number of nodes in the spatial discretization of $d$, and governed by two primary factors: Firstly, the eigenvalue decay rate of the covariance-preconditioned Hessian, estimated via the randomized trace estimator, and secondly the convergence behavior of the INCG optimizer.
+Figure 8 illustrates the behavior of the chance constraint function in (9) with the number of steps in the continuation scheme, the gradient norm 53 and the absolute values of the generalized eigenvalues for the last step of chance convergence discussed in Algorithm 2.
+These metrics are presented for two design parameter dimensions, corresponding to the optimal design problem depicted in Figure 6 (c).
 These results demonstrate that the quadratic approximation enables rapid decay of the critical chance, with the gradient norm converging below the tolerance threshold within 100 iterations for both mesh resolutions.
@@
 Additionally, Figure 8(c) shows a rapid eigenvalue decay, unaffected by the dimension of uncertain parameters, supporting the scalability of the randomized eigensolver used for trace approximation.
-1st                     2nd                     10th                     20th
-Refer to caption  Refer to caption  Refer to caption  Refer to caption
-40th                   60th                   80th                   100th
-Figure 10: Eigenfunctions of the covariance-preconditioned Hessian of the design objective corresponding to eigenvalue decay for the scenario in Figure 7 (c).
-Figure 10 depicts the eigenfunctions of the covariance-preconditioned Hessian for the design objective.
-Eigenvectors associated with larger eigenvalues capture broader spatial scales of the design parameter, while those linked to smaller eigenvalues become increasingly oscillatory and less significant.
-Together, these dimension-independent factors suggest that leveraging quadratic Taylor approximations and stochastic optimization techniques enables a scalable approach for PDE-constrained optimal design under uncertainty with the enforcement of chance constraints, regardless of the dimensionality of the uncertain parameters.
+Together, these dimension-independent factors suggest that leveraging quadratic Taylor approximations and stochastic optimization techniques enables a scalable approach for optimal design under high-dimensional uncertainty with the enforcement of chance constraints.
 5.4 Optimal design: effect of chance constraint
 (a)                              (b)                              (c)
-Figure 11: The top row shows the optimal design results obtained at the mean of the uncertain sample $\tilde{m}=0$ for different values of critical chance: (a) $\alpha_{c}=0.05$ (b) $\alpha_{c}=0.08$ and (c) $\alpha_{c}=0.1$.
+Figure 9: Effect of the critical chance $\alpha_{c}$ on optimal design.
+The top row shows the optimal design results obtained at the mean of the uncertain sample $\tilde{m}=0$ for different values of critical chance: (a) $\alpha_{c}=0.05$ (b) $\alpha_{c}=0.08$ and (c) $\alpha_{c}=0.1$.
 The bottom row represents the von Mises stress for the corresponding optimal designs in the top row.
-In all results $T_{cr}=22.5\>\text{MPa},L_{c}=0.25,\sigma=0.25,\beta_{V}=0.1$ and $\beta_{R}=1\times 10^{-5}$.
+In all results $T_{cr}=22.5\>\text{MPa},L_{CR}=0.25,\sigma=0.25,\beta_{V}=0.1$ and $\beta_{RG}=1\times 10^{-5}$.
 (a)                              (b)                              (c)
-Figure 12: The top row shows the optimal design results obtained at the mean of the uncertain sample $\tilde{m}=0$ for different values of limiting critical stress: (a) $T_{cr}=20\>\text{MPa}$ (b) $T_{cr}=22.5\>\text{MPa}$ and (c) $T_{cr}=25\>\text{MPa}$.
+Figure 10: Effect of the limiting critical stress $T_{cr}$ on optimal design.
+The top row shows the optimal design results obtained at the mean of the uncertain sample $\tilde{m}=0$ for different values of limiting critical stress: (a) $T_{cr}=20\>\text{MPa}$ (b) $T_{cr}=22.5\>\text{MPa}$ and (c) $T_{cr}=25\>\text{MPa}$.
 The bottom row represents the von Mises stress for the corresponding optimal designs in the top row.
-In all results $\alpha_{c}=0.05,L_{c}=0.25,\sigma=0.25,\beta_{V}=0.1$ and $\beta_{R}=1\times 10^{-5}$.
+In all results $\alpha_{c}=0.05,L_{CR}=0.25,\sigma=0.25,\beta_{V}=0.1$ and $\beta_{RG}=1\times 10^{-5}$.
 The chance constraint function defined in (9) incorporates parameters such as $\alpha_{c}$ and $T_{cr}$, which govern the enforcement of constrain to mitigate stress concentrations within the domain.
@@
 The critical stress limit was set at $T_{cr}=22.5\>\text{MPa}$, while $\alpha_{c}$ was varied among 0.05, 0.08, and 0.1.
-This figure presents the optimal design, and von Mises stress distribution evaluated at the mean of the uncertain parameter $\tilde{m}=0$.
+This figure presents the optimal design, and von Mises stress distribution evaluated at the mean of the uncertain parameter $\bar{m}=0$.
 The results indicate that increasing critical chance $\alpha_{c}$ expands the regions of mechanically stronger material (highlighted in red), effectively reducing localized stresses across the component.
@@
 The rows show the porosity $\phi_{f}$, solid displacement magnitude $||\mathbf{u}_{s}||$, magnitude of effective solid stress $||\mathbf{T}_{s}||$, , solid temperature $\theta_{s}$, fluid temperature $\theta_{f}$ and fluid pressure $p$.
-The limiting critical stress, $T_{cr}$, defines the maximum permissible von Mises stress within the beam structure.
-Figure 12 demonstrates the influence of $T_{cr}$ on the optimal designs, evaluated at the mean of the uncertain parameter $\tilde{m}=0$, along with the corresponding von Mises stress distributions.
+The limiting critical stress, $T_{cr}$, defines the maximum permissible von Mises stress for the beam structure.
+Figure 10 shows the influence of $T_{cr}$ on the optimal designs, evaluated at the mean of the uncertain parameter $\bar{m}=0$, along with the corresponding von Mises stress distributions.
 The findings indicate that increasing the critical stress $T_{cr}$ results in the allocation of stronger insulating material (depicted in blue) across a larger region of the domain.
@@
 Thus, a higher value of $T_{cr}$ improves insulation efficiency at the expense of reduced mechanical stability.
-Figure 13 shows the state parameters and porosity field evaluated at the mean of the uncertain parameter $\tilde{m}=0$ and three samples for the optimal design in Figure 12 (b).
-The states include the solid displacement magnitude $||\mathbf{u}_{s}||$, magnitude of effective solid stress $||\mathbf{T}_{s}||$, , solid temperature $\theta_{s}$, fluid temperature $\theta_{f}$ and fluid pressure $p$.
+Figure 11 shows the states and porosity field evaluated at $\bar{m}=0$ and three samples for the scenario in Figure 10 (b).
+The states include the solid displacement magnitude $||\mathbf{u}_{s}||$, effective solid stress $||\mathbf{T}_{s}||$, solid temperature $\theta_{s}$, fluid temperature $\theta_{f}$ and fluid pressure $p$.
 5.5 Three-dimensional design of beam-insulator system
@@
 (e) (f)
-Figure 14: Convergence of critical chance $\alpha_{c}$ with the results obtained for optimal design at mean $\tilde{m}=0$ for steps in adaptive optimization discussed in Algorithm 3 from step 1 to step 5 shown in (b)-(f).
-The uncertain parameter $m$ has a correlation length of $L_{c}=0.25$ and a variance of $\sigma^{2}={0.1}^{2}$,the critical chance is $\alpha_{c}=0.1$, and the critical stress considered is $T_{cr}=15\>\text{MPa}$.
+Figure 12: Evolution of design pattern with continuation scheme.
+Optimal design obtained at each step of the continuation scheme with increasing $\omega$ as shown in Figure 3.
+Random field as initial condition is shown in (a) and the results obtained for optimal design at mean $\tilde{m}=0$ for steps in adaptive optimization from step 1 to step 5 shown in (b)-(f).
+The uncertain parameter $m$ has a correlation length of $L_{CR}=0.25$ and a variance of $\sigma^{2}={0.1}^{2}$,the critical chance is $\alpha_{c}=0.1$, and the critical stress considered is $T_{cr}=12.5\>\text{MPa}$.
 We demonstrate the efficiency and scalability of the proposed optimization under uncertainty framework by applying it to a three-dimensional design problem.
@@
 In this case, a uniform circular load is also applied to the middle of the top surface, simulating the influence of additional building envelope assemblies.
-The uncertain parameter $m$ is modeled with a horizontal correlation length of $L_{c}=0.25$ and a variance of $\sigma^{2}={0.1}^{2}$.
-For the chance constraint, we set a critical chance of $\alpha_{c}=0.1$ and a critical stress of $T_{cr}=15\>\text{MPa}$.
-The finite element discretization of the 3D domain resulted in 389,044 design parameters.
+The uncertain parameter $m$ is modeled with a horizontal correlation length of $L_{CR}=0.25$ and a variance of $\sigma^{2}={0.1}^{2}$.
+For the chance constraint, we set a critical chance of $\alpha_{c}=0.1$ and a critical stress of $T_{cr}=12.5\>\text{MPa}$.
+The finite element discretization of the 3D domain resulted in 919,464 design parameters.
 Refer to caption Refer to caption
 (a) (b)
-Figure 15: Final design obtained after five steps of the continuation scheme, achieving convergence of the critical chance $\alpha_{c}$.
-The 3D domain’s finite element discretization, resulted in 389044 design parameters.
-Figure 14 illustrates the optimal designs obtained at each step of the continuation scheme outlined in Algorithm 3.
+Figure 13: Optimal design patterns for the 3D scenario.
+Final design obtained after five steps of the continuation scheme, achieving convergence of the critical chance $\alpha_{c}$.
+The 3D domain’s finite element discretization, resulted in 919464 design parameters.
+Figure 12 illustrates the optimal designs obtained at each step of the continuation scheme for a random field as initial condition, with increasing $\omega$ as shown in Figure 3.
+We tested different initial conditions and arrived at the same final pattern, indicating that for the scenario represented in Figure 12, we achieve global optima.
 Regions shown in red correspond to material with low porosity ($\phi_{f}=0.1$).
-As the continuation scheme progresses, the red regions gradually decrease, reducing mechanical support in the domain and increasing the p-norm of von Mises stress.
+As the continuation scheme progresses, the red regions start to form, increasing mechanical support in the domain and decreasing the p-norm of von Mises stress.
 Consequently, the probability in (10) approaches the critical chance threshold of $\alpha_{c}=0.1$.
-By step 5, the final design, shown in Figure 15, emerges, featuring a column-like structure of low-porosity material $\phi_{f}=0.1$, providing support aligned with the direction of the applied external load.
+By step 5, the final design, shown in Figure 13, emerges, featuring a column-like structure of low-porosity material $\phi_{f}=0.1$, providing support aligned with the direction of the applied external load, and also along the left wall to provide support against the stress due to thermal expansion of the beam.
 6 Conclusions
-This study introduces an efficient and scalable computational framework for chance-constrained optimal design under high-dimensional uncertainty governed by partial differential equations (PDEs).
+This study introduces an efficient and scalable computational framework for chance-constrained optimal design under high-dimensional uncertainty governed by finite element models.
 The framework optimizes the spatial distribution of a spatially correlated and uncertain porosity field in additively manufactured aerogel thermal breaks to maximize thermal insulation performance while mitigating stress concentrations.
@@
 Furthermore, it would enable efficient approximation of both the PDE state and its sensitivities with respect to the design parameters, facilitating scalable and computationally feasible gradient-based optimization for complex systems.
-This study addresses the computational challenges of PDE-constrained optimization under high-dimensional uncertainty by leveraging intrinsic dimensionality and incorporating chance constraints to prevent stress concentrations.
+This study addresses the computational challenges of optimization under high-dimensional uncertainty by leveraging intrinsic dimensionality and incorporating chance constraints to prevent stress concentrations.
 The framework enhances optimal design reliability, offering a robust and scalable approach while paving the way for future advancements to refine and extend its capabilities.
+Data Availability Statement
+Some or all data, models, or code generated or used during the study are available in a repository online.
 Acknowledgments
@@
 National Science Foundation (NSF) CAREER Award CMMI-2143662.
-The authors express their appreciation for the helpful discussions with Dr.
-Peng Chen, from Georgia Institute of Technology, regarding the problem formulation and software libraries.
-Additionally, the authors would like to acknowledge the support provided by the Center for Computational Research at the University at Buffalo.
-Appendix A Adaptive optimization
-Input : Initial design parameter $d_{0}$, smoothing parameter $\omega_{0}$, penalty parameter $\gamma_{0}$, power parameters $\sigma_{\omega}$ and $\sigma_{\gamma}$, $\hat{f}=f_{\text{QUAD}}(d)$
-Output : Design gradient $d_{opt}$
+The authors express their appreciation for Dr.Peng Chen, from Georgia Institute of Technology for valuable discussions on problem formulation and software libraries.
+Furthermore, the authors acknowledge the support provided by the Center for Computational Research at the University at Buffalo.
+Appendix A Gradient and Hessian action of design objective and chance constraint
+This section outlines the evaluations of the $m$-gradient, $m^{f}$-gradient, $m$-Hessian, and $m^{f}$-Hessian at the mean of the uncertain parameters $\bar{m}$, which are essential for constructing the quadratic Taylor approximations of the design objective and chance constraint, as described in Section 3.1.
+$m$-gradient: The gradient $Q^{m}(\bar{m})$ acting in the direction $\tilde{m}=m-\bar{m}$ is expressed as [24],
+$\langle\tilde{m},\bar{Q}^{m}\rangle=\langle\tilde{m},\partial_{m}\bar{Q}+\partial_{m}{\bar{r}^{Q}}\rangle,$ (31)
+where $r^{Q}=r(\mathbf{u},\mathbf{v}^{Q},m,d)$ and the adjoint variable $\mathbf{v}^{Q}$ is considered as the Lagrange multiplier associated with the state equation with respect to the objective function $Q$.
+Here $(\bar{\cdot})$ is the quantities at the mean of $m$, and $(\tilde{\cdot})$ indicates directional variables referred to the variables that define the direction along which perturbation is considered to compute the change in the functional to obtain the derivative of the functional.
+In order to obtain the $m$-gradient, we need to obtain the multipliers $\mathbf{u}$ and $\mathbf{v}^{Q}$, which can be obtained using the state problem,
+[…]
+and the adjoint problem,
+[…]
+$m$-Hessian: The Hessian $Q^{mm}(\bar{m})$ acting on $\hat{m}$ can be expressed as [24],
+$\begin{split}\langle\tilde{m},\bar{Q}^{mm}\,\hat{m}^{Q}\rangle=&\langle\tilde{m},\partial_{m\mathbf{v}}\bar{r}^{Q}\,\hat{\mathbf{v}}^{Q}+\partial_{m\mathbf{u}}\bar{r}^{Q}\,\hat{\mathbf{u}}^{Q}+\partial_{m\mathbf{u}}\bar{Q}\,\hat{\mathbf{u}}^{Q}+\partial_{mm}\bar{r}^{Q}\,\hat{m}^{Q}\\ +\partial_{mm}\bar{Q}\,\hat{m}^{Q}\rangle.\end{split}$ (34)
+Thus, computing $m$-Hessian requires evaluation of the Lagrange multipliers $\hat{\mathbf{u}}^{Q}$ and $\hat{\mathbf{v}}^{Q}$ by solving the incremental state problem:
+[…]
+and the incremental adjoint problem:
+Find $\hat{\mathbf{v}}^{Q}$, such that
+$\langle\tilde{\mathbf{u}},\partial_{\mathbf{uv}}\bar{r}^{Q}\,\hat{\mathbf{v}}^{Q}\rangle=-\langle\tilde{\mathbf{u}},\partial_{\mathbf{uu}}\bar{r}^{Q}\,\hat{\mathbf{u}}+\partial_{\mathbf{uu}}\bar{Q}\,\hat{\mathbf{u}}^{Q}+\partial_{\mathbf{u}m}\bar{r}^{Q}\,\hat{m}^{Q}+\partial_{\mathbf{u}m}\bar{Q}\,\hat{m}^{Q}\rangle.$ (36)
+$m^{f}$-gradient: The gradient of $f$ at $\bar{m}$ acting in the direction $\tilde{m}$ defined as [21]:
+$\langle\tilde{m},\bar{f}^{m}\rangle=\langle\tilde{m},\partial_{m}\bar{f}+\partial_{m}\bar{r}^{f}\rangle$ (37)
+The computation of $m^{f}$-gradient requires the lagrange multiplier $\mathbf{v}^{f}$ which is computed through the adjoint problem as:
+Find $\mathbf{v}^{f}$, such that
+$\langle\tilde{\mathbf{u}},\partial_{\mathbf{u}}\bar{r}^{f}\rangle=-\langle\tilde{\mathbf{u}},\partial_{\mathbf{u}}\bar{f}\rangle,$ (38)
+$m^{f}$-Hessian: The action of Hessian of $f$ can be computed at $\tilde{m}$ as [21]:
+$\begin{split}\langle\tilde{m},\bar{f}^{mm}\hat{m}^{f}\rangle=\langle\tilde{m},\partial_{m\mathbf{v}}\bar{r}^{f}\hat{\mathbf{v}}^{f}+\partial_{m\mathbf{u}}\bar{r}^{f}\hat{\mathbf{u}}^{f}+\partial_{m\mathbf{u}}\bar{f}\hat{\mathbf{u}}^{f}+\partial_{mm}\bar{r}^{f}\hat{m}^{f}+\partial_{mm}\bar{f}\>\hat{m}^{f}\rangle\quad.\end{split}$ (39)
+Thus, the solution of $m^{f}$-Hessian requires the computation of Lagrange multipliers $\hat{\mathbf{u}}^{f}$ and $\hat{\mathbf{v}}^{f}$, by solving the incremental state problem,
+[…]
+and the incremental adjoint problem,
+[…]
+Also, in order to compute $m^{f}$-Hessian, we solve a generalized eigenvalue problem in (22) for a finite number of eigenvalues $N^{f}_{\mathrm{eig}}$.
+Appendix B Gradient of design functional
+The Lagrangian formalism is employed to derive the $d$-gradient by considering a meta-Lagrangian functional
+$\mathcal{L}_{\mathrm{QUAD}}$ $(d,\mathbf{u},\mathbf{v}^{*},\{{\lambda_{j}^{Q}}\},\{{\psi_{j}}^{Q}\},\{{\hat{\mathbf{u}}_{j}^{Q}}\},\{{\hat{\mathbf{v}}_{j}}^{Q}\},$ (42)
+$\left(\mathbf{u}^{Q}\right)^{*},\left(\mathbf{v}^{Q}\right)^{*},\{\left(\lambda_{j}^{Q}\right)^{*}\},\{\left(\psi_{j}^{Q}\right)^{*}\},\{\left(\hat{\mathbf{u}}_{j}^{Q}\right)^{*}\},\{\left(\hat{\mathbf{v}}_{j}^{Q}\right)^{*}\},$
+$\{{\lambda_{j}^{f}}\},\{{\psi_{j}}^{f}\},\{{\hat{\mathbf{u}}_{j}}^{f}\},\{{\hat{\mathbf{v}}_{j}}^{f}\},\left(\mathbf{u}^{f}\right)^{*},\left(\mathbf{v}^{f}\right)^{*},\{\left(\lambda_{j}^{f}\right)^{*}\},\{\left(\psi_{j}^{f}\right)^{*}\},\{\left(\hat{\mathbf{u}}_{j}^{f}\right)^{*}\},\{\left(\hat{\mathbf{v}}_{j}^{f}\right)^{*}\})$
+$=\mathcal{J}_{\mathrm{QUAD}}(d)+\langle\mathbf{v}^{*},\partial_{\mathbf{v}}\bar{r}^{Q}\rangle$
+$+\tau_{\gamma}(\mathbb{E}[l_{\omega}(f_{\text{QUAD}}(m,d))]-\alpha_{c})$
+$+\langle\left(\mathbf{u}^{f}\right)^{*},\partial_{\mathbf{u}}\bar{r}_{f}+\partial_{\mathbf{u}}\bar{f}\rangle$
+$+\sum^{N_{\mathrm{eig}}^{f}}_{j=1}\langle\left(\psi_{j}^{f}\right)^{*},(\bar{f}^{mm}-{\lambda_{j}^{f}}\mathcal{K}^{-1})\psi_{j}^{f}\rangle$
+$+\sum^{N_{\mathrm{eig}}^{f}}_{j=1}\left(\lambda_{j}^{f}\right)^{*}(\langle{\psi_{j}}^{f},\mathcal{K}^{-1}{\psi_{j}}^{f}\rangle-1)+\sum^{N_{\mathrm{eig}}^{f}}_{j=1}\langle\left(\hat{\mathbf{v}}_{j}^{f}\right)^{*},\partial_{\mathbf{vu}}\bar{r}_{f}\,{\hat{\mathbf{u}}_{j}}^{f}+\partial_{\mathbf{v}m}\bar{r}_{f}\,{\psi_{j}}^{f}\rangle$
+$+\sum^{N_{\mathrm{eig}}^{f}}_{j=1}\langle\left(\hat{\mathbf{u}}_{j}^{f}\right)^{*},\partial_{\mathbf{uv}}\bar{r}_{f}\,{\hat{\mathbf{v}}_{j}}^{f}+\partial_{\mathbf{uu}}\bar{r}_{f}\,\left(\mathbf{u}^{f}\right)^{*}+\partial_{\mathbf{uu}}\bar{f}\,{\hat{\mathbf{u}}_{j}}^{f}+\partial_{\mathbf{u}m}\bar{f}\,{\psi_{j}}^{f}+\partial_{\mathbf{u}m}\bar{r}_{f}\,{\psi_{j}}^{f}\rangle$
+$+\langle\left(\mathbf{u}^{Q}\right)^{*},\partial_{\mathbf{u}}\bar{r}^{Q}+\partial_{\mathbf{u}}\bar{Q}\rangle$
+$+\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\langle\left(\psi_{j}^{Q}\right)^{*},(\bar{Q}^{mm}-{\lambda_{j}^{Q}}\mathcal{K}^{-1})\psi_{j}^{Q}\rangle$
+$+\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\left(\lambda_{j}^{Q}\right)^{*}(\langle{\psi_{j}}^{Q},\mathcal{K}^{-1}{\psi_{j}}^{Q}\rangle-1)+\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\langle{\hat{\mathbf{v}}_{j}}^{{Q}^{*}},\partial_{\mathbf{vu}}\bar{r}_{Q}\,{\hat{\mathbf{u}}_{j}}^{Q}+\partial_{\mathbf{v}m}\bar{r}_{Q}\,{\psi_{j}}^{Q}\rangle$
+$+\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\langle\left(\hat{\mathbf{u}}_{j}^{Q}\right)^{*},\partial_{\mathbf{uv}}\bar{r}_{Q}\,{\hat{\mathbf{v}}_{j}}^{Q}+\partial_{\mathbf{uu}}\bar{r}_{Q}\,\left(\mathbf{u}^{Q}\right)^{*}+\partial_{\mathbf{uu}}\bar{Q}\,{\hat{\mathbf{u}}_{j}}^{Q}+\partial_{\mathbf{u}m}\bar{Q}\,{\psi_{j}}^{Q}+\partial_{\mathbf{u}m}\bar{r}_{Q}\,{\psi_{j}}^{Q}\rangle$
+where $(\cdot^{*})$ represents the adjoint variables, $(\hat{\cdot})$ denotes the incremental variables.
+The terms on the right hand side of (42), correspond to the following: the quadratic approximation of the design objective (18), the state and adjoint problems for chance constraint $f$ and design objective $Q$, the generalized eigenvalue problems for chance constraint $f$ and design objective $Q$, the incremental state and adjoint problems for chance constraint $f$ and design objective $Q$.
+For obtaining the design gradient, we need to compute five sets of functional derivatives to solve for the corresponding ten sets of adjoint variables,
+$\left(\mathbf{u}^{f}\right)^{*},\left(\mathbf{u}^{Q}\right)^{*},\left(\mathbf{v}^{f}\right)^{*},\left(\mathbf{v}^{Q}\right)^{*},\left(\psi_{j}^{f}\right)^{*},\left(\psi_{j}^{Q}\right)^{*},\left(\hat{\mathbf{u}^{f}_{j}}\right)^{*},\left(\hat{\mathbf{u}}^{Q}_{j}\right)^{*},\left(\hat{\mathbf{v}^{f}_{j}}\right)^{*},\left(\hat{\mathbf{v}^{Q}_{j}}\right)^{*}\}$ in (42), while the adjoint variable set for the eigenvalue set $\{{\lambda_{j}^{f}}^{*},{\lambda_{j}^{Q}}^{*}\}$ is not required due to its independency of the design parameter $d$.
+Once we have defined the functional $\mathcal{L}_{\mathrm{QUAD}}$, we need to apply the condition that the first variation of Lagrangian as zero with respect to Lagrange multipliers for both design objective $Q$ and chance constraint function $f$.
+We start by equating the variation of (42) as zero with respect to $\{\lambda_{j}^{f}\}$,
+$\{\left(\psi^{f}_{j}\right)^{*}\}=\frac{1}{2}\bigg{(}1+2\beta_{V}\,\lambda^{f}_{j}\bigg{)}\,\{\psi^{f}_{j}\},\quad j=1,\cdots,N^{f}_{\mathrm{eig}}.$ (43)
+Next, we equate the first variation of (42) as zero with respect to $\mathbf{v}_{j}^{f}$ and $\mathbf{u}_{j}^{f}$ which gives the incremental state problem,
+$\textit{Find $\hat{\mathbf{u}}^{f}$, such that}\quad\big{<}\tilde{\mathbf{v}},\partial_{\mathbf{vu}}\bar{r}^{f}\,\mathbf{u}^{f*}_{j}\big{>}=-\big{<}\tilde{\mathbf{v}},\partial_{\mathbf{v}m}\bar{r}^{f}\,\psi^{{f}^{*}}_{j}\big{>}.$ (44)
+and linear adjoint problem:
+Find $\left(\mathbf{v}^{f}\right)^{*}$, such that (45)
+$\langle\tilde{\mathbf{v}},\partial_{\mathbf{uv}}\bar{r}^{f}\,\left(\mathbf{u}^{f}\right)^{*}\rangle=$ $-\sum_{n=1}^{N_{eig}^{f}}\langle\tilde{\mathbf{v}},\partial_{\mathbf{v}m\mathbf{u}}\bar{r}^{f}\hat{\mathbf{u}}_{n}^{f}{\left(\psi_{n}^{f}\right)}^{*}+\partial_{\mathbf{v}mm}\bar{r}^{f}\psi_{n}^{f}{\left(\psi_{n}^{f}\right)}^{*}\rangle$
+$-\sum_{n=1}^{N_{eig}^{f}}\langle\tilde{\mathbf{v}},\partial_{\mathbf{vuu}}\bar{r}^{f}\hat{\mathbf{u}}_{n}^{f}{\left(\hat{\mathbf{u}}_{n}^{f}\right)}^{*}+\partial_{\mathbf{vu}m}\bar{r}^{f}\psi_{n}^{f}{\left(\hat{\mathbf{u}}_{n}^{f}\right)}^{*}\rangle$
+$-\langle\tilde{\mathbf{v}},\partial_{\mathbf{v}m}\bar{r}^{f}m^{f}\rangle.$
+Similarly, we set the first variation of the (42) to zero for the Lagrange multipliers related to design objective $Q$.
+Firstly, we equate the first variation of (42) w.r.t eigenvalues $\{\lambda^{Q}_{j}\}$ to zero,
+$\{{\psi^{Q}_{j}}^{*}\}=\frac{1}{2}\bigg{(}1+2\beta_{V}\,\lambda_{j}^{Q}\bigg{)}\,\{\psi_{j}^{Q}\},\quad j=1,\cdots,N_{\mathrm{eig}}^{Q}.$ (46)
+Furthermore, we let the first variation of (42) as zero w.r.t $\mathbf{v}^{Q}_{j}$ and $\mathbf{u}^{Q}_{j}$ leading to the incremental state problem:
+$\textit{Find $\hat{\mathbf{u}}^{Q}$, such that}\quad\big{<}\tilde{\mathbf{v}},\partial_{\mathbf{vu}}\bar{r}^{Q}\,\left(\mathbf{u}^{Q}_{j}\right)^{*}\big{>}=-\big{<}\tilde{\mathbf{v}},\partial_{\mathbf{v}m}\bar{r}^{Q}\,\left(\psi^{Q}_{j}\right)^{*}\big{>}.$ (47)
+and the incremental adjoint problem:
+$\begin{split}\textit{Find $\hat{\mathbf{v}}^{Q}$, such that}&\quad\\ \big{<}\tilde{\mathbf{u}},\partial_{\mathbf{uv}}\bar{r}^{Q}\,\left(\mathbf{v}^{Q}_{j}\right)^{*}\big{>}=-\big{<}\tilde{\mathbf{u}},\partial_{\mathbf{uu}}\bar{r}^{Q}\,{\left(\hat{\mathbf{u}}^{Q}_{j}\right)}^{*}+\partial_{\mathbf{uu}}\bar{Q}\,&{\left(\hat{\mathbf{u}}^{Q}_{j}\right)}^{*}+\partial_{\mathbf{u}m}\bar{r}^{Q}\,\left(\psi^{Q}_{j}\right)^{*}\big{>},\end{split}$ (48)
+similar to (35) and (36), respectively. In combination with (46), we can express the adjoint variables for both the incremental state and adjoint problems separately for each eigenvector,
+$\{\left(\hat{\mathbf{u}}\right)^{*}_{j}\}$ $=\frac{1}{2}\bigg{(}1+2\beta_{V}\,\lambda^{Q}_{j}\bigg{)}\,\{\hat{\mathbf{u}}^{Q}_{j}\},\quad j=1,\cdots,N_{\mathrm{eig}}^{Q},$ (49)
+$\{\hat{\mathbf{v}}^{*}_{j}\}$ $=\frac{1}{2}\bigg{(}1+2\beta_{V}\,\lambda^{Q}_{j}\bigg{)}\,\{\hat{\mathbf{v}}^{Q}_{j}\},\quad j=1,\cdots,N_{\mathrm{eig}}^{Q}.$ (50)
+Finally, we equate the first variation of the Lagrangian (42) to zero with respect to adjoint or state, we obtain the linear state problem:
+Find ${\mathbf{u}^{Q}}^{*}$, such that
+$\langle\tilde{\mathbf{v}},\partial_{\mathbf{vu}}\bar{r}^{Q}\,\left(\mathbf{u}^{Q}\right)^{*}\rangle=$ $-2\beta_{V}\langle\tilde{\mathbf{v}},\partial_{\mathbf{v}m}\bar{r}^{Q}\,(\mathcal{K}\partial_{m}\bar{r}^{Q})\rangle$ (51)
+$-\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\langle\tilde{\mathbf{v}},\partial_{\mathbf{v}mu}\bar{r}^{Q}\,\hat{\mathbf{u}}^{Q}_{j}\,\left(\psi^{Q}_{j}\right)^{*}+\partial_{\mathbf{v}mm}\bar{r}^{Q}\,\psi^{Q}_{j}\,\left(\psi^{Q}_{j}\right)^{*}\rangle$
+$-\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\langle\tilde{\mathbf{v}},\partial_{\mathbf{vuu}}\bar{r}^{Q}\,\hat{\mathbf{u}}^{Q}_{j}\,\left(\hat{\mathbf{u}}^{Q}_{j}\right)^{*}+\partial_{\mathbf{vu}m}\bar{r}^{Q}\,\psi^{Q}_{j}\,\left(\hat{\mathbf{u}}^{Q}_{j}\right)^{*}\rangle,$
+and linear adjoint problem:
+Find $\left(\mathbf{v}^{Q}\right)^{*}$, such that
+$\langle\tilde{\mathbf{u}},\partial_{\mathbf{uv}}\bar{r}^{Q}\,\left(\mathbf{v}^{Q}\right)^{*}\rangle=$ $-\langle\tilde{\mathbf{u}},\partial_{\mathbf{u}}\bar{Q}\rangle-2\beta_{V}\langle\tilde{\mathbf{u}},\partial_{\mathbf{u}m}\bar{r}^{Q}\,(\mathcal{K}\partial_{m}\bar{r}^{Q})\rangle$ (52)
+$-\langle\tilde{\mathbf{u}},\partial_{\mathbf{uu}}\bar{r}^{Q}\,\left(\mathbf{u}^{Q}\right)^{*}+\partial_{\mathbf{uu}}\bar{Q}\,\left(\mathbf{u}^{Q}\right)^{*}\rangle$
+$-\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\langle\tilde{\mathbf{u}},\partial_{\mathbf{u}m\mathbf{v}}\bar{r}^{Q}\,\hat{\mathbf{v}}_{j}\,\left(\psi^{Q}_{j}\right)^{*}$
+$+\partial_{\mathbf{u}m\mathbf{u}}\bar{r}^{Q}\,\hat{\mathbf{u}}_{j}\,\left(\psi^{Q}_{j}\right)^{*}+\partial_{\mathbf{u}mm}\bar{r}^{Q}\,\psi^{Q}_{j}\,\left(\psi^{Q}_{j}\right)^{*}\rangle$
+$-\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\langle\tilde{\mathbf{u}},\partial_{\mathbf{uvu}}\bar{r}^{Q}\,\hat{\mathbf{u}}^{Q}_{j}\,\left(\hat{\mathbf{v}}^{Q}_{j}\right)^{*}+\partial_{\mathbf{uv}m}\bar{r}^{Q}\,\psi^{Q}_{j}\,\left(\hat{\mathbf{v}}^{Q}_{j}\right)^{*}\rangle$
+$-\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\langle\tilde{\mathbf{u}},\partial_{\mathbf{uuv}}\bar{r}^{Q}\,\hat{\mathbf{v}}^{Q}_{j}\,\left(\hat{\mathbf{u}}^{Q}_{j}\right)^{*}+\partial_{\mathbf{uuu}}\bar{r}^{Q}\,\hat{\mathbf{u}}^{Q}_{j}\,\left(\hat{\mathbf{u}}^{Q}_{j}\right)^{*}$
+$+\partial_{\mathbf{uuu}}\bar{Q}\,\hat{\mathbf{u}}^{Q}_{j}\,\left(\hat{\mathbf{u}}^{Q}_{j}\right)^{*}+\partial_{\mathbf{uu}m}\bar{r}^{Q}\,\psi^{Q}_{j}\,\left(\hat{\mathbf{u}}^{Q}_{j}\right)^{*}\rangle.$
+The computation of $d$-gradient means that we have to compute how the cost functional $\mathcal{J}_{\mathrm{QUAD}}$ changes with small variations in $d$.
+This can be done by computing the Fréchet derivative of $\mathcal{J}_{\mathrm{QUAD}}$ w.r.t $d$, which can be expressed as $D_{d}\mathcal{J}_{\mathrm{QUAD}}(d)$.
+However, the computation of this Fréchet derivative $D_{d}\mathcal{J}_{\mathrm{QUAD}}(d)$ becomes computationally prohibitive when the design parameter $d$ is high-dimensional.
+Hence, we compute the gradient through the partial derivative of the lagrangian $\mathcal{L}_{\mathrm{QUAD}}$ with respect to design parameter $d$, which is $\partial_{d}\mathcal{L}_{\mathrm{QUAD}}$, as it only requires the solution of lagrange multipliers, which are independent of the dimension of design parameter $d$.
+Also, since the total derivative of $\mathcal{J}_{\mathrm{QUAD}}$ along a random direction $\tilde{d}$ is equal to the partial derivative of lagrangian $\mathcal{L}_{\mathrm{QUAD}}$ w.r.t $d$, as it depends on design parameter only.
+Therefore, the $d$-gradient can be derived as,
+$\langle\tilde{d},D_{d}\mathcal{J}_{\mathrm{QUAD}}(d)\rangle=\langle\tilde{d},\partial_{d}\mathcal{L}_{\mathrm{QUAD}}\rangle.$ (53)
 1
-2while $||d_{k}-d_{k-1}||\leq\mathcal{\epsilon_{\text{out}}}$ or $k<k_{\text{max}}$; $d_{k+1}=\text{Inexact Newton Conjugate-Gradient}(d_{k},\mathcal{J}(d),\nabla_{d}\mathcal{J}(d),\epsilon_{\text{in}})$.; Evaluate approximate chance $\hat{f}_{k+1}$ at $d_{k+1}$.; Update $\omega_{k+1}=\sigma_{\omega}\omega_{k}$, $\gamma_{k+1}=\sigma_{\gamma}\gamma_{k}$;
-3end while
-4Obtain the design gradient $d_{opt}$;
-Algorithm 3 Adaptive Optimization
+Input : Design objective (7) and weak form of the PDE (4).
+Output : Gradient of the quadratic approximation of the design objective $\mathcal{J}_{\mathrm{QUAD}}(d)$ with respect to the design parameter $d$.
+2
+3Compute the $m$-gradient (31);
+4
+5Compute the $m$-Hessian (34);
+6
+7Compute the $m_{f}$-gradient (37);
+8
+9Compute the $m_{f}$-Hessian (39);
+10
+11Obtain the generalized eigenpairs using the double-pass randomized algorithm;
+12
+13Solve for $\mathcal{J}_{\mathrm{QUAD}}$ by (30);
+14
+15Find the solution to the linear state problem (51);
+16
+17Find the solution to the linear adjoint problem (45);
+18
+19Solve for the $d$-gradient (53)
+Algorithm 1 Analytical design gradient
+Appendix C Adaptive optimization
 The gradient-based scheme uses a continuation scheme that gradually increases the smoothing parameter $\omega$ and penalty parameter $\gamma$ to achieve convergence [21].
 An initial value of the smoothing parameter $\omega_{0}$ and the penalty parameter $\gamma_{0}$ is defined for the continuation scheme.
-These parameters are updated at the end of each iteration using scaling parameters $\sigma_{\omega}$ and $\sigma_{\gamma}$ for smoothing and penalty parameters, respectively.
-For $k^{th}$ iteration , the updated smoothing parameter is $\omega_{k+1}=\sigma_{\omega}\omega_{k}$ and the updated penalty parameter is $\gamma_{k+1}=\sigma_{\gamma}\gamma_{k}$.
+These parameters are updated at the end of each iteration using scaling parameters $\Lambda_{\omega}$ and $\Lambda_{\gamma}$ for smoothing and penalty parameters, respectively.
+For $k^{th}$ iteration , the updated smoothing parameter is $\omega_{k+1}=\Lambda_{\omega}\omega_{k}$ and the updated penalty parameter is $\gamma_{k+1}=\Lambda_{\gamma}\gamma_{k}$.
 In an outer loop, we update the parameters $\omega$ and $\gamma$, whereas an Inexact Newton Conjugate-Gradient (INCG) is applied in the inner loop to solve the optimization problem (28).
-The termination criterion for the outer loop is if the maximum number of iterations $k_{\text{max}}$ is reached.
+The termination criterion for the outer loop is if the maximum number of iterations $k_{\text{M}}$ is reached.
 For each iteration, the approximated chance is computed through the quadratic approximation $f_{\text{QUAD}}$ of the chance function as given in (9).
-The output is the optimal value of design parameter $d_{\text{opt}}$.
-The algorithm for this adaptive optimization is given in Algorithm 1.
+The output is the optimal value of design parameter $d_{\text{opm}}$.
+Input : Initial design parameter $d_{0}$, smoothing parameter $\omega_{0}$, penalty parameter $\gamma_{0}$, scaling parameters $\Lambda_{\omega}$ and $\Lambda_{\gamma}$, $\hat{f}=f_{\text{QUAD}}(d)$
+Output : Optimal Design $d_{opm}$
+1
+2while $||d_{k}-d_{k-1}||\geq\mathcal{\epsilon_{\text{out}}}$ or $k<k_{\text{M}}$; $d_{k+1}=\text{Inexact Newton Conjugate-Gradient}(d_{k},\mathcal{J}(d),\nabla_{d}\mathcal{J}(d),\epsilon_{\text{in}})$.; Evaluate approximate chance $\hat{f}_{k+1}$ at $d_{k+1}$.; Update $\omega_{k+1}=\Lambda_{\omega}\omega_{k}$, $\gamma_{k+1}=\Lambda_{\gamma}\gamma_{k}$;
+3end
+4Obtain the optimal design $d_{opm}$;
+Algorithm 2 Adaptive Optimization